Bibliography

article

23/03-04 16:19

[aagedal2001a] Jan Øyvind Aagedal. Quality of Service Support in Development of Distributed Systems. PhD thesis, Department for Informatics, University of Oslo, June 2001.

[abadi1992a] Martín Abadi and Leslie Lamport. An old-fashioned recipe for real time. SRC Research Reports 91, DEC’s System Research Center, October 1992. (Presented at REX Workshop, The Netherlands 1991, Real-Time: Theory in Practice).

[abadi1993a] Martín Abadi and Leslie Lamport. Conjoining specifications. SRC Research Reports 118, DEC’s System Research Center, December 1993.

[abbadi1989a] Amr El Abbadi and Sam Toueg. The group paradigm for concurrency control protocols. IEEE Transactions on Knowledge and Data Engineering, 1(3):376–386, September 1989.

[abbott1990a] Russel J. Abbott. Resourceful systems for fault tolerance, reliability, and safety. ACM Computing Surveys, 22(1):36–68, March 1990.

[accetta1986a] Mike Accetta, Robert Baron, William Bolosky, David Golub, Richard Rashid, Avadis Tevanian, and Michael Young. Mach: A new kernel foundation for UNIX development. In Proceeding of the USENIX 1986 Summer Conference, pages 93– 112, Atlanta, 1986. USENIX.

[adam1994a] Joel F. Adam, Michael Ismert Henry H. Houh, and David L. Tennenhouse. A network architecture for distributed multimedia systems. In Proceedings of the International Conference on Multimedia Computing and Systems, Boston, MA, May 1994.

[agha1990a] Gul Agha. The structure and semantics of actor languages. In J. W. de Bakker, W. P. de Roever, and G. Rozenberg, editors, The Proceedings of REX School/Workshop Foundations of Object-Oriented Languages, volume 489 of Lecture Notes in Computer Science, pages 1–59, Noordwijkerhout, The Netherlands, May/June 1990. Springer-Verlag.

[agha1998a] Gul Agha. Compositional development from reusable components requires connectors for managing both protocols and resources. In Workshop on Compositional Software Architectures, Monterey, California, January 1998. (position paper).

[ahanger1995a] G. Ahanger, D. Benson, and T. D. C. Little. Video query formulation. In IS&T/SPIE Symposium on Electronic Imaging Science and Technology, Storage and Retrieval for Image and Video Databases III, volume 2420, pages 280–291, San Jose, February 1995.

[ahuja1996a] R. Ahuja, S. Keshav, and H. Saran. Design, implementation, and performance of a native-mode ATM transport layer. In Proceeding IEEE INFOCOM ’96, March 1996.

[ait-kaci1991a] Hassan Aït-Kaci and Andreas Podelski. Towards a meaning of LIFE. Technical Report R-R 11, Digital Paris Research Laboratory, June 1991.

[akselsen1993a] Sigmund Akselsen, Arne Kjetil Eidsvik, and Trine Folkow. Telemedicine and ISDN. IEEE Communications Magazine, pages 46–51, January 1993.

[aksit1994a] Mehmet Aksit, Jan Bosch, William van der Sterren, and Lodewijk Bergmans. Realtime specification inheritance anomalies and real-time filters. In Mario Tokoro and Remo Pareschi, editors, Proceedings of the Eighth European Conference on Object-Oriented Programming, volume 821 of Lecture Notes in Computer Science, pages 386–407, Bologna, Italy, July 1994. Springer-Verlag.

[aksit1998a] Mehmet Aksit and Bedir Tekinerdogan. Solving the modeling problems of objectoriented languages by composing multiple aspects using composition filters. In AOP’98 workshop position paper, 1998.

[alles1995a] Anthony Alles. ATM internetworking. Technical report, Cisco Systems, Inc., May 1995.

[alur1991a] Rajeev Alur and David L. Dill. The theory of timed automata. In J. W. de Bakker, C. Huizing, W. P. de Roever, and G. Rozenberg, editors, Proceedings of the REX Workshop 1991, Real-Time: Theory in Practice, volume 600 of Lecture Notes in Computer Science, pages 45–73, Mook, The Netherlands, June 1991. Springer-Verlag.

[alur1991b] Rajeev Alur and Thomas A. Henzinger. Logics and models of real time: A survey. In J. W. de Bakker, C. Huizing, W. P. de Roever, and G. Rozenberg, editors, Proceedings of the REX Workshop 1991, Real-Time: Theory in Practice, volume 600 of Lecture Notes in Computer Science, pages 74–106, Mook, The Netherlands, June 1991. Springer-Verlag.

[alur1994a] Rajeev Alur and David L. Dill. A theory of timed automata. Theoretical Computer Science, 126:183–235, 1994.

[amenyo1993a] John-Thones Amenyo, Aurel A. Lazar, and Giovanni Pacifici. Proactive cooperative scheduling and buffer managment for multimedia networks. Multimedia Systems, 1(1):37–49, 1993.

[analyti1995a] Anastasia Analyti and Stavros Christodoulakis. Multimedia object modelling and content-based querying. In Advanced Course: Multimedia Databases in Perspective, pages 213–237, June 1995.

[ancona1997a] Massimo Ancona, Walter Cazzola, Gabriella Dodero, and Vittoria Gianuzzi. Communication modeling by channel reification. In Workshop on Advances in Languages for User Modeling, in Sixth International Conference on User Modeling, pages 1–9, Chia Laguna, Sardinia Italia, June 1997.

[anderson1995a] Ross Anderson and Roger Needham. Programming satan’s computer. In Jan van Leeuwen, editor, Computer Science Today — Recent Trends and Developments, volume 1000 of Lecture Notes in Computer Science, pages 426–440. Springer-Verlag, 1995.

[andrews1983a] Gregory R. Andrews and Fred B. Schneider. Concepts and notations for concurrent programming. ACM Computing Surveys, 15(1):3–43, March 1983.

[andrews1987a] Gregory R. Andrews, Richard D. Schlichting, Roger Hayes, and Titus D. M. Purdin. The design of the Saguaro distributed operating system. IEEE Transactions on Software Engineering, SE-13(1):104–118, January 1987.

[anger1989a] Frank D. Anger. On Lamport’s interprocess communication model. ACM Transactions on Programming Languages and Systems, 11(3):404–417, July 1989.

[anshus1991a] Otto J. Anshus. Introduction to MacroScope — a distributed operating system. Technical report, Department of Computer Science, University of Tromsø, April 1991.

[apteker1995a] Ronnie T. Apteker, James Fisher, Valentin S. Kisimov, and Hanoch Neishlos. Video acceptability and frame rate. IEEE Multimedia, 2(3):32–40, Fall 1995.

[armbruster1992a] Heinrich Armbrüster and Klaus Wimmer. Networks: High-performance computing, high-capacity storage, and high-speed communication. IEEE Journal on Selected Areas in Communications, 10(9):1382–1396, December 1992.

[arnold1996a] David Arnold, Andy Bond, Martin Chilvers, and Richard Taylor. Hector: Distributed objects in Python. In Proceeding of the 4th International Python Conference, Livermore, California, June 1996.

[arnold1997a] David Arnold, Andy Bond, and Martin Chilvers. Hector: Distributed objects in Python. Dr. Dobb’s Sourcebook on Distributed Objects, pages 13–18, January/February 1997.

[artymiak2000a] Jacek Artymiak. Python programming for beginners. Linux Journal – Strictly On-Line, (73), May 2000.

[ashford1998a] Colin Ashford. The TINA service composition architecture. In Workshop on Compositional Software Architectures, Monterey, California, January 1998. (position paper).

[asperti1993a] Andrea Asperti. Linear logic, comonads and optimal reductions. Technical report, Dipartimento di Matematica, Universitá di Bologna, September 1993.

[asplin1993a] Jo Asplin. Distributed parallel volume rendering of atmospheric model output. In Visualisering’93, Trondheim, 1993.

[assenmacher1993a] Holger Assenmacher, Thomas Breitbach, Peter Buhler, Volker Hubsch, and Reinhard Schwarz. PANDA — supporting distributed programming in C++. In Oscar M. Nierstraz, editor, ECOOP’93 — The 7th European Conference on Object-Oriented Programming, volume 707 of Lecture Notes in Computer Science, pages 361–383, Kaiserlautern, Germany, July 1993. Springer-Verlag.

[assumpcao1994a] Jecel Mattos de Assumpção, Jr. and Sergio Takeo Kufuji. Bootstrapping the object oriented operating system Merlin: Just add reflection. Technical report, Laboratorio de Sistemas Integraveis, University of Sao Paulo, Brazil, 1994.

[aurrecoechea1995a] Cristina Aurrecoechea, Andrew Campbell, Alexandros Eleftheriadis, and Hisaya Hadama. Meeting end-to-end QoS challenges for adaptive digital video flows. Lancaster University Report MPG-95-06, Center for Telecommunication Research, Columbia University, 1995.

[aurrecoechea1995b] Cristina Aurrecoechea, Andrew Campbell, and Linda Hauw. A review of quality of service architectures. ACM Multimedia Systems Journal, November 1995. Lancaster University Report MPG-95-10.

[aycock1998a] John Aycock. Compiling little languages in python. In Proceedings of the 7th International Python Conference, pages 66–77, 1998.

[babaoglu1981a] Özalp Babaoğlu and William Joy. Converting a swap-based system to do paging in an architecture lacking page-referenced bits. In Proceedings of the 8th Symposium on Operating System Principles. ACM, 1981.

[babaoglu1988a] Özalp Babaoğlu, Pat Stephenson, and Rogério Drummond. Reliable broadcast and communication models: Tradeoffs and lower bounds. Distributed Computing, 2:177–189, 1988.

[babaoglu1991a] Özalp Babaoğlu, Lorenzo Alvisi, Alessandro Amoroso, and Renzo Davoli. Paralex: An environment for parallel programming in distributed systems. Technical Report UBLCS-91-01, Department of Mathematics, University of Bologna, February 1991.

[babaoglu1993a] Özalp Babaoğlu and Keith Marzullo. Consistent global states of distributed systems: Fundamental concepts and mechanisms. Technical Report UBLCS-93-1, Laboratory for Computer Science, University of Bologna, January 1993.

[babaoglu1993b] Özalp Babaoğlu and Sam Toueg. Understanding non-blocking atomic commitment. Technical Report UBLCS-93-2, Laboratory for Computer Science, University of Bologna, January 1993.

[babaoglu1995a] Özalp Babaoğlu, Renzo Davoli, Luigi Giachini, and Paolo Sabattini. The inherent cost of strong-partial view-synchronous communication. In Jean-Michel Hélary and Michel Raynal, editors, Distributed Algorithms, 9th International Workshop, WDAG ’95, volume 972 of Lecture Notes in Computer Science, pages 72–86, Le Mont-Saint-Michel, France, September 1995. Springer-Verlag.

[bacon1995a] Jean Bacon, John Bates, Richard Hayton, and Ken Moody. Using events to build distributed applications. In Proceedings of the 2nd International Conference on Services for Distributed and Networked Environments, 1995.

[baiceanu1996a] Veronica Baiceanu, Crispin Cowan, Dylan McNamee, Calton Pu, and Jonathan Walpole. Multimedia applications require adaptive CPU scheduling. In Workshop on Resource Allocation Problems in Multimedia Systems, Washington DC, 1996.

[baker1992a] Rusti Baker, Alan Downing, Kate Finn, and Earl Rennison. Multimedia processing model for a distributed multimedia i/o system. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 164–175, La Jolla, California, USA, November 1992. Springer-Verlag.

[bakker1997a] Arno Bakker, Ihor Kuz, and Maarten van Steen. Towards a taxonomy of distributed-object models. In Proceeding of the Third Annual ASCI Conference, pages 22–27, Heijen, The Netherlands, June 1997.

[bal1989a] Henri E. Bal, Jennifer G. Steiner, and Andrew S. Tanenbaum. Programming languages for distributed computing systems. ACM Computing Surveys, 21(3):261– 322, September 1989.

[bal1991a] Henri E. Bal. A comparative study of five paralell programming languages. In Distributed Open Systems in Perspective — Proceedings of the Spring 1991 EurOpen Conference, pages 209–228, Tromsø, Norway, May 1991. EurOpen.

[ballesteros1998a] Franscisco J. Ballesteros, Sergio Arévalo, Fabio Kon, and Roy H. Campbell. Towards a grand unified framework for mobile objects. In III ECOOP Workshop on Mobility and Replication, Brussels, Belgium, July 1998.

[ballesteros1999a] Franscisco J. Ballesteros, Christopher K. Hess, Fabio Kon, Sergio Arévalo, and Roy H. Campbell. Object-orientation in Off++ – a distributed adaptable microkernel. In 2nd ECOOP Workshop on Object-Orientation and Operating Systems, Lisbon, Portugal, June 1999.

[banerjea1994a] Anindo Banerjea, Domenico Ferrari, Bruce A. Mah, Mark Moran, Dinesh C. Verma, and Hui Zhang. The Tenet real-time protocol suite: Design, implementation, and experiences. Technical Report TR-94-059, International Computer Science Institute, Berkeley, CA, November 1994.

[barr1991a] Michael Barr. Accessible categories and models of linear logic. Technical report, Department of Mathematics and Statistics, McGill University, Montreal, Quebec, Canada, June 1991.

[barr1992a] Michael Barr. Terminal coalgebras in well-founded set theory. Technical report, Department of Mathematics and Statistics, McGill University, Montreal, Quebec, Canada, 1992.

[barr1992b] Michael Barr. Functional set theory. Technical report, Department of Mathematics and Statistics, McGill University, Montreal, Quebec, Canada, 1992.

[barr1992c] Michael Barr. HSP type theorems in the category of posets. Technical report, Department of Mathematics and Statistics, McGill University, Montreal, Quebec, Canada, 1992.

[barr1992d] Michael Barr. Algebraically compact functors. Technical report, Department of Mathematics and Statistics, McGill University, Montreal, Quebec, Canada, 1992.

[barr1992e] Michael Barr. Functional semantics and HSP type theorems. Technical report, Department of Mathematics and Statistics, McGill University, Montreal, Quebec, Canada, 1992.

[barr1992f] Michael Barr. Fuzzy models of linear logic. Technical report, Department of Mathematics and Statistics, McGill University, Montreal, Quebec, Canada, 1992.

[barr1992g] Michael Barr. *-autonomous categories and linear logic. Technical report, Department of Mathematics and Statistics, McGill University, Montreal, Quebec, Canada, 1992.

[bartoli1992a] Alberto Bartoli. Segment and object managment in Pegasus. Technical report, Computer Laboratory, University of Cambridge and Faculty of Computer Science, University of Twente, September 1992. Draft.

[barzilai1997a] Tsipora Barzilai, Dilip Kandlur, Ashish Mehra, Debanjan Saha, and Steve Wise. Design and implementation of an RSVP-based quality of service architecture for integrated services internet. In Proceedings of the 17th International Conference on Distributed Systems, Baltimore, Maryland, USA, May 1997. IEEE Computer Society Press.

[bates1994b] John Bates and Jean Bacon. A development platform for multimedia applications in a distributed, ATM network environment. In Proceedings of the 1st International Conference on Multimedia Computing and Systems, 1994.

[bates1995a] John Bates and Jean Bacon. Supporting interactive presentation of distributed multimedia. Multimedia Tools and Applications, 1(1):47–78, 1995.

[batlivala19??a] Nariman Batlivala, Barry Gleeson, Jim Hamrick, Scott Lurndal, Darren Price, James Soddy, and Vadim Abrossimov. Experience with SVR4 over Chorus. Unisys Corporation and Chorus Systèmes.

[baugsto1989a] Bjørn Arild W. Baugstø and Jarle Fredrik Greipsland. Parallel sorting methods for large data volumes on a hypercube database computer. In Proceedings of International Workshop on Database Machines, IWDM’89. Springer-Verlag, 1989.

[baumgartner1993a] Gerald Baumgartner and Vincent F. Russo. Signatures: A C++ extension for type abstraction and subtype polymorphism. Technical Report CSD-TR-93-059, Department of Computer Science, Purdue University, September 1993.

[bearman1993a] Mirion Y. Bearman. ODP-trader. In J. De Meer, B. Mahr, and S. Storp, editors, Open Distributed Processing II, Proceedings of the IFIP TC6/WG6.1 International Conference on Distributed Processing, IFIP Transactions, volume C-20, pages 37– 51, Berlin, Germany, September 1993. Elsevier Science B. V.

[beaudry1996a] Don Beaudry. MESS Tutorial, 1996.

[becker1997a] Christian R. Becker and Kurt Geihs. MAQS — management for adaptive QoSenabled services. In IEEE Workshop on Middleware for Distributed Real-Time Systems and Services, San Francisco, USA, 1997.

[becker1998a] Christian R. Becker and Kurt Geihs. Quality of service — aspects of distributed programs. In International Workshop on Aspect-Oriented Programming at ICSE’98, Kyoto, Japan, 1998.

[beeri1990a] Catriel Beeri. A formal approach to object-oriented databases. Data & Knowledge Engineering, 5:353–382, 1990.

[bellini1995a] A. Bellini, D. Giuli, and F Pirri. Experimental multimedia broadband telecommunication service for tele-education application. In F. Williams, H. Brä, and J. Nolan, editors, Proceedings of the Fourth International Conference Broadband Island ’95, Global Broadband & Beyond, pages 163–, Dublin, Ireland, September 1995.

[bellman1995a] Alessandro Bellman. A system for multi-media: The first ring. In F. Williams, H. Brä, and J. Nolan, editors, Proceedings of the Fourth International Conference Broadband Island ’95, Global Broadband & Beyond, pages 313–324, Dublin, Ireland, September 1995.

[bellovin1989a] S. M. Bellovin. Security problems in the TCP/IP protocol suite. ACM Computer Communication Review, 19(2):32–48, 1989.

[bennett1995a] Franzer Bennett and Andy Hopper. Developments in networked multimedia. In Proceedings NEC Research Symposium in Multimedia Computing, June 1995.

[berger1992a] Emery Berger. FP + OOP = Haskell. Technical report, Department of Computer Science, The University of Texas at Austin, March 1992.

[bernhardt1995a] Christoph Bernhardt and Ernst Biersack. The server array: A novel architecture for a scalable video server. Technical report, Institut Eurécom, February 1995.

[berra1992a] P. B. Berra, C. Y. R. Chen, A. Ghafoor, and T. D. C. Little. Issues in networking and data management of distributed multimedia systems. In Symposium on High-Performance Distributed Computing, Syracuse NY, September 1992.

[bershad1990a] Brian N. Bershad, Thomas E. Anderson, Edward D. Lazowska, and Henry M. Levy. Lightweight remote procedure call. ACM Transactions on Computer Systems, 8(1):37–55, February 1990.

[berson1995a] Steven Berson, Leana Golubchik, and Richard R. Muntz. Fault tolerant design of multimedia servers. In Michael J. Carey and Donovan A. Schneider, editors, Proceedings of the 1995 ACM SIGMOD International Conference on Management of Data, pages 364–375, San Jose, California, May 1995.

[bettati1995a] R. Bettati, D. Ferrari, A. Gupta, W. Heffner, W. Howe, M. Moran, Q. Nguyen, and R. Yavatkar. Connection establishment for multi-party real-time communication. In Proceedings of the Fifth International Workshop on Network and Operating System Support for Digital Audio and Video, Durham, NH, April 1995.

[betz1991a] David Betz. Your own tiny object-oriented language. Dr. Dobb’s Journal, pages 26–33, 86–89, September 1991.

[beuche1997a] Danilo Beuche. An approach for managing highly configurable operating systems. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Object-Orientation and Operating Systems, volume 1357 of Lecture Notes in Computer Science, pages 531–536, Jyväskylä, Finland, June 1997. Springer-Verlag.

[beugnard1999a] Antoine Beugnard and Robert Ogor. Encapsulation of protocols and services in medium components to build distributed applications. In Engineering Distributed Objects Workshop, ICSE’99, Los Angeles, May 1999. Position Paper.

[birman1987a] Kenneth P. Birman and Thomas A. Joseph. Reliable communication in the presence of failures. ACM Transactions on Computer Systems, 5(1):47–76, February 1987.

[birman1987b] Ken Birman and Thomas Joseph. Exploiting virtual synchrony in distributed systems. Operating Systems Review, 21(5):123–138, 1987. (11th ACM Symposium on Operating System Principles, ACM SIGOPS).

[birman1989a] Ken Birman and Keith Marzullo. ISIS and the META project. Sun Technology, pages 90–104, Summer 1989.

[birman1991a] Kenneth P. Birman. The process group approach to reliable distributed computing. Technical report, Department of Computer Science, Cornell University, July 1991.

[birman1991b] Kenneth P. Birman. ISIS prblems with solutions, 1991.

[birman1991c] Ken Birman. Some notes on threads in ISIS, September 1992.

[birman1991d] Kenneth P. Birman, Robert Cooper, and Barry Gleeson. Programming with process groups: Group and multicast semantics. Technical Report TR-91-1185, Department of Computer Science, Cornell University, January 1991.

[birrell1982a] Andrew D. Birrell, Roy Levin, Roger M. Needham, and Michael D. Schroeder. Grapevine: And exercise in distributed computing. Communications of the ACM, 25(4):260–274, April 1982.

[birrell1984a] Andrew D. Birrell and Bruce Jay Nelson. Implementing remote procedure call. ACM Transactions on Computer Systems, 2(1):39–59, February 1984.

[birrell1986a] Andrew. D. Birrell, Butler W. Lampson, Roger M. Needham, and Michael D. Schroeder. A global authentication service without global trust. In Proceedings of the 1986 IEEE Symposium on Security and Privacy, pages 223–230, Oakland, California, April 1986.

[birrell1987a] A. D. Birrell, J. V. Guttag, J. J. Horning, and R. Levin. Synchronization primitives for a multiprocessor: A formal specification. Operating Systems Review, 21(5):94– 102, 1987.

[birrell1989a] Andrew D. Birrell. An introduction to programming with threads. SRC Research Reports 35, DEC’s System Research Center, January 1989.

[black1987a] Andrew Black, Norman Hutchinson, Eric Jul, Henry Levy, and Larry Carter. Distribution and abstract types in Emerald. IEEE Transactions on Software Engineering, SE-13(1):65–76, January 1987.

[black1990a] David L. Black. Scheduling support for concurrency and parallelism in the Mach operating system. Computer, pages 35–43, May 1990.

[blair1985a] Gordon S. Blair, Jon R. Malone, and John A. Mariani. A critique of UNIX. Software—Practice and Experience, 15(12):1125–1139, December 1985.

[blair1992a] Gordon S. Blair, Geoff Coulson, Nigel Davis, and Neil Williams. The role of operating systems in object-oriented distributed multimedia platforms. In Proceedings of the Second ACM/IEEE Workshop on Object Orientation in Operating Systems, Dourdan, France, September 1992.

[blair1992b] Gordon S. Blair, Nigel Davis, Geoff Coulson, and Neil Williams. Abstractions for continuous media in hetrogeneous distributed systems: The role of groups. In Proceedings of the 5th ACM SIGOPS Workshop on Models and Paradigms for Distributed Systems Structuring, Mont Saint-Michel, France, September 1992.

[blair1992c] G. Blair, G. Coulson, P. Auzimour, L. Hazard, F. Horn, and J. B. Stefani. An integrated platform and computational model for open distributed multimedia applications. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 223–236, La Jolla, California, USA, November 1992. Springer-Verlag.

[blair1993a] Gordon Blair, Lynne Blair, Howard Bowman, and Amanda Chetwynd. Formal support for the specification and construction of distributed multimedia systems (the Tempo project), final project deliverable. Lancaster University Report MPG-93-23, School of Engineering, Computing and Mathematical Sciences, Lancaster University, December 1993.

[blair1993b] Gordon S. Blair, Geoff Coulson, and Nigel Davis. System support for multimedia applications: An assessment of the state of the art. Lancaster University Report MPG-93-29, Distributed Multimedia Research Group, Department of Computing, Lancaster University, 1993.

[blair1993c] Gordon Blair, Andrew Campbell, Geoff Coulson, Francisco Garcia, David Hutchinson, Andrew Scott, and Doug Shepherd. A network interface unit to support continuos media. IEEE Journal on Selected Areas in Communications, 11(2):264–275, February 1993.

[blair1994a] G. S. Blair, G. Coulson, M. Papathomas, Philippe Robin, J. B. Stefani, F. Horn, and L. Hazard. A hybrid approach to real-time synchronisation in distributed multimedia systems. Lancaster University Report MPG-94-21, Distributed Multimedia Research Group, Department of Computing, Lancaster University, 1993.

[blair1994b] G. S. Blair, M. Papathomas, G. Coulson, Philippe Robin, J. B. Stefani, F. Horn, and L. Hazard. Supporting real-time multimedia behaviour in opend dstributed systems: An approach based on synchronous languages. Lancaster University Report MPG-94-13, Distributed Multimedia Research Group, Department of Computing, Lancaster University / Centre National d’Etude des Telecommunications, France, April 1994.

[blair1996a] Gordon S. Blair, Geoff Coulson, Michael Papathomas, Philippe Robin, Jean-Bernard Stefani, F. Horn, and Laurent Hazard. A programming model and system infrastructure for real-time synchronization in distributed multimedia systems. IEEE Journal on Selected Areas in Communications, 14(1):249–263, January 1996.

[blair1997a] Gordon Blair. Notes on reflective middleware. Technical report, Department of Computing, Lancaster University, 1997.

[blair1997b] Gordon Blair, Lynne Blair, and Jean-Bernard Stefani. A specification architecture for multimedia systems in open distributed processing. Computer Networks and ISDN Systems, 29(Special Issue on Specification Architecture):473–500, 1997.

[blair1997c] Gordon Blair, Geoff Coulson, Nigel Davies, Philippe Robin, and Tom Fitzpatrick. Adaptive middleware for mobile multimedia applications. In Proceedings of the 8th International Workshop on Network and Operating System Support for Digital Audio and Video (NOSSDAV ’97), St. Louis, Missouri, U.S, May 1997.

[blair1998c] Gordon S. Blair, Geoff Coulson, Philippe Robin, and Michael Papathomas. An architecture for next generation middleware. In Middleware’98, September 1998.

[blair1998e] Lynne Blair and Gordon S. Blair. The impact of aspect-oriented programming on formal methods. In ECOOP ’98 Workshop on Aspect-Oriented Programming, volume 1543 of Lecture Notes in Computer Science, page 436, July 1998. Position Paper.

[blair1998g] Gordon S. Blair, Geoff Coulson, Frank Eliassen, Thomas Plagemann, and Weihai Yu. CORBAng: Towards next generation middleware. A proposal for a collaboration research project between norway and britain in 1988, Lancaster University, University of Tromsø, University of Oslo, 1998.

[blair1998i] Gordon S. Blair and Geoff Coulson. The case for reflective middleware. Distributed Multimedia Research Group Report MPG-98-38, Distributed Multimedia Research Group, Lancaster University, 1998.

[blair1999c] Gordon S. Blair, Fabio Costa, Geoff Coulson, Hector Duran, Nikos Parlavantzas, Fabien Delpiano, Bruno Dumant, Franćois Horn, and Jean-Bernard Stefani. The design of a resource-aware reflective middleware architecture. In Proceeding of the 2nd International Conference on Meta-Level Architectures and Reflection (Reflection’99), volume 1616 of Lecture Notes in Computer Science, pages 115–134, St-Malo, France, 1999. Springer-Verlag.

[blair2000d] Gordon S. Blair. On the failure of middleware to support multimedia applications. In 7th International Workshop on Interactive Distributed Multimedia Systems and Telecommunication Services, October 2000.

[blair2000e] Lynne Blair, Trevor Jones, and Gordon S. Blair. Stochastically enhanced timed automata. In Proceedings of the Fourth International Workshop on Formal Methods for Open Object-Based Distributed System (FMOODS’99), Standford, California, 2000.

[blefari-melazzi1996a] Nicola Bléfari-Melazzi. Modeling and analysis of MPEG video sources for performance evaluation of broadband integrated networks. In Bernhard Plattner, editor, Broadband Communications — 1996 International Zurich Seminar on Digital Communications, IZS’96, volume 1044 of Lecture Notes in Computer Science, pages 131–148, Zurich, Switzerland, February 1996. Springer-Verlag.

[blum1994a] Christian Blum, Refik Molva, and Erich Rütsche. A terminal-based approach for distributed multimedia applications. Technical report, Institut Eurécom, 1994.

[blum1996a] Christian Blum and Refik Molva. A software platform for distributed multimedia applications. In Proceedings of the International Workshop on Multimedia Software Development, pages 10–21, Berlin, Germany, March 1996. IEEE Computer Society Press.

[blum1997a] Christian Blum and Refik Molva. A CORBA-based platform for distributed multimedia applications. In Proceedings of Multimedia Computing and Networking ’97, San José, CA, February 1997.

[bochmann1977a] Gregor V. Bochmann and Jan Gecsei. A unified method for the specification and verification of protocols. In Bruce Gilchrists, editor, Information Processing 77 — Proceedings of IFIP Congress 77, volume 7 of IFIP Congress Series, Toronto, August 1977. North-Holland Publishing Company.

[bochmann1996a] Gregor V. Bochmann, Brigitte Kerherve, Abdelhakim Hafid, Petre Dini, and Anne Pons. Architectural design of adaptive distributed multimedia systems. In Proceedings of the International Workshop on Multimedia Software Development, pages 31–40, Berlin, Germany, March 1996. IEEE Computer Society Press.

[bochmann1997a] Gregor V. Bochmann and Abdelhakim Hafid. Some principles for quality of service management. Distributed System Engineering Journal, 4:16–27, 1997.

[bocking1995a] Stefan Böcking. TIP’s performance quality of service. IEEE Communications Magazine, 33(8):74–81, August 1995.

[bolognesi1987a] Tommaso Bolognesi and Ed Brinksma. Introduction to the ISO specification language LOTOS. Computer Networks and ISDN Systems, (14):25–59, 1987.

[bonde1994a] Allen R. Bonde, Jr. and Sumit Ghosh. A comparative study of fuzzy versus ‘fixed’ thresholds for robust queue managment i cell-switching networks. IEEE/ACM Transactions on Networking, 2(4):337–344, August 1994.

[bosch1993a] Peter Bosch, Sape Mullender, and Tage Stabell-Kulø. Huygens file server: Storage architecture. Pegasus paper 93-3, University of Cambridge Computer Laboratory/University of Twente Faculty of Computer Science, August 1993.

[bosch1994a] Peter Bosch and Arne Helme. Secure personal computing in open environments, January 1994. Extended Abstract (draft).

[bosch1997a] Jan Bosch, Wolfgang Weck, and Clemens Szyperski. 2nd workshop on componentoriented programming (wcop’97), summary. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Component-Oriented Programming, volume 1357 of Lecture Notes in Computer Science, pages 323–326, Jyväskylä, Finland, June 1997. Springer-Verlag.

[bosch1997b] Jan Bosch. Adapting object-oriented components. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Component-Oriented Programming, volume 1357 of Lecture Notes in Computer Science, pages 379–383, Jyväskylä, Finland, June 1997. Springer-Verlag.

[botafogo1995a] Rodrigo Botafogo and Daniel Mossé. The MORENA model for hypermedia authoring and browsing. In Proceedings of the International Conference on Multimedia Computing and Systems, pages 42–49, Washington, D.C., May 1995. IEEE Computer Society Press.

[bouajjani1995a] Ahmed Bouajjani and Yassine Lakhnech. Temporal logic + timed automata: Expressiveness and decidability. In Insup Lee and Scott A. Smolka, editors, CON-CUR ’95: Concurrency Theory, 6th International Conference, volume 962 of Lecture Notes in Computer Science, pages 531–545, Philadelphia, PA, USA, August 1995. Springer-Verlag.

[boucher1995a] A. Boucher, Z. Yaar, E. J. Rubin, J. D. Palmer, and T. D. C. Little. Design and performance of a multi-stream MPEG-I system layer encoder/player. In IS&T/SPIE Symposium on Electronic Imaging Science and Technology, Multimedia Computing and Networking 1995, San Jose, February 1995.

[bowen1993a] Jonathan Bowen. Formal methods in safty-critical standards. Technical report, Oxford University Computing Laboratory, 1993.

[bowman1994a] Howard Bowman, Lynne Blair, Gordon S. Blair, and Amanda G. Chetwynd. A formal description technique supporting expressions of quality of service and media synchronization. In D. Hutchison, A. Danthine, H. Leopold, and G. Coulson, editors, Multimedia Transport and Teleservice, International COST 237 Workshop, volume 882 of Lecture Notes in Computer Science, pages 145–167, Vienna, Austria, November 1994. Springer-Verlag.

[box1999a] Don Box. Windows 2000 brings significant refinements to the COM(+) programming model. Microsoft System Journal, 14(5):17–34, May 1999.

[bracha1987a] Gabriel Bracha and Sam Toueg. Distributed deadlock detection. Distributed Computing, 2:127–139, 1987.

[brant1998a] John Brant, Brian Foote, Ralph E. Johnson, and Donald Roberts. Wrappers to the reuse. In Eric Jul, editor, ECOOP’98 – Object-Oriented Programming, volume 1445 of Lecture Notes in Computer Science, pages 396–417, Brussels, Belgium, July 1998. Springer-Verlag.

[brassard1995a] Gilles Brassard. A quantum jump in computer science. In Jan van Leeuwen, editor, Computer Science Today — Recent Trends and Developments, volume 1000 of Lecture Notes in Computer Science, pages 1–14. Springer-Verlag, 1995.

[bratbergsengen1989a] Kjell Bratbergsengen and Torgrim Gjelsvik. The development of the CROSS8 and HC16-186 parallel (database) computers. In Proceedings of International Workshop on Database Machines, IWDM’89. Springer-Verlag, 1989.

[bray1997a] Mike Bray. Middleware. Software technology review, Software Engineering Institute, Carnegie Mellon University, June 1997.

[bricker1991a] Allan Bricker, Michel Gien, Marc Guillemont, Jim Lipkis, Douglas Orr, and Marc Rozier. A new look at microkernel-based UNIX operating systems: Lessons in performance and compatibility. In Distributed Open Systems in Perspective — Proceedings of the Spring 1991 EurOpen Conference, pages 13–32, Tromsø, Norway, May 1991. EurOpen.

[briggs1994a] Ted L. Briggs and John Werth. A specification language for object-oriented analysis and design. In Mario Tokoro and Remo Pareschi, editors, Proceedings of the Eighth European Conference on Object-Oriented Programming, volume 821 of Lecture Notes in Computer Science, pages 365–385, Bologna, Italy, July 1994. Springer-Verlag.

[brinch-hansen1970a] Per Brinch Hansen. The nucelus of a multiprogramming system. Communications of the ACM, 13(4):238–250, April 1970.

[brinch-hansen1972a] Per Brinch Hansen. Structured multiprogramming. Communications of the ACM, 15(7):574–578, July 1972.

[brinch-hansen1978a] Per Brinch Hansen. Distributed processes: A concurrent programming concept. Communications of the ACM, 21(11):934–941, November 1978.

[brinksma1991a] Ed Brinksma. What is the method in formal methods? Technical Report TIOS 91-29, MemorandaInformatica 91-79, Tele-Informatics Group, Department of Computer Science, University of Twente, November 1991.

[brinksma1991b] Ed Brinksma and Pim Kars. From data structure to porcess structure. Technical Report TIOS 91-11, MemorandaInformatica 91-38, Department of Computer Science, University of Twente, 1991.

[brookes1994a] Wayne Brookes and Jadwiga Indulska. A type management system for open distributed processing. Technical report, CRC for Distributed Systems Technology, Department of Computer Science, The University of Queensland, 1994.

[brookes1994b] Wayne Brookes and Jadwiga Indulska. A formal specification of a basic ODPbased type manager. Technical report, CRC for Distributed Systems Technology, Department of Computer Science, The University of Queensland, 1994.

[brose1998a] Gerald Brose. Reflection in CORBA, Java and JacORB. In Proceedings JIT. Springer-Verlag, November 1998.

[browne1993a] Shirley Browne. Communication and synchronization issues in distributed multimedia database systems. In Nabil R. Adam and Bharat K. Bhargava, editors, Advanced Database Systems, volume 759 of Lecture Notes in Computer Science, pages 381–396. Springer-Verlag, 1993.

[broy1982a] Manfred Broy and Martin Wirsing. Partial abstract types. Acta Informatica, 18:47– 64, 1982.

[broy1992a] Manfred Broy. Experiences with software specification and verification using LP, the larch proof assistant. SRC Research Reports 93, DEC’s System Research Center, November 1992.

[bruce1997a] Kim B. Bruce, Leaf Peteresen, and Adrian Fiech. Subtyping is not a good “match” for object-oriented languages. In Mehmet Aksit and Satoshi Matsuoka, editors, ECOOP’97 — The 11th European Conference on Object-Oriented Programming, volume 1241 of Lecture Notes in Computer Science, pages 104–127, Jyväskylä, Finland, June 1997. Springer-Verlag.

[bryans1995a] Jeremy Bryans, Jim Davies, and Steve Schneider. Towards a denotational semantic for ET-LOTOS. In Insup Lee and Scott A. Smolka, editors, CONCUR ’95: Concurrency Theory, 6th International Conference, volume 962 of Lecture Notes in Computer Science, pages 269–283, Philadelphia, PA, USA, August 1995. Springer-Verlag.

[bryhni1996a] Haakon Bryhni, Hilde Lovett, Erling Maartmann-Moe, Dag Solvoll, and Tryggve Sørensen. On-demand regional television over the internet. In Proceedings Multimedia’96, New York, NY, USA, November 1996. ACM Press.

[bucci1994a] Giacomo Bucci, Riccardo Detti, Valdo Pasqui, and Stefano Nativi. Sharing multimedia data over a client-server network. IEEE Multimedia, 1(3):44–55, Fall 1994.

[buchi1997a] Martin Büchi and Emil Sekerinski. Formal methods for component software: The refinement calculus perspective. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Component-Oriented Programming, volume 1357 of Lecture Notes in Computer Science, pages 332–337, Jyväskylä, Finland, June 1997. Springer-Verlag.

[bulterman1992a] Dick C. A. Bulterman. Synchronization of multi-sourced multimedia data for heterogeneous target systems. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 119–129, La Jolla, California, USA, November 1992. Springer-Verlag.

[bulterman1993a] Dick C. A. Bulterman. Specification and support of adaptable networked multimedia. Multimedia Systems, 1(2):68–76, 1993.

[bulterman1995a] Dick C. A. Bulterman and Lynda Hardman. Multimedia authoring tools: State of the art and research challenges. In Jan van Leeuwen, editor, Computer Science Today — Recent Trends and Developments, volume 1000 of Lecture Notes in Computer Science, pages 575–591. Springer-Verlag, 1995.

[bunker1994a] Ted Bunker. Is it 1984? LAN Magazine, 9(8), August 1994.

[burkow1994a] Tatjana M. Burkow. Operating system support for distributed multimedia applications; a survey of current research. Pegasus paper 94-8, University of Cambridge Computer Laboratory/University of Twente Faculty of Computer Science, June 1994.

[burkow1995a] Tatjana M. Burkow and Sape J. Mullender. Multimedia systems. In Advanced Course: Multimedia Databases in Perspective, pages 73–128, June 1995.

[burrows1989a] Michael Burrows, Martín Abadi, and Roger Needham. A logic of authentication. SRC Research Reports 39, DEC’s System Research Center, February 1989.

[calvin1995a] Licia Calvin and Walter Geerts. ITEM: A multimedia application in the modern language curriculum. In F. Williams, H. Brä, and J. Nolan, editors, Proceedings of the Fourth International Conference Broadband Island ’95, Global Broadband & Beyond, pages 471–473, Dublin, Ireland, September 1995.

[campbell1992a] Andrew Campbell, Geoff Coulson, Francisco García, and David Hutchison. Orchestration services for distributed multimedia synchronisation. In A. Danthine and O. Spaniol, editors, High Performance Networking, IV, Proceedings of the IFIP TC6/WG6.4 Fourth International Conference on High Performance Networking, IFIP Transactions, volume C-14, pages 153–168, Liège, Belgium, December 1992. Elsevier Science B. V.

[campbell1993a] Andrew Campbell, Geoff Coulson, Francisco García, David Hutchison, and Helmut Leopold. Integrated quality of service for multimedia communications. Lancaster University Report MPG-93-17, Department of Computing, Lancaster University/Alcatel ELIN Research, 1994.

[campbell1993b] Andrew Campbell, Geoff Coulson, and David Hutchison. A multimedia enhanced transport service in a quality of service architecture. In D. Shepherd, G. Blair, G. Coulson, N. Davies, and F. Garcia, editors, Proceedings of the 4th International Workshop on Network and Operating System Support for Digital Audio and Video, volume 846 of Lecture Notes in Computer Science, pages 124–137, Lancaster, U.K., November 1993. Springer-Verlag.

[campbell1994a] Andrew Campbell, Geoff Coulson, and David Hutchison. A quality of service architecture. Lancaster University Report MPG-94-08, Department of Computing, Lancaster University, 1994.

[campbell1994b] Andrew Campbell, Geoff Coulson, and David Hutchison. Flow managment in a quality of service architecture. In S. Fdida, editor, High Performance Networking, V — Proceedings of the IFIP TC6/WG6.4 Fifth International Conference on High Performance Networking, volume C-26 of IFIP Transactions, pages 201–218, Grenoble, France, 1994. North-Holland Publishing Company.

[campbell1995a] Andrew Campbell, Geoff Coulson, and David Hutchison. Supporting adaptive flows in a quality of service architecture. Lancaster University Report MPG-95-19, Department of Computing, Lancaster University, 1995.

[campbell1995b] Andrew Campbell, Cristina Aurrecoechea, Linda Hauw, and David Hutchison. BT URI project; WP3/D1: A review of quality of service architectures. Lancaster University Report MPG-95-20, Department of Computing, Lancaster University, 1995.

[campbell1995c] Andrew Campbell, Cristina Aurrecoechea, and Linda Hauw. Architectural perspectives on QoS management in distributed multimedia systems. Lancaster University Report MPG-95-23, Center for Telecommunication Research, Columbia University, 1995.

[carreira1995a] M. Carreira, J. Casebolt, G. Desrosiers, and T. D. C. Little. Capture-time indexing paradigm, authoring tool, and browsing environment for digital broadcast video. In IS&T/SPIE Symposium on Electronic Imaging Science and Technology, Multimedia Computing and Networking 1995, San Jose, February 1995.

[cavanaugh1992a] John David Cavanaugh and Timothy J. Salo. Internetworking with ATM WANs. Technical report, Minnesota Supercomputing Center, Inc., December 1992.

[cazzola1998a] Walter Cazzola. Evaluation of object-oriented reflective models. In ECOOP’98 Workshop on Reflective Object-Oriented Programming Systems, Brussels, Belgium, July 1998.

[cazzola2000a] Walter Cazzola and Massimo Ancona. mChaRM: a reflective middleware for communication-based reflection. Report DISI-TR-00-09, DISI, Università degli Studi di Genova, May 2000.

[cen1995a] Shanwei Cen, Calton Pu, Richard Staehli, Crispin Cowan, and Jonathan Walpole. A distributed real-time MPEG video audio player. In NOSSDAV ’95, 1995.

[chambers1999a] Craig Chambers and Weimin Chen. Efficient multiple and predicate dispatching. In Proceedings of OOPSLA’99, November 1999.

[champine1990a] George A. Champine, Daniel E. Geer, and William N. Ruh. Project Athena as a distributed computer system. Computer, pages 40–50, September 1990.

[chandy1985a] K. Mani Chandy and Leslie Lamport. Distributed snapshots: Determining global states of distributed systems. ACM Transactions on Computer Systems, 3(1):63–75, February 1985.

[chandy1998a] K. Mani Chandy, Adam Rifkin, and the Infospheres Group. Caltech Infospheres project. In Workshop on Compositional Software Architectures, Monterey, California, January 1998. (position paper).

[chang1984a] Jo-Mei Chang and N. F. Maxemchuk. Reliable broadcast protocols. ACM Transactions on Computer Systems, 2(3):251–273, August 1984.

[chang1995a] Shih-Fu Chang and David G. Messerschmitt. Manipulation and compositing of MC-DCT compressed video. IEEE Journal on Selected Areas in Communications, 13(1):1–11, January 1995.

[chappell1998a] David Chappell. The trouble with CORBA. Object News, April 1998.

[chassot1995a] Christophe Chassot, Michel Fournier, Michel Diaz, and André Lozes. Service definition of a multimeda partial order connection. In D. Hutchison, H. Christiansen, G. Coulson, and A. Danthine, editors, Teleservices and Multimedia Communications — Second International COST 237 Workshop, volume 1052 of Lecture Notes in Computer Science, pages 118–137, Copenhagen, Denmark, November 1995. Springer-Verlag.

[chatterjee1997a] Saurav Chatterjee. A quality of service based allocation and routing algorithm for distributed, hetrogenous real time systems. In Proceedings of the 17th International Conference on Distributed Systems, Baltimore, Maryland, USA, May 1997. IEEE Computer Society Press.

[chen1993a] H.-J. Chen and T. D. C. Little. Physical storage organizations for time-dependent multimedia data. In Proceedings of the 4th International Conference on Foundations of Data Organization and Algorithms, FODO’93, Evanston, IL, October 1993.

[chen1995a] Zhigang Chen, See-Mong Tan, Roy H. Campbell, and Yongcheng Li. Real time video and audio in the world wide web. In WWW’95, 1995. (submitted).

[chen1995b] H.-J. Chen, A. Krishnamurthy, T. D. C. Little, and D. Venkatesh. A scalable videoon-demand service for the provision of VCR-like functions. In Proceedings of the 2nd International Conference on Multimedia Comuting Systems, Washington DC, May 1995.

[chen1995c] H.-J. Chen, T. D. C. Little, and D. Venkatesh. A storage and retrieval technique for the provision of scalable MPEG video. Journal of Parallel and Distributed Computing, 30(2):180–189, November 1995.

[chen1998a] Shigang Chen and Klara Nahrstedt. An overview of quality-of-service routing for the next generation high-speed networks: Problems and solutions. IEEE Network Magazine, 1998.

[cheriton1983a] David R. Cheriton and Willy Zwaenepoel. The distributed V kernel and its performance for diskless workstations. Operating Systems Review, 17(5):129–140, October 1983. (Proceedings of the 9th Symposium on Operating System Principles, October 1983).

[cheriton1988a] David R. Cheriton. The V distributed system. Communications of the ACM, 31(3):314–333, March 1988.

[cheriton1993a] David R. Cheriton and Dale Skeen. Understanding the limitations of causally and totally ordered communication. In Proceedings of the 14th Symposium on Operating System Principles, Asheville, North Carolina, December 1993. ACM.

[chesson1989a] Greg Chesson. XTP/PE design considerations. In IFIP WG6.1/6.4 Workshop on Protocols for High-Speed Networks, May 1989.

[chesson1991a] Greg Chesson. The evolution of XTP. In Proceedings of the Third International Conference on High Speed Networking, North-Holland, 1991.

[chiariglione1995a] Leonardo Chiariglione. MPEG: A technological basis for multimedia applications. IEEE Multimedia, 2(1):85–89, Spring 1995.

[chiba1993a] Shigeru Chiba and Takashi Masuda. Designing an extensible distributed language with a meta-level architecture. In Oscar M. Nierstrasz, editor, ECOOP ’93 — Object Oriented Programming, volume 707 of Lecture Notes in Computer Science, pages 482–501, Kaiserslautern, Germany, July 1993. Springer-Verlag.

[chiba1997a] Shigeru Chiba, Takashi Nishimura, Kenichi Kourai, Atsushi Ohnoki, and Takashi Masuda. Weak protection for reflective operating systems. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Reflective Real-Time Object-Oriented Programming and Systems, volume 1357 of Lecture Notes in Computer Science, pages 62–67, Jyväskylä, Finland, June 1997. Springer-Verlag.

[chiba1998a] Shigeru Chiba and Michiaki Tatsubori. A yet another java.lang.class. In ECOOP’98 Workshop on Reflective Object-Oriented Programming Systems, Brussels, Belgium, July 1998.

[chiueh1999a] Tzi cker Chiueh, Chitra Venkatramani, and Michael Vernick. The intergartion of real-time i/o and network support in the Stony Brook video server. IEEE Network, 13(2):30–36, March/April 1999.

[cho1995a] H. S. Cho, M. R. Fry, A. Senevirante, and V. Witana. Towards a hybrid scheme for application adaptivity. In D. Hutchison, H. Christiansen, G. Coulson, and A. Danthine, editors, Teleservices and Multimedia Communications — Second International COST 237 Workshop, volume 1052 of Lecture Notes in Computer Science, pages 177–191, Copenhagen, Denmark, November 1995. Springer-Verlag.

[cho1998a] Kenjiro Cho. A framework for alternate queueing: Towards traffic management by PC-UNIX based routers. In Proceedings of USENIX 1998 Annual Technical Conference, New Orleans, LA, June 1998.

[cho1999a] Sehyeong Cho and Youngmee Shin. Multimedia service interworking over hetrogeneous networking environments. IEEE Network, 13(2):61–69, March/April 1999.

[cho1999b] Kenjiro Cho. Managing traffic with ALTQ. In Proceedings of USENIX 1999 Annual Technical Conference: FREENIX Track, Monterey CA, June 1999.

[chockler1996a] Gregory V. Chockler, Nabil Huleihel, Idit Keidar, and Danny Dolev. Multimedia multicast transport service for groupware. In Proceedings of the TINA Conference on the Convergence of Telecommunications and Distributed Computing Technologies, pages 43–54, September 1996.

[chou1992a] Stephen T.-C. Chou and Hideyuki Tokuda. System support for dynamic QOS control of continuos media communication. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 364–368, La Jolla, California, USA, November 1992. Springer-Verlag.

[choukair1997a] Zièd Choukair and Antoine Beugnard. Real-time object-oriented distributed processing with COREMO. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Object-Oriented Technology for Telecommunications Services Engineering, volume 1357 of Lecture Notes in Computer Science, pages 45–50, Jyväskylä, Finland, June 1997. Springer-Verlag.

[christopher1993a] Wayne A. Christopher, Steven J. Procter, and Thomas E. Anderson. The Nachos instructional operating system. Technical report, Computer Science Division, University of California, 1993.

[chu1997a] Hao hua Chu and Klara Nahrstedt. Soft real time scheduling server in UNIX operating system. In Ralf Steinmetz and Lars C. Wolf, editors, IDMS’97 — 4th International Workshop on Interactive Distributed Multimedia Systems and Telecommunication Services, volume 1309 of Lecture Notes in Computer Science, Darmstadtm Germany, September 1997. Springer-Verlag.

[clarke1998a] Michael Clarke and Geoff Coulson. An architecture for dynamically extensible operating systems. In Proceedings of the 4th International Conference on Configurable Distributed Systems (ICCDS ’98), Annapolis, Maryland, US, May 1998.

[clavel1996a] Manuel G. Clavel and José Meseguer. Axiomatizing reflective logics and languages. In Gregor Kiczales, editor, Proceedings of the Reflection’96 Conference, pages 263– 288, April 1996.

[cointe1996a] Pierre Cointe. Reflective languages and metalevel architectures. ACM Computing Surveys, 28A(4), December 1996.

[contreras1998a] José L. Contreras and Jean-Louis Sourrouille. Adaptive active object. In ECOOP’98 Workshop on Reflective Object-Oriented Programming Systems, Brussels, Belgium, July 1998.

[corbato1965a] F. J. Corbató and V. A. Vyssotsky. Introduction and overview of the multics system. In Proceedings of AFIPS Fall Joint Computer Conference, pages 185–196, 1965.

[costa-carmo1992a] Luiz F. Rust da Costa Carmo, Pierre de Saqui-Sannes, and Jean-Pierre Courtiat. Basis synchronisation concepts in multimeda systems. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 94–105, La Jolla, California, USA, November 1992. Springer-Verlag.

[costa1998a] Fabio Costa, Gordon S. Blair, and Geoff Coulson. Experiments with reflective middleware. In ECOOP’98 Workshop on Reflective Object-Oriented Programming Systems, volume 1543 of Lecture Notes in Computer Science, pages 390–391, Brussels, Belgium, July 1998.

[costa1998b] Fabio Costa, Katia B. Saikoski, Tom Kristensen, Hans-Ole Rafaelsen, Knut-Olav Tra, Frank Eliassen, and Thomas Plageman. Binding framework. Workshop report, CORBAng project, December 1998.

[costa1999a] Fabio Costa, Hector Duran, Nikos Parlavantzas, Katia B. Saikoski, Gordon S. Blair, and Geoff Coulson. The role of reflective middleware in supporting the engineering of complex applications. In Walter Cazzola, Robert J. Stroud, and Francesco Tisato, editors, OOPSLA’99 Workshop on Object Oriented Reflection and Software Engineering, pages 63–76, Denver, Colorado, US, November 1999.

[costa2000a] Fabio Costa and Gordon S. Blair. The role of meta-information management in reflective middleware. In ECOOP’2000 Workshop on Reflection and Metalevel Architectures, Sophia Antipolis and Cannes, France, June 2000.

[coulson1992a] G. Coulson, G. S. Blair, N. Davies, and N. Williams. Extensions to ANSA for multimedia computing. Computer Networks and ISDN Systems, (25):305–323, 1992.

[coulson1993a] Geoff Coulson, Gordon Blair, and Philippe Robin. Micro-kernel support for continuous media in distributed systems. Lancaster University Report MPG-93-04, Department of Computing, Lancaster University, 1993.

[coulson1993b] Geoff Coulson, Gordon Blair, Philippe Robin, and Doug Shepherd. Extending the Chorus micro-kernel to support continuous media applications. In D. Shepherd, G. Blair, G. Coulson, N. Davies, and F. Garcia, editors, Proceedings of the 4th International Workshop on Network and Operating System Support for Digital Audio and Video, volume 846 of Lecture Notes in Computer Science, pages 54–65, Lancaster, U.K., November 1993. Springer-Verlag.

[coulson1994a] Geoff Coulson, Andrew Campbell, Philippe Robin, Gordon Blair, Michael Papathomas, and David Hutchison. The design of a QoS controlled ATM based communications system i Chorus. Lancaster University Report MPG-94-05, Distributed Multimedia Research Group, Department of Computing, Lancaster University, 1994.

[coulson1994b] G. Coulson, G. S. Blair, P. Robin, and D. Shepherd. Supporting continuous media applications in a micro-kernel environment. Lancaster University Report MPG-94-16, Department of Computing, Lancaster University, 1994.

[coulson1994c] Geoff Coulson and Gordon Blair. Meeting the real-time synchronisation requirements of multimedia in open distributed processing. Distributed System Engineering Journal, 1994.

[coulson1995a] G. Coulson, G. S. Blair, J. B. Stefani, F. Horn, and L. Hazard. Supporting the real-time requirements of continuos media in open distributed processing. Computer Networks and ISDN Systems, 27(8):1231–1246, July 1995.

[coulson1995b] Geoff Coulson and Gordon Blair. Architectural principles and techniques for distributed multimedia applications support in operating systems. Operating Systems Review, 29(4):17–24, October 1995.

[coulson1995c] Geoff Coulson, Andrew Campbell, David Hutchison, and Frankie Garcia. BT URI project; WP2/D2: Session acceptance and QoS management in end-systems. Lancaster University Report MPG-95-21, Department of Computing, Lancaster University, 1995.

[coulson1998a] Geoff Coulson. Configurable performance and predictability in a soft realtime/multimedia middleware platform. Distributed multimedia research group internal report, Computing Department, Lancaster University, 1998.

[coulson1998b] G. Coulson. A distributed object platform infrastructure for multimedia applications. Computer Communications, 1998. (to appear).

[coulson1998c] G. Coulson. A configurable multimedia middleware platform. Distributed multimedia research group internal report, Computing Department, Lancaster University, 1998.

[courtait1991a] JP. Courtiat and M. Diaz. Time in state based formal description techniques for distributed systems. In J. W. de Bakker, C. Huizing, W. P. de Roever, and G. Rozenberg, editors, Proceedings of the REX Workshop 1991, Real-Time: Theory in Practice, volume 600 of Lecture Notes in Computer Science, pages 149–175, Mook, The Netherlands, June 1991. Springer-Verlag.

[cowan1995a] Crispin Cowan, Shanwei Cen, Jonathan Walpole, and Calton Pu. Adaptive methods for distributed video presentation. ACM Computing Surveys, 27(4):580–583, December 1995.

[craigen1993a] Dan Craigen, Susan Gehart, and Ted Ralston. An international survey of industrial applications of formal methods — volume 1 purpose, approach, analysis, and conclusions. NISTGCR 93/626, U.S. Department of Commerece, Technology Administration, National Institute of Standards and Technology, Computer Systems Laboratory, March 1993.

[craigen1993b] Dan Craigen, Susan Gehart, and Ted Ralston. An international survey of industrial applications of formal methods — volume 2 case studies. NISTGCR 93/626, U.S. Department of Commerece, Technology Administration, National Institute of Standards and Technology, Computer Systems Laboratory, March 1993.

[cramer1992a] Andreas Cramer, Manny Farber, Brian McKellar, and Ralf Steinmetz. Experiences with the Heidelberg multimedia communication system. In A. Danthine and O. Spaniol, editors, High Performance Networking, IV, Proceedings of the IFIP TC6/WG6.4 Fourth International Conference on High Performance Networking, IFIP Transactions, volume C-14, pages 199–218, Liège, Belgium, December 1992. Elsevier Science B. V.

[cranor199?a] Charles D. Cranor and Gurudatte M. Parulkar. Design of universal continuos media i/o. Technical report, Washington University, 199?

[crawley1997a] S. Crawley, S. Davis, J. Indulska, S. McBride, and K. Raymond. Meta information management. In Howard Bowman and John Derrick, editors, IFIP TC6 WG6.1 International Workshop on Formal Methods for Open Object-based Distributed Systems (FMOODS ’97), volume 2, pages 194–202, Canterbury, Kent, UK, July 1997. International Federation for Information Processing (IFIP), Chapman & Hall.

[cregut1994a] Pierre Crégut and David B. MacQueen. An implementation og higher-order functors. Technical report, CNET/France Telecom, AT&T Bell Laboratories, 1994.

[crowcroft1993a] Jon Crowcroft. Remote procedure call: Not a panacea for distributed computing problems. Draft, February 1993.

[cunningham1994a] H. Conrad Cunningham. Notes on functional programming with Gofer. Technical Report UMCIS-1994-01, Software Methods Research Group, Department of Computer and Information Science, University of Missisippi, January 1994.

[dale1990a] Kåre Dale. Introduksjon til ANSA testbenk. Technical report, Seksjon for Datafag, Institutt for Matematiske Realfag, Universitetet i Tromsø, 1990.

[dalen1995a] Terje Dalen and Gisle Aas. Kommunikasjonsprotokoll for leveranse av videodata i ATM. Rapport IT291/1-95, Norut IT, January 1995.

[dalseng19??a] John I. Dalseng. Algebraisk spesifikasjon, funksjonell programmering og program verifikasjon. Seksjon for Datafag, IMR, Universitetet i Tromsø.

[damaskos1994a] Spiridon Damaskos and Anastassios Gavras. A simplified QoS model for multimedia protocols over ATM. In S. Fdida, editor, High Performance Networking, V — Proceedings of the IFIP TC6/WG6.4 Fifth International Conference on High Performance Networking, volume C-26 of IFIP Transactions, pages 241–257, Grenoble, France, 1994. North-Holland Publishing Company.

[danielsen1990a] Thore Danielsen, Wiggo Finnset, Frode Flægstad, Gunnar Hartvigsen, and Roar Steen. Ratatosk — et adaptivt brukergrensesnitt for datamaskinbasert kommunikasjon. TF-rapport R35/90, Teledirektoratets forskningsavdeling, 1990.

[danthine1992a] Andrè Danthine, Yves Baguete, Guy Leduc, and Luc Lèonard. The OSI 95 connection-mode transport service — the enhanced QoS. In A. Danthine and O. Spaniol, editors, High Performance Networking, IV, Proceedings of the IFIP TC6/WG6.4 Fourth International Conference on High Performance Networking, IFIP Transactions, volume C-14, pages 235–252, Liège, Belgium, December 1992. Elsevier Science B. V.

[danthine1993a] Andrè Danthine, Yves Baguete, Luc Lèonard, and Guy Leduc. An enhancement of the QoS concept. Missing section 6 on QoS of the ECFF guidelines document ISO/IEC JTC1/SC6/WG4/N, Institut d’Electricité Montefiore B28, Université de Liège, Belgium, January 1993.

[danzig1991a] Peter B. Danzig and Sugih Jamin. tcplib: A library of TCP internetwork traffic characteristics. Technical Report USC-CS-91-495, Computer Science Department, University of Southern California, 1991.

[das1997a] Sajal K. Das, Rajeev Jayaram, and Sanjoy K. Sen. An optimistic quality-of-service provisioning scheme for cellular networks. In Proceedings of the 17th International Conference on Distributed Systems, Baltimore, Maryland, USA, May 1997. IEEE Computer Society Press.

[dasgupta1988a] Partha Dasgupta, Richard J. LeBlanc, Jr., and William F. Appelbe. The Clouds distributed operating system. In The 8th International Conference on Distributed Computing Systems. IEEE, Computer Society Press, 1988.

[day1977a] Robert A. Day. How to write a scientific paper. IEEE Transactions on Professional Communication, PC-20(1):32–37, June 1977.

[day1995a] Mark Day, Robert Gruber, Barbara Liskov, and Andrew C. Myers. Subtypes vs. where clauses: Constraining parametric polymorphism. In OOPSLA 1995, pages 156–168, 1995.

[decina1997a] Maurizio Decina and Vittorio Trecordi. Convergence of telecommunications and computing to networking models for integrated services and applications. Proceedings of the IEEE, 85(12), December 1997.

[degener1994a] Jutta Degener. Digital speech compression. Dr. Dobb’s Journal, pages 30–34, December 1994.

[deicke1997a] J. Deicke, U. Mayer, and M. Glesner. An object-oriented client/server architecture for video-on-demand applications. In Ralf Steinmetz and Lars C. Wolf, editors, IDMS’97 — 4th International Workshop on Interactive Distributed Multimedia Systems and Telecommunication Services, volume 1309 of Lecture Notes in Computer Science, Darmstadtm Germany, September 1997. Springer-Verlag.

[delacroix1995a] Joelle Delacroix and Christian Santellani. A test bed for distributed real-time scheduling experimentation based on the CHORUS micro-kernel. In Proceedings of European Research Seminar on Advances in Distributed Systems, ERSADS’95, L’Alpe-d’-Huez, April 1995.

[delgrossi1993a] Luca Delgrossi, Ralf Guido Herrtwich, Carsten Vogt, and Lars C. Wolf. Reservation protocols for internetworks: A comparison of ST-II and RSVP. In Proceedings of the Fourth International Workshop on Network and Operating System Support for Digital Audio and Video, Lancaster, UK, November 1993.

[dellafera19??a] C. Anthony DellaFera, Mark W. Eichin, Robert S. French, David C. Jedlinsky, John T. Kohl, and William E. Sommerfeld. The Zephyr notification service. Project Athena.

[dermler1992a] G. Dermler and K. Froitzheim. JVTOS — a reference model for a new multimedia service. In A. Danthine and O. Spaniol, editors, High Performance Networking, IV, Proceedings of the IFIP TC6/WG6.4 Fourth International Conference on High Performance Networking, IFIP Transactions, volume C-14, pages 183–197, Liège, Belgium, December 1992. Elsevier Science B. V.

[dermler1995a] Gabriel Dermler, Walter Fiederer, Ingo Barth, and Kurt Rothermel. A framework for negotiable quality of service in distributed multimedia systems. Fakultätsbericht 10/1995, Fakultät Informatik, Universität Stuttgart, December 1995.

[dermler1996a] Gabriel Dermler, Walter Fiederer, Ingo Barth, and Kurt Rothermel. A negotiation and resource reservation protocol (NRP) for configurable multimedia applications. In Proceedings of Multimedia ’96, pages 113–116, 1996.

[dijkstra1968a] E. W. Dijkstra. Co-operating sequential processes. In F. Genuys, editor, Programing Languages, pages 43–112. Academic Press, 1968.

[dijkstra1968b] Edsger W. Dijkstra. The structure of the “THE”-multiprogramming system. Communications of the ACM, 11(5), May 1968.

[dijkstra1972a] E. W. Dijkstra. Hierarchical ordering of sequential processes. In C. A. R. Hoare and R. H. Perrott, editors, Operating Systems Techniques, Proceeding of a Seminar Held at Queen’s University, pages 72–93, Belfast, 1972. Academic Press.

[dijkstra1975a] Edsger W. Dijkstra. Guarded commands, nondeterminancy and formal derivation of programs. Communications of the ACM, 18(8):453–457, August 1975.

[dijkstra1982b] Edsger W. Dijkstra. How do we tell truths that might hurt? In Selected Writings on Computing: A personal Perspective. Springer-Verlag, 1982.

[dijkstra1986a] Edsger W. Dijkstra. A belated proof of self-stabilization. Distributed Computing, 1:5–6, 1986.

[dillema1994a] F. W. Dillema. Intelligent user interfaces — why so many computer scientists have fear of flying. Technical report, University of Twente, August 1994.

[dodani1989a] Mahesh H. Dodani, Charles E. Hughes, and J. Michael Moshell. Separation of powers. Byte, pages 255–262, March 1989.

[dolev1994a] Danny Dolev and Dalia Malki. The design of the Transis system. In K. P. Birman, F. Mattern, and A. Schiper, editors, Theory and Practice in Distributed Systems, International Workshop, Selected Papers, volume 938 of Lecture Notes in Computer Science, pages 83–98, Dagstuhl Castle, Germany, September 1994. Springer-Verlag.

[dolgicer1998a] Max Dolgicer. History repeats itself in middleware debates. Technical report, International System Group, Inc., October 1998.

[dong1993a] Jin Song Dong and Roger Duke. An object-oriented approach to the formal specification of ODP trader. In J. De Meer, B. Mahr, and S. Storp, editors, Open Distributed Processing II, Proceedings of the IFIP TC6/WG6.1 International Conference on Distributed Processing, IFIP Transactions, volume C-20, pages 341–352, Berlin, Germany, September 1993. Elsevier Science B. V.

[doorn1995a] L. van Doorn, P. Homburg, and A. S. Tanenbaum. Paramecium: An extensible object-based kernel. In Proceedings Hot Topics in Operating Systems V, pages 86–89. ACM, 1995.

[doorn1997a] Leendert van Doorn and Andrew S. Tanenbaum. FlexRTS: An extensible Orca run-time system. In Proceeding of the Third ASCI Conference, May 1997.

[dourish1996a] Paul Dourish, Annette Adler, and Brian Cantwell Smith. Organising user interfaces around reflective accounts. In Gregor Kiczales, editor, Proceedings of the Reflection’96 Conference, pages 235–243, April 1996.

[drayton1991a] L. Drayton, A. Chetwynd, and G. S. Blair. An introduction to LOTOS through a worked example. Distributed Multimedia Research Group Report MPG-91-18, Computing Department, Lancaster University, September 1991.

[dubois1996a] Paul F. DuBois, Konrad Hinsen, and James Hugunin. Numerical Python. Technical report, Lawrence Livermore National Laboratory and Department of Chemistry, University of Montreal and Laboratory of Computer Science, Massachusetts Institute of Technology, 1996.

[duran2000a] Hector Duran and Gordon S. Blair. A resource management framework for adaptive middleware. In 3th IEEE International Symposium on Object-oriented Real-time Distributed Computing, Newport Beach, California, USA, March 2000.

[duran2000b] Hector Duran and Gordon S. Blair. Configuring and reconfiguring resources in middleware. In 1st International Symposium on Advanced Distributed Systems, Guadalajara, Jalisco, Mexico, March 2000.

[dyer19??a] Stephen P. Dyer. The Hesoid name server. Project Athena.

[eckerson1995a] Wayne W. Eckerson. Three-tier client/server architecture. Open Information Systems, 10(1):3–22, January 1995.

[effelsberg1995a] Wolfgang Effelsberg. Communication support for distributed multimedia systems. In Advanced Course: Multimedia Databases in Perspective, pages 129–154, June 1995.

[ehley1995a] Lynnae Ehley, Mohammad Ilyas, and Boko Furth. A survey of multimedia synchronization techniques. In Borko Furth and Milan Milenkovic, editors, A Guided Tour of Multimedia Systems, pages 230–255. IEEE Computer Society Press, 1995.

[ehrich1990a] H.-D. Ehrich, J. A. Goguen, and A. Sernadas. A categorical theory of objects as observed processes. In J. W. de Bakker, W. P. de Roever, and G. Rozenberg, editors, The Proceedings of REX School/Workshop Foundations of Object-Oriented Languages, volume 489 of Lecture Notes in Computer Science, pages 203–228, Noordwijkerhout, The Netherlands, May/June 1990. Springer-Verlag.

[eichin1989a] Mark W. Eichin and Jon A. Rochlis. With microscope and tweezers: An analysis of the internet virus of november 1988. Technical report, Massachusetts Institute of Technology, February 1989. (A version presented at the 1989 IEEE Symposium on Research in Security and Privacy).

[eliassen1994a] Frank Eliassen, Randi Karlsen, and Weihai Yu. Distribuerte integrerte multimedia informasjonssystemer (DIME). Prosjektforslag, Seksjon for Informatikk, Universitetet i Tromsø, 1994.

[eliassen1995a] Frank Eliassen and John R. Nicol. Polymorphic typing of continuous media flows and its application to QoS brokerage. Technical Report TR-0303-07-95-380, GTE Laboratories Incorporated, July 1995.

[eliassen1996a] Frank Eliassen and John R. Nicol. Supporting interoperation of continuos media objects. Theory and Practice of Object Systems, 1996. (to appear).

[eliassen1996b] Frank Eliassen and John R. Nicol. A flexible type checking model for stream interface binding. In Proceedings of the International Workshop on Multimedia Software Development, pages 52–60, Berlin, Germany, March 1996. IEEE Computer Society Press.

[eliassen1997a] F. Eliassen. A conformance relationship for stream interfaces. In Howard Bowman and John Derrick, editors, IFIP TC6 WG6.1 International Workshop on Formal Methods for Open Object-based Distributed Systems (FMOODS ’97), volume 2, pages 399–408, Canterbury, Kent, UK, July 1997. International Federation for Information Processing (IFIP), Chapman & Hall.

[eliassen1999a] F. Eliassen and H. O. Rafaelsen. A conformance relationship supporting selection of explicit stream bindings. Technical report, Department of Informatics, University of Oslo and Department of Computer Science, Univeristy of Tromsø, 1999. Submitted IDMS’99.

[eliassen2000a] Frank Eliassen, Tom Kristensen, Thomas Plageman, and Hans Ole Rafaelsen. MULTE-ORB: Adaptive QoS aware binding. In Workshop on Reflective Middleware at Middleware 2000, New York, USA, April 2000.

[elsaadany1994a] A. Elsaadany, M. Singhal, and M. T. Liu. Multimedia on local area network. Technical Report OSU-CISRC-3/94-TR11, Department of Computer and Information Science, The Ohio State University, 1994.

[elstad1994b] Kristian Elstad. Bilde komprimering ved bruk av Wavelet transform koding. Rapport basert på videre arbeid med diplomoppgaven, Seksjon for Informatikk, Institutt for Matematiske Realfag, Universitetet i Tromsø, August 1994.

[elvang-goransson1991a] Morten Elvang-Gøransson, Herman Ruge Jervell, Eric Monteiro, and Arild Waaler. Propositional linear logic: Wiring semantics. Technical report, University of Oslo, May 1991. (Submitted to Information Processing Letters).

[elvang-goransson1991b] Morten Elvang-Gøransson, Eric Monteiro, and Arild Waaler. Automated proof search in propositional linear logic. Technical report, University of Oslo, May 1991. (Extended Abstract for CSL’91).

[engberg1990a] Uffe Engberg and Glynn Winskel. Petri nets as models of linear logic. In Proceedings of CAAP’90, volume 431 of Lecture Notes in Computer Science. Springer-Verlag, 1990.

[engler1995a] Dawson R. Engler and M. Frans Kaashoek. Exterminate all operating system abstractions. In HotOS V, 1995.

[engler1995b] Dawson R. Engler, M. Frans Kaashoek, and James O’Toole, Jr. Exokernel: An operating system architecture for application-level resource management. In Proceedings of the Fifteenth Symposium on Operating Systems Principles, December 1995.

[eun1994a] Seongbae Eun, Eun Suk No, Hyung Chul Kim, Hyunsoo Yoon, and Seung Ryoul Maeng. Eventor: An authering system for interactive multimedia applications. Multimedia Systems, 2(3):129–140, September 1994.

[fagin1987a] Ronald Fagin and Joseph Y. Halpern. I’m OK if you’re OK: On the notion of trusting communication. In Proceedings of the Symposium on Logic in Computing Science, pages 280–292. IEEE, 1987.

[fairbairns1993a] Robin Fairbairns. Kernel work package. Summary report, Esprit BRA 6865 — Pegasus, November 1993.

[fallmyr1997a] Terje Fallmyr and Tage Stabell-Kulø. Qos applied to security in mobile computing. Computer Science Technical Report 97-29, Department of Computer Science, University of Tromsø, June 1997.

[farhat1995a] Jocelyne Farhat, Isabelle Demeure, and Franco Gasperoni. Automatic QoS support for distributed multimedia applications. In Proceedings of European Research Seminar on Advances in Distributed Systems, ERSADS’95, L’Alpe-d’-Huez, April 1995.

[farmer19??a] Dan Farmer and Wietse Venema. Improving the security of your site by breaking into it. Sun Microsystems.

[farooqui1995a] Kazi Farooqui, Luigi Logrippo, and Jan de Meer. The ISO reference model for open distributed processing: an introduction. Computer Networks and ISDN Systems, 27(8):1215–1229, July 1995.

[farrow1986a] Rik Farrow. Security issues and strategies for users. Unix/World, pages 65–71, April 1986.

[farrow1986b] Rik Farrow. Security for superusers, or how to break the unix system. Unix/World, pages 65–70, May 1986.

[fedaoui1993a] Linda Fédaoui, Wassim Tawbi, and Eric Horlait. Distributed multimedia systems quality of service in ODP framework of abstraction: A first study. In J. De Meer, B. Mahr, and S. Storp, editors, Open Distributed Processing II, Proceedings of the IFIP TC6/WG6.1 International Conference on Distributed Processing, IFIP Transactions, volume C-20, pages 265–274, Berlin, Germany, September 1993. Elsevier Science B. V.

[fedaoui1994a] Linda Fédaoui, Aruna Senevirante, and Eric Horlait. Implementation of a endto-end quality of service managment scheme. In D. Hutchison, A. Danthine, H. Leopold, and G. Coulson, editors, Multimedia Transport and Teleservice, International COST 237 Workshop, volume 882 of Lecture Notes in Computer Science, pages 124–144, Vienna, Austria, November 1994. Springer-Verlag.

[fedaoui1994b] Linda Fédaoui and Eric Horlait. Interfacing a quality of service manager for distributed multimedia applications. In Pedro Veiga and Dipak Khakhar, editors, Proceeding of the IFIP TC6 International Conference on Information Networks and Data Communication, volume C-23 of IFIP Transactions, pages 89–98, Funchal, Madeira Island, Portugal, April 1994. North-Holland Publishing Company.

[feldmeier1993a] David C. Feldmeier. A fremework of architectural concepts for high-speed communication systems. IEEE Journal on Selected Areas in Communications, 11(4):480–488, May 1993.

[ferber1989a] Jacques Ferber. Computational reflection in class based object oriented languages. In OOPSLA ’89 Proceedings, pages 317–326, October 1989.

[ferrari1990a] Domenico Ferrari. Client requirements for real-time communication services. In Proceedings of the International Conference on Information Technology, number 11 in IEEE Communications Magazine, pages 65–72, Tokyo, Japan, October 1990.

[ferrari1992a] Domenico Ferrari, Jean Ramaekers, and Giorgio Ventre. Client-network interaction in quality of service communication environments. In A. Danthine and O. Spaniol, editors, High Performance Networking, IV, Proceedings of the IFIP TC6/WG6.4 Fourth International Conference on High Performance Networking, IFIP Transactions, volume C-14, pages 221–234, Liège, Belgium, December 1992. Elsevier Science B. V.

[ferrari1994a] Domenico Ferrari, Anindo Banerjea, and Hui Zhang. Network support for multimedia — a discussion of the Tenet approach. Computer Networks and ISDN Systems, 26:1267–1280, 1994.

[ferrari1995a] Domenico Ferrari, Amit Gupta, and Giorgio Ventre. Distributed advance reservation of real-time connections. In Proceedings of the Fifth International Workshop on Network and Operating System Support for Digital Audio and Video, Durham, NH, April 1995.

[filipski1986a] Alan Filipski and James Hanko. Making UNIX secure. Byte, pages 113–128, April 1986.

[fischer1997a] Clemens Fischer. CSP-OZ: A combination of Object-Z and CSP. In Howard Bowman and John Derrick, editors, IFIP TC6 WG6.1 International Workshop on Formal Methods for Open Object-based Distributed Systems (FMOODS ’97), volume 2, pages 423–438, Canterbury, Kent, UK, July 1997. International Federation for Information Processing (IFIP), Chapman & Hall.

[fitzpatrick1995a] Geraldine Fitzpatrick, William J. Tolone, and Simon M. Kaplan. Work, locales and distributed social worlds. In H. Marmolin, Y. Sundblad, and K. Schmidt, editors, Proceedings of 1995 European Conference on Computer-Supported Cooperative Work, pages 1–16. Kluwer Academic Publishers, 1995.

[fitzpatrick1998a] Tom Fitzpatrick, Gordon S. Blair, Geoff Coulson, Nigel Davies, and Philippe Robin. Supporting adaptive multimedia applications through open bindings. In Proceedings of the 4th International Conference on Configurable Distributed Systems (ICCDS ’98), Annapolis, Maryland, US, May 1998.

[fokkinga1992a] Maarten M. Fokkinga. Calculate categorically! Formal Aspects of Computing, 4(2):673–692, 1992.

[fokkinga1993a] Maarten Fokkinga and Lambert Meertens. Adjunctions. Technical report, University of Twente, June 1993.

[fokkinga1994a] Maarten M. Fokkinga. A gentle introduction to category theory, the calculation approach. Technical Report Version of June 6, 1994, University of Twente, June 1994.

[fongen1992a] Anders Fongen. Perl, alt i ett. PCW Nettverk, 2:70–72, 1992.

[fongen1994a] Anders Fongen. Rydd opp i kaoset med ISDN. PCW Nettverk, 3:8–14, 1994.

[forman1994a] Ira R. Forman, Scott Danforth, and Hari Madduri. Composition of before/after metaclasses in SOM. In OOPSLA ’94 Proceeding, pages 427–439, October 1994.

[forman1996a] Ira R. Forman and Scott H. Danforth. Inheritance of metaclass constraints in SOM. In Gregor Kiczales, editor, Proceedings of the Reflection’96 Conference, pages 185– 202, April 1996.

[fritzsche1996a] J. Christian Fritzsche. Multimedia building blocks for distributed applications. In Proceedings of the International Workshop on Multimedia Software Development, pages 41–49, Berlin, Germany, March 1996. IEEE Computer Society Press.

[fry1993a] Michael Fry, Anthony Richards, and Aruna Senevirante. Framework for implementing the next generation of communication protocols. In D. Shepherd, G. Blair, G. Coulson, N. Davies, and F. Garcia, editors, Proceedings of the 4th International Workshop on Network and Operating System Support for Digital Audio and Video, volume 846 of Lecture Notes in Computer Science, pages 138–148, Lancaster, U.K., November 1993. Springer-Verlag.

[fulton1996a] Jim Fulton. Extension classes, Python extension types become classes. Technical report, Digital Creations, 1996.

[furth1994a] Borko Furth. Multimedia systems: An overview. IEEE Multimedia, 1(1):47–59, Spring 1994.

[gabriel1991a] Richard P. Gabriel. Lisp: Good news, bad news, and how to win big. Technical report, Lucid, Inc., 1991.

[gaines1978a] R. Stockton Gaines and Norman Z. Shapiro. Some security principles and their application to computer security. Operating Systems Review, 12(3):19–28, 1978.

[gallimore1989a] R. M. Gallimore, D. Coleman, and V. Stavridou. UMIST OBJ: a language for executable program specifications. The Computer Journal, 32(5):413–421, 1989.

[ganger1995a] Gregory Ganger and Yale Patt. Soft updates: A solution to the metadata update problem in file systems. Technical Report CSE-TR-254-95, Computer Science and Engineering Division, University of Michigan, 1995.

[garcia1996a] F García, D. Hutchison, and N. Yeadon. QoS support for distributed multimedia communications. In IFIP/IEEE International Conference on Distributed Platforms, Dresden, Germany, 1996.

[garey1979a] Michael R. Garey and David S. Johnson. Computers and Intractability — A Guide to the Theory of NP-Completeness, chapter 1–3,7. A Series of Books in the Mathematical Sciences. W. H. Freeman and Company, 1979.

[garlan1998a] David Garlan. Higher-order connectors. In Proceedings of Workshop on Compositional Software Architectures, January 1998.

[gasser1989a] Morrie Gasser, Andy Goldstein, Charlie Kaufman, and Butler Lampson. The Digital distributed system security architecture. In Proceedings of 1989 National Computer Security Conference, 1989.

[gay1995a] Valérie Gay, Peter Leydekkers, and Robert Huis in ’t Veld. Specification of multiparty audio and video interaction based on the reference model for open distributed processing. Computer Networks and ISDN Systems, 27(8):1247–1262, July 1995.

[geihs1994a] Kurt Geihs, Birgitte Bär, and Arno Puder. Towards open service environments. In K. P. Birman, F. Mattern, and A. Schiper, editors, Theory and Practice in Distributed Systems, International Workshop, Selected Papers, volume 938 of Lecture Notes in Computer Science, pages 153–163, Dagstuhl Castle, Germany, September 1994. Springer-Verlag.

[gelernter1985a] David Gelernter. Generative communication in Linda. ACM Transactions on Programming Languages and Systems, 7(1):80–112, January 1985.

[gerber1991a] Richard Gerber and Insup Lee. Specifications and analysis of resource-bound realtime systems. In J. W. de Bakker, C. Huizing, W. P. de Roever, and G. Rozenberg, editors, Proceedings of the REX Workshop 1991, Real-Time: Theory in Practice, volume 600 of Lecture Notes in Computer Science, pages 371–396, Mook, The Netherlands, June 1991. Springer-Verlag.

[gerber1994a] Richard Gerber. Languages and tools for real-time systems: Problems, solutions and opportunities. University of Maryland Technical Report CS-TR-3362, UMIACS-TR-94-117, Department of Computer Science, University of Maryland, October 1994.

[german1992a] Steven M. German and A. Prasad Sistla. Reasoning about systems with many processes. Journal of the Association for Computing Machinery, 39(3):675–735, July 1992.

[gettys19??a] James Gettys. Problems implementing window systems in UNIX. Digital Equipment Corporation, Project Athena, Massachusetts Institute of Technology.

[geyer1996a] Werner Geyer, Christoph Bernhardt, and Ernst Biersack. A synchronization scheme for stored multimedia streams. In Berthold Butscher, Eckhard Moeller, and Herwart Pusch, editors, Interactive Distributed Multimedia Systems and Services — European Workshop IDMS’96, volume 1045 of Lecture Notes in Computer Science, pages 277–295, Berlin, Germany, March 1996. Springer-Verlag.

[ghafoor1993a] Arif Ghafoor and P. Bruce Berra. Multimedia database systems. In Nabil R. Adam and Bharat K. Bhargava, editors, Advanced Database Systems, volume 759 of Lecture Notes in Computer Science, pages 397–411. Springer-Verlag, 1993.

[gibbon1993a] J. F. Gibbon and T. D. C. Little. Real-time data delivery for multimedia networks. In Proceedings of the 18th Annual Conference on Local Computer Networks, pages 7–16, Minneapolis, MN, September 1993.

[gibbs1992a] Simon Gibbs. Application construction and component design in an object-oriented multimedia framework. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 394–398, La Jolla, California, USA, November 1992. Springer-Verlag.

[gibson1989a] Michael Lucas Gibson. The CASE philosophy. Byte, pages 209–218, April 1989.

[gien1991a] Michel Gien. Next generation operating systems architecture. In Proceedings of the International Workshop on Operating Systems of the 90s and Beyond, volume 563 of Lecture Notes in Computer Science, Dagstuhl Castle, Germany, July 1991. Springer-Verlag.

[gien1992a] Michel Gien. From operating systems to cooperative operating environments. Technical Report CS/TR-92-82, Chorus Systèmes, October 1992.

[gien1995a] Michel Gien. Evolution of the CHORUS open microkernel architecture: The STREAM project. In Proceedings of 5th. IEEE Workshop on Future Trends in Distributed Computing Systems, FTDCS’95, Cheju Island, Korea, August 1995.

[girard1987a] J.-Y. Girard. Linear logic. Theoretical Computer Science, 50:1–102, 1987.

[glass1989a] Brett Glass. The token ring. Byte, pages 363–376, January 1989.

[goedicke1997a] Michael Goedicke and Torsten Meyer. Design and evaluation of distributed component-oriented software systems. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Component-Oriented Programming, volume 1357 of Lecture Notes in Computer Science, pages 348–352, Jyväskylä, Finland, June 1997. Springer-Verlag.

[goedicke1999a] Michael Goedicke and Torsten Meyer. Towards dynamic semantic-directed configuration management. In Engineering Distributed Objects Workshop, ICSE’99, Los Angeles, May 1999. Position Paper.

[goguen1987a] Joseph A. Goguen and Josè Meseguer. Unifying functional, object-oriented and relational programming with logical semantics. In Shriver and P. Wegner, editors, Research Directions in Object-Oriented Programming, pages 417–477. MIT Press, 1978.

[goldberg1997a] Jack Goldberg and Robert J. Stroud. Adaptive fault-tolerant systems and reflective architectures. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Reflective Real-Time Object-Oriented Programming and Systems, volume 1357 of Lecture Notes in Computer Science, pages 80–88, Jyväskylä, Finland, June 1997. Springer-Verlag.

[golding1991a] Richard Golding and Darrell D. E. Long. Accessing replicated data in a largescale distributed system. Technical report, Concurrent Systems Laboratory. Baskin Center for Computer Engineering & Information Sciences, University of California, January 1991.

[golding19??a] Richard A. Golding, Darrell D. E. Long, and John Wilkes. The refdbms distributed bibliographic database system.

[golm1997a] Michael Golm and Jürgen Kleinöder. Implementing real-time actors with Meta-Java. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Reflective Real-Time Object-Oriented Programming and Systems, volume 1357 of Lecture Notes in Computer Science, pages 68–73, Jyväskylä, Finland, June 1997. Springer-Verlag.

[golub1990a] David Golub, Randall Dean, Alessandro Forin, and Richard Rashid. Unix as an application program. In Proceedings of the Summer 1990 USENIX Conference, pages 87–95, 1990.

[gomaa1999a] H. Gomaa and G. Farrukh. Software engineering of a distributed object architecture for federated client/server systems. In Engineering Distributed Objects Workshop, ICSE’99, Los Angeles, May 1999. Position Paper.

[gong1995a] Yihong Gong, Lim Teck Sin, Chua Hock Chuan, Hongjiang Zhang, and Masao Sakauchi. Automatic parsing af TV soccer programs. In Proceedings of the International Conference on Multimedia Computing and Systems, pages 167–174, Washington, D.C., May 1995. IEEE Computer Society Press.

[goos1972a] Gerhard Goos. Some basic principles in structuring operating systems. In C. A. R. Hoare and R. H. Perrott, editors, Operating Systems Techniques, Proceeding of a Seminar Held at Queen’s University, pages 94–100, Belfast, 1972. Academic Press.

[gopalakrishnan1990a] Ganesh Gopalakrishnan and Prabhat Jain. Some recent asynchronous system design methodologies. Technical Report UU-CS-TR-90-016, Department of Computer Science, University of Utah, October 1990. (Submitted to ACM Computing Surveys).

[gopalakrishnan1994a] R. Gopalakrishnan and G. Parulkar. Efficient quality of service support in multimedia computer operating systems. Technical Report WUCS-94-26, Department of Computer Science, Washington University, 1994.

[gopalakrishnan1994b] R. Gopalakrishnan and G. Parulkar. Application level protocol implementations to provide quality-of-service guarantees at endsystems. In 9th IEEE Computer Communications Conference, 1994.

[gopalakrishnan1995a] R. Gopalakrishnan and G. Parulkar. A framework for QoS guarantees for multimedia applications within and endsystem. In German and Swiss Computer Science Soc. Conference, September 1995.

[gopalakrishnan1995b] R. Gopalakrishnan and G. Parulkar. Real-time upcalls: A mechanism to provide real-time processing guarantees. Technical Report WUCS-95-06, Department of Computer Science, Washington University, 1995.

[gopalakrishnan1995c] R. Gopalakrishnan and G. Parulkar. RMDP — a real-time CPU scheduling algorithm to provide QoS guarantees for protocol processing. In IEEE RTAS’95, Chicago, May 1995.

[gopalakrishnan1995d] R. Gopalakrishnan and G. Parulkar. Quality of service support for protocol processing within endsystems. In High-Speed Networks for Multimedia Applications Workshop, Dagstuhl, July 1995.

[gopalakrishnan1996a] R. Gopalakrishnan and G. Parulkar. Bringing real-time scheduling and practice closer for multimedia computing. In SIGMETRICS’96, Philadelphia, May 1996.

[gopalakrishnan1996b] R. Gopalakrishnan and G. Parulkar. Efficient user space protocol implementations with QoS guarantees using real-time upcalls. Technical Report WUCS-96-11, Department of Computer Science, Washington University, 1996.

[gorn196?a] Saul Gorn. Specification languages for mechanical languages and their processors — a baker’s dozen. Communications of the ACM, ??(?):532–542, 196?

[goscinski1993a] A. Goscinski and Y. Ni. Object trading in open systems. In J. De Meer, B. Mahr, and S. Storp, editors, Open Distributed Processing II, Proceedings of the IFIP TC6/WG6.1 International Conference on Distributed Processing, IFIP Transactions, volume C-20, pages 145–156, Berlin, Germany, September 1993. Elsevier Science B. V.

[gowing1996a] Brendan Gowing and Vinny Cahill. Meta-object protocols for C++: The Iguana approach. In Gregor Kiczales, editor, Proceedings of the Reflection’96 Conference, pages 137–152, April 1996.

[grampp1984a] F. T. Grampp and R. H. Morris. UNIX operating system security. AT&T Bell Laboratories Technical Journal, 63(8):1649–1672, October 1984.

[gray1981a] Jim Gray. The transaction concept: Virtues and limitations. In Seventh International Conference on Very Large Data Bases. IEEE, September 1981.

[grosky1994a] William I. Grosky. Multimedia information systems. IEEE Multimedia, 1(1):12–24, Spring 1994.

[grunbacher1999a] Andreas Grünbacher and Mehdi Jazayeri. Protocol-based runtime monitoring of dynamic distributed systems. In Engineering Distributed Objects Workshop, ICSE’99, Los Angeles, May 1999. Position Paper.

[guerraoui1995a] Rachid Guerraoui. Revisting the relationship between non-blocking atomic commitment and consensus. In Jean-Michel Hélary and Michel Raynal, editors, Distributed Algorithms, 9th International Workshop, WDAG ’95, volume 972 of Lecture Notes in Computer Science, pages 87–100, Le Mont-Saint-Michel, France, September 1995. Springer-Verlag.

[guerraoui1998a] Rachid Guerraoui, Pascal Felber, Benoît Garbinato, and Karim Mazouni. System support for object groups. In Proceeding OOPSLA’98, Vancouver, Canada, October 1998.

[guha1994a] Aloke Guha and Allalaghatta Pavan. Real-time support of continuous and variable bit rate traffic on ATM network. In Proceedings of the 14th International Conference on Distributed Computing Systems, pages 268–275, Poznan, Poland, June 1994. IEEE Computer Society Press.

[guha1995a] Aloke Guha, Allalaghatta Pavan, Jonathan C. L. Liu, and Barry A. Roberts. Controlling the process with distributed multimedia. IEEE Multimedia, 2(2):20–29, Summer 1995.

[guillemont1991a] Marc Guillemont, Jim Lipkis, Doug Orr, and Marc Rozier. A second-generation micro-kernel based UNIX; lessons in performance and compatibility. In Proceedings of the Winter 1991 USENIX Conference, pages 13–22, 1991.

[gurchom1995a] M. M. W. A. van Gurchom, W. B. Teeuw, and M. A. W. Houtsma. Criticla success factors for multimedia applications. In Advanced Course: Multimedia Databases in Perspective, pages 337–368, June 1995.

[gutfreund1995a] S. Gutfreund, J. R. Nicol, V. Phuah, and F. Eliassen. ATOMS: ATM support mechanisms for advanced multimedia application research. Technical Report TR-0297-06-95-380, GTE Laboratories Incorporated, June 1995.

[guttag1976a] John V. Guttag, Ellis Horowitz, and David R. Musser. The design of data type specifications. In Proceedings Second International Conference on Software Engineering, pages 414–420, San Francisco, October 1976.

[guttag1977a] John Guttag. Abstract data types and the development of data structures. Communications of the ACM, 20(6):396–404, June 1977.

[guttag1978a] J. V. Guttag and J. J. Horning. The algebraic specification of abstract data types. Acta Informatica, 10:27–52, 1978.

[guttag1978b] John V. Guttag, Ellis Horowitz, and David R. Musser. Abstract data types and software validation. Communications of the ACM, 21(12):1048–1064, December 1978.

[guttag1985a] J. V. Guttag, J. J. Horning, and J. M. Wing. Piece I — the larch family of specification languages. In Larch in Five Easy Pieces (SRC Research Reports 5), pages 3–21. DEC’s System Research Center, July 1985.

[habert1990a] Sabine Habert. COOL: Kernel support for object-oriented environments. In Norman Meyrowitz, editor, Proceedings of ECOOP’90/OOPSLA’90, volume 25(10) of Sigplan Notices, pages 269–277, Ottawa, Canada, October 1990. ACM Press.

[hafid1995a] Abdelhakim Hafid and Gregor V. Bochmann. An approach to quality of service management for distributed multimedia applications. In Procceedings of the Third IFIP International Conference on Open Distributed Computing, Brisbane, Australia, 1995.

[hafid1996a] Abdelhakim Hafid, Gregor V. Bochmann, and Brigitte Kerhervé. A quality of service negotiation procedure for distributed multimedia presentational applications. In Proceedings of the Fifth international IFIP Symposium of High Performance of Distributed Processing (HPDC-5), Syracuse, USA, 1996.

[hafid1998a] Abdelhakim Hafid, Gregor V. Bochmann, and Rachida Dssouli. Distributed multimedia application and quality of service: A review. Electronic Journal on Networks and Distributed Processing, (6):1–50, February 1998.

[haitzler1999a] Carsten Haitzler. Esound. Technical report, Red Hat Software, Inc., 1999.

[halpern1985a] Joseph Y. Halpern and Ronald Fagin. A formal model of knowledge, action, and communication in distributed systems: Preliminary report. In Procceedings of the 4th ACM Symposium on Principles of Distributed Computing, pages 224–236, Ontario, August 1985.

[hamilton1993a] Graham Hamilton and Panos Kougiouris. The Spring nucleus: A microkernel for objects. Technical Report SMLI TR-93-14, Sun Microsystems Laboratories, Inc., April 1993.

[hanssen1997a] Øyvind Hanssen.  FlexiNet — QoS investigation.  Technical Report

APM.1997.01.00, ANSA Phase III, APM, June 1997.

[hanssen1997b] Øyvind Hanssen. FlexiNet — extensible kernel investigation. Technical Report APM.2002.01.00, ANSA Phase III, APM, July 1997.

[hanssen1997c] Øyvind Hanssen.  FlexiNet — binding framework.  Technical Report

APM.2057.02.00, ANSA Phase III, APM, October 1997.

[hanssen1998a] Øyvind Hanssen and Frank Eliassen. Towards a QoS aware binding model. In SYBEN’98, Zürich, May 1998.

[hanssen1999a] Øyvind Hanssen. A framework for policy bindings. Technical report, Faculty of Engineering, Agder College, 1999. Submitted DOA’99.

[haraszti1998a] Péter Haraszti, Brendan Gowing, and Vinny Cahill. The use of reflective programming in telecommunication management applications. In ECOOP’98 Workshop on Reflective Object-Oriented Programming Systems, Brussels, Belgium, July 1998.

[harel1995a] David Harel. Will i be pretty, will i be rich? some thoughts on theory vs. practice in system engineering. In Insup Lee and Scott A. Smolka, editors, CONCUR ’95: Concurrency Theory, 6th International Conference, volume 962 of Lecture Notes in Computer Science, pages 264–268, Philadelphia, PA, USA, August 1995. Springer-Verlag.

[harman1995a] Dale D. Harman, Gang Huang, Gi-Hong Im, Mai-Huong Nguyen, Jean-Jacques Werner, and Michael K. Wong. Local distribution for IMTV. IEEE Multimedia, 2(3):14–23, Fall 1995.

[hartley1972a] D. F. Hartley. Techniques in the Titan operating system. In C. A. R. Hoare and R. H. Perrott, editors, Operating Systems Techniques, Proceeding of a Seminar Held at Queen’s University, pages 271–276, Belfast, 1972. Academic Press.

[hartvigsen1988a] Gunnar Hartvigsen and Dag Johansen. StormCast — a distributed artifical intelligence application for serve storm forecasting. In Proceedings of Distributed Computer Control Systems 1988, pages 99–102, Vitznau, Switzerland, 1988. Pergamon Press.

[hartvigsen1989a] Gunnar Hartvigsen and Dag Johansen. Cooperation through information interchange in StormCast. In Proceedings of the 9th IFAC Workshop on Distributed Computer Control Systems, DCCS-89, Tokyo, Japan, September 1989.

[hartvigsen1990a] Gunnar Hartvigsen and Dag Johansen. Cooperation in a distributed artificial intelligence environment — the StormCast application. CS Research Report 90-05, Department of Computer Science, University of Tromsø, June 1990.

[hartvigsen1990c] Gunnar Hartvigsen and Dag Johansen. A multiagent architecture for a distributed artificial intelligence application. CS Research Report 90-04, Department of Computer Science, University of Tromsø, June 1990.

[hartvigsen1990d] Gunnar Hartvigsen and Dag Johansen. Co-operation in a distributed artificial intelligence environment — the StormCast application. Engineering Applications of AI, 3:229–237, September 1990.

[hartvigsen1990e] Gunnar Hartvigsen and Dag Johansen. The StormCast application — version 1.8. CS Research Report 90-09, Department of Computer Science, University of Tromsø, December 1990. (Preliminary).

[hartvigsen1995a] Gunnar Hartvigsen, Arne Helme, and Stig Johansen. A secure architecture for software agents: The virtual secretary approach. Technical report, Department of Computer Science, University of Tromsø and Faculty of Computer Science / SPA, University of Twente, 1995.

[hauck1998a] Franz J. Hauck, Ulrich Becker, Martin Geier, Erich Meier, Uwe Rastofer, and Martin Steckermeier. AspectIX: A middleware for aspect-oriented programming. In ECOOP ’98 Workshop on Aspect-Oriented Programming, July 1998. Position Paper.

[hayter1994a] Mark Hayter and Derek McAuley. Multimedia streams and the CPU cache. Technical report, University of Cambridge, Computer Laboratory, March 1994.

[hayton1997a] Richard Hayton.  FlexiNet open ORB framework.  Technical Report

APM.2047.01.00, ANSA Phase III, APM, 1997.

[helbig1996a] Tobias Helbig and Kurt Rothermel. An architecture for a distributed stream synchronization service. In Berthold Butscher, Eckhard Moeller, and Herwart Pusch, editors, Interactive Distributed Multimedia Systems and Services — European Workshop IDMS’96, volume 1045 of Lecture Notes in Computer Science, pages 297–314, Berlin, Germany, March 1996. Springer-Verlag.

[helbig1997a] Tobias Helbig, Stefan Tretter, and Dirk Trossen. Combining CORBA and ITU-T.120 to an efficient conferencing service. In Ralf Steinmetz and Lars C. Wolf, editors, IDMS’97 — 4th International Workshop on Interactive Distributed Multimedia Systems and Telecommunication Services, volume 1309 of Lecture Notes in Computer Science, Darmstadtm Germany, September 1997. Springer-Verlag.

[helme1990a] Arne Helme. Introduksjon til bruk av HP 64000-UX design center — utviklingsmiljø for maskinær programmering. Technical report, Seksjon for Datafag, Universitetet i Tromsø, 1990.

[helme1990b] Arne Helme. HP 64000 design center — utviklingsmiljø for maskinær programmering. Technical report, Seksjon for Datafag, Universitetet i Tromsø, 1990.

[helme1990c] Arne Helme. Evaluering av HP 64000-UX — utviklingsmiljø for maskinær programmering. Technical report, Seksjon for Datafag, Universitetet i Tromsø, 1990.

[helme1992b] Arne Helme. Kausal og atomisk multicast, en innføring. Notat fra seminar om distribuerte systemer, Seksjon for Datafag, Universitetet i Tromsø, March 1992.

[helme1993a] Arne Helme. Trip report, Palo Alto, May 1993.

[helme1993b] Arne Helme. –Trust– security support for ubiquitos computing. Broadcast Technical Report 93-XX, Faculty of Computer Science, University of Twente, July 1993. (Draft).

[helme1993c] Arne Helme. Secure ubiquitous computing, November 1993.

[helme1995a] Arne Helme and Erik Plaggenmarsch. SCIP: A secure network layer under Unix. Technical report, Faculty of Computer Science / SPA, University of Twente, May 1995.

[helme1997b] Arne Helme and Tage Stabell-Kulø. Security functions for a file repository. Operating Systems Review, 31(2):3–8, April 1997.

[herman1995a] I. Herman, G. J. Reynolds, and J. Van Loo. PREMO: an emerging standard for multimedia presentation. CWI research report CS-R9554, Department of Interactive Systems, CWI, 1995.

[hernadez1997a] Juan Hernádez, Michael Papathomas, Juan M. Murillo, and Fernando Sánchez. Coordinating concurrent objects: How to deal with the coordination aspect? Technical report, Computer Science Department, Extremadura University and Computer Department, Lancaster University, 1997.

[herrmann1992a] Bènèdicte Herrmann and Laurent Philippe. Chorus/MiX, a distributed UNIX, on multicomputers. In Proceedings of Transputer’92, Arc et Senans, France, May 1992.

[herrtwich1992a] Ralf Guido Herrtwich and Luca Delgrossi. Beyond ST-II: Filfilling the requirements of multimedia communication. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 25–31, La Jolla, California, USA, November 1992. Springer-Verlag.

[hess1999a] Christopher K. Hess and Roy H. Campbell. Media streaming protocol: An adaptive protocol for the delivery of audio and video over the internet. In IEEE International Conference on Multimedia Computing Systems, Florence, Italy, July 1999.

[hildebrand1992a] Dan Hildebrand. An architectural overview of QNX. In Prooceedings of the USENIX Workshop on Micro-Kernels and other Kernel Architectures, pages 113– 126, 1992.

[hill1994a] Jonathan M. D. Hill and Keith Clarke. An introduction to category theory, category theory monads, and their relationship to functional programming. Technical Report QMW-DCS-681, Department of Computer Science, Queen Mary & Westfield College, University of London, August, 1994.

[hirzalla1995a] Naél Hirzalla, Bne Falchuk, and Ahmed Karmouch. A temporal model for interactive multimedia scenarios. IEEE Multimedia, 2(3):24–31, Fall 1995.

[hoare1969a] C. A. R. Hoare. An axiomatic basis for computer programming. Communications of the ACM, 12(10):576–583, October 1969.

[hoare1972a] C. A. R. Hoare. Operating systems: Their pupose, objectives, functions and scope. In C. A. R. Hoare and R. H. Perrott, editors, Operating Systems Techniques, Proceeding of a Seminar Held at Queen’s University, pages 11–19, Belfast, 1972. Academic Press.

[hoare1972b] C. A. R. Hoare. Towards a theory of parallel programming. In C. A. R. Hoare and R. H. Perrott, editors, Operating Systems Techniques, Proceeding of a Seminar Held at Queen’s University, pages 61–71, Belfast, 1972. Academic Press.

[hoare1972c] C. A. R. Hoare and D. C. S. Allison. Incomputability. ACM Computing Surveys, 4(3):169–178, September 1972.

[hoare1974a] C. A. R. Hoare. Monitors: An operating system structuring concept. Communications of the ACM, 17(10):549–557, October 1974.

[hoare1981a] C. A. R. Hoare. The emperor’s old clothes. Communications of the ACM, 24(2):75– 83, February 1981.

[hof2000a] Markus A. Hof. Using reflection for composable message semantics. In ECOOP’2000 Workshop on Reflection and Metalevel Architectures, Sophia Antipolis and Cannes, France, June 2000.

[hoffmann199?a] Frank Oliver Hoffmann and Luca Delgrossi. ST-II for the heidelberg transport system: Protocol, design and implementation. Technical report, IBM European Networking Center, 199?

[hoffner1994a] Yigal Hoffner. Monitoring in distributed systems. Architecture Report APM.1008.01, ANSA, Architecture Projects Managment Limited, October 1994.

[hogan1988a] Carole B. Hogan. Protection imperfect: The security of some computing environments. Operating Systems Review, 22:7–27, July 1988.

[holder1997a] Ophir Holder and Israel Ben-Shaul. A reflective model for mobile software objects. In Proceedings of the 17th International Conference on Distributed Systems, Baltimore, Maryland, USA, May 1997. IEEE Computer Society Press.

[hondt1997a] Koen De Hondt, Carine Lucas, and Patrick Steyaert. Reuse contracts as component interface descriptions. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Component-Oriented Programming, volume 1357 of Lecture Notes in Computer Science, pages 338–342, Jyväskylä, Finland, June 1997. Springer-Verlag.

[hong1999a] James Won-Ki Hong, Jong-Seo Kim, and Jae-Kyu Park. A CORBA-based quality of service management framework for distributed multimedia services and applications. IEEE Network, 13(2):70–79, March/April 1999.

[howes1995a] Timothy A. Howes and Mark C. Smith. A scalable, deployable directory service framework for the internet. CITI Technical Report 95-7, Center for Information Technology Integration, University of Michigan, July 1995.

[howes1995b] Timothy A. Howes. The lightweight directory access protocol: X.500 lite. CITI Technical Report 95-8, Center for Information Technology Integration, University of Michigan, July 1995.

[howes1995c] Timothy A. Howes. An X.500 and LDAP database: Design and implementation. Technical report, Center for Information Technology Integration, University of Michigan, 1995.

[hoydalsvik1993a] Stein Høydalsvik. Status StormCast-samarbeidet. Rapport, NIT-Nord, July 1993.

[hutchison1994a] David Hutchison, Geoff Coulson, Andrew Campbell, and Gordon S. Blair. Quality of service managment in distributed systems. In Morris Sloman, editor, Network and Distributed Systems Management, chapter 11, pages 273–302. Addison-Wesley, 1994.

[huynh1995a] Khoa D. Huynh and Taghi M. Khoshgoftaar. A performance analysis of an object-based I/O architecture in a video server environment. Multimedia Systems, 3(4):162–177, September 1995.

[hvasshovd1991a] Svein O. Hvasshovd, Tore Sæter, and Øystein Torbjørnsen. Critical issues in the design of a fault-tolerant multiprocessor database server. In Proceedings of PRFTS’91, 1991.

[iannella1995a] Renato Iannella. Internet resource discovery with BabyOIL. In Proceedings of AUUG Queensland Summer Technical Conference 95. AUUG, 1995.

[icaza1999a] Miguel de Icaza and Jonathan Blandford. Components in the GNOME project. Technical report, Red Hat Software, Inc., July 1999.

[indulska1993a] J. Indulska, M. Bearman, and K. Raymond. A type managment system for an ODP trader. In J. De Meer, B. Mahr, and S. Storp, editors, Open Distributed Processing II, Proceedings of the IFIP TC6/WG6.1 International Conference on Distributed Processing, IFIP Transactions, volume C-20, pages 169–180, Berlin, Germany, September 1993. Elsevier Science B. V.

[ingalls1997a] Dan Ingalls, Ted Kaehler, John Maloney, Scott Wallace, and Alan Kay. Back to the future: The storu of Squeak, a practical smalltalk written in itself. Technical report, Apple Computer, 1997.

[ingold1992a] Daniel P. Ingold. MEGAPHONE: A multimedia application based om objectoriented communication. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 357–362, La Jolla, California, USA, November 1992. Springer-Verlag.

[inouye1997a] Jon Inouye, Shanwei Cen, Calton Pu, and Jonathan Walpole. System support for mobile multimedia applications. In NOSSDAV’97, pages 143–154, St. Louis, Missouri, May 1997.

[isg1997a] International System Group. Middleware white paper. Technical report, International System Group, Inc., February 1997.

[ishikawa1996a] Yutaka Ishikawa Atsushi Hori, Mitsuhisa Sato, Motohiko Matsuda, Jórg Nolte, Hiroshi Tezuka, Hiroki Konaka, Munenori Maeda, and Kazuto Kubota. Design and implementation of metalvel architecture in C++: MPC++ approach. In Gregor Kiczales, editor, Proceedings of the Reflection’96 Conference, pages 153–166, April 1996.

[itoh1995a] Jun ichiro Itoh, Yasuhiko Yokote, and Rodger Lea. Using meta-objects to support optimization in the Apertos operating system. In USENIX Conference on Object-Oriented Technologies (COOTS), Monterey, California, June 1995.

[jacquemot1994a] C. Jacquemot, F. Herrmann, P.S. Jensen, P. Gautron, J. Mukerji, H.G. Baumgarten, and H. Hartlage. An overview of COOL and COOL-IDL. Ouverture OU/TR-94-3, Chorus Systèmes, Novell/USG, Siemens AG, March 1994.

[jacquemot1994b] C. Jacquemot, F. Herrmann, P.S. Jensen, P. Gautron, J. Mukerji, H.G. Baumgarten, and H. Hartlage. COOL: The CHORUS CORBA compliant framework. In Proceedings of Spring COMPCON 94, San Francisco, CA, 1994.

[jacquemot1994c] C. Jacquemot, F. Herrmann, P.S. Jensen, P. Gautron, J. Mukerji, H.G. Baumgarten, and H. Hartlage. COOL-IDL extensions to C++ to support distributed programming. In Advanced Workshop on Distributed Computing, USENIX C++ Conference, Cambridge, April 1994.

[jaeschke1991a] Rex Jaeschke. Standard C — a status report. Dr. Dobb’s Journal, pages 16–26, August 1991.

[jakobsen2003a] Anna-Brith Jakobsen and Randi Karlsen. Towards a reflective transactional middleware platform for advanced applications. In NIK 2003, November 2003.

[jammel1977a] Alfons J. Jammel and Helmut G. Stiegler. Managers versus monitors. In Bruce Gilchrists, editor, Information Processing 77 — Proceedings of IFIP Congress 77, volume 7 of IFIP Congress Series, Toronto, August 1977. North-Holland Publishing Company.

[jansen1992a] Pierre G. Jansen and Fred Gansevles. Priority inheritance in real-time micro kernels. In Proceedings of the International Symposium on Computer and Information Sciences VII, pages 211–220, Antalya, Turkey, 1992.

[jansen1994a] Pierre G. Jansen and Paul Sijben. Real-time in multimedia: Opportunistic scheduling or quality of service contracts? In Distributed Multimedia Systems and Applications, pages 194–197, Honolulu, Hawaii, August 1994.

[jansen1994b] Pierre G. Jansen and Paul Sijben. Kernel support in a distributed multimedia environment. In First International workshop on Real-Time Computing Systems and Applications, pages 117–121, Seoul National University, Seoul, Korea, December 1994.

[jansen1996a] Pierre G. Jansen and Emiel Wijgerink. Flexible scheduling by deadline inheritance in soft real time kernels. Memoranda Informatica 96-04, Computer Science Department, University of Twente, January 1996.

[jansen1996b] Pierre G. Jansen and Emiel Wijgerink. Flexible real time scheduling of continuous media streams: A multimedia experiment. In Multimedia 96 Conference, pages 323–330, Yokohama, Japan, March 1996.

[janssen1994a] Bill Janssen and Mike Spreitzer. ILU: Inter-language unification via object modules. In OOPSLA ’94 Workshop on Multi-Language Object Models, 1994.

[jardetzky1995a] Paul W. Jardetzky, Cormac J. Sreenan, and Roger M. Needham. Storage and synchronization for distributed continuous media. Multimedia Systems, 3(4):151– 161, September 1995.

[jifeng1987a] He Jifeng and C. A. R. Hoare. Algebraic specification and proof of a distributed algorithm. Distributed Computing, 2:1–12, 1987.

[johansen1988a] Dag Johansen. Weather forecasting — distributed in nature. In Kiril Boyanov and Rasko Angelinov, editors, Proceedings of the IFIP TC6/TC8 Open Symposium on Network Information Processing Systems, pages 197–203, Sofia, Bulgaria, May 1988. IFIP.

[johansen1991a] Dag Johansen and Gunnar Hartvigsen. StormCast — a distributed application. In Distributed Open Systems in Perspective — Proceedings of the Spring 1991 EurOpen Conference, pages 273–286, Tromsø, Norway, May 1991. EurOpen.

[johansen1991b] Dag Johansen and Arne Munch-Ellingsen. A comparative evaluation of different platforms for distributed computing. In Proceedings of NIK’91, 1991.

[johansen1993a] Dag Johansen. StormCast — a ‘computing the future’ approach. Technical report, Department of Computer Science, University of Tromsø, March 1993.

[johansen1993b] Dag Johansen. A distributed approach to the design of application. In Proceedings of ICCI’93, Canada, May 1993.

[johnson1990a] David S. Johnson. A catalog of complexity classes. In J. van Leeuwen, editor, Handbook of Theoretical Comuter Science, chapter 2, pages 67–161. Elsevier Science B. V., 1990.

[johnson1991a] Brad Curtis Johnson. A distributed computing environment framework: An OSF perspective. In Distributed Open Systems in Perspective — Proceedings of the Spring 1991 EurOpen Conference, pages 69–87, Tromsø, Norway, May 1991. EurOpen.

[jokstad1997a] H. Jokstad and M. N. Louka. The Software Bus Communication System: User’s Guide. Institutt for Energiteknikk, OECD Halden Reactor Project, 2.0 edition, June 1997.

[jones1977a] Anita K. Jones. The narrowing gap between languages systems and operating systems. In Bruce Gilchrists, editor, Information Processing 77 — Proceedings of IFIP Congress 77, volume 7 of IFIP Congress Series, Toronto, August 1977. North-Holland Publishing Company.

[jones1993a] Michael B. Jones. Adaptive real-time resource management supporting modular composition of digital multimedia services. In D. Shepherd, G. Blair, G. Coulson, N. Davies, and F. Garcia, editors, Proceedings of the 4th International Workshop on Network and Operating System Support for Digital Audio and Video, volume 846 of Lecture Notes in Computer Science, pages 21–28, Lancaster, U.K., November 1993. Springer-Verlag.

[jones1993b] Simon L. Peyton Jones, John Hughes, and John Launchbury. How to give a good research talk. Sigplan Notices, 28(11), November 1993.

[joseph1991a] Mathai Joseph. Problems, promises and performance: Some questions for realtime system specification. In J. W. de Bakker, C. Huizing, W. P. de Roever, and G. Rozenberg, editors, Proceedings of the REX Workshop 1991, Real-Time: Theory in Practice, volume 600 of Lecture Notes in Computer Science, pages 315– 324, Mook, The Netherlands, June 1991. Springer-Verlag.

[kaashoek1989a] M. Frans Kaashoek, Andrew S. Tanenbaum, Susan Flynn Hummel, and Henri E. Bal. An efficient reliable broadcast protocol. Operating Systems Review, 23(4):5–19, October 1989.

[kaashoek1996a] M. Frans Kaashoek, Dawson R. Engler, Gregory R. Ganger, and Deborah A. Wallach. Server operating systems. In Proceedings of the 1996 SIGOPS European Workshop, pages 141–148, Connemara, Ireland, September 1996.

[kackenhoff1994a] Rolf Käckenhoff and Detlef Merten. MOSS as a multimedia-object server. In Ralf Steinmetz, editor, Multimedia: Advanced Teleservices and High-Speed Communication Architectures, Second International Workshop, IWACA ’94, volume 868 of Lecture Notes in Computer Science, pages 413–425, Heidelberg, Germany, September 1994. Springer-Verlag.

[kahn1988a] Kenneth M. Kahn and Mark S. Miller. Language design and open systems. Technical report, Knowledge Systems Area, Intelligent System Laboratory, Xerox Palo Alto Research Center, 1988. (Published in The Ecology of Computation, B. Huberman (ed.), North Holland, 1988).

[kalbarczyk1999a] Zbigniew T. Kalbarczyk, Saurabh Bagchi, Keith Whisnant, and Ravishankar K. Iyer. Chamelon: A software infrastructure for adaptive fault tolerance. IEEE Transactions on Parallel and Distributed Systems, 10(6), June 1999.

[kappner1992a] Thomas Käppner, Dietmar Hehman, and Ralf Steinmetz. An introduction to HeiMAT: The Heidelberg multimedia application toolkit. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 405–416, La Jolla, California, USA, November 1992. Springer-Verlag.

[kappner1994a] Thomas Käppner and Lars C. Wolf. Media scaling in distributed multimedia object services. In Ralf Steinmetz, editor, Multimedia: Advanced Teleservices and High-Speed Communication Architectures, Second International Workshop, IWACA ’94, volume 868 of Lecture Notes in Computer Science, pages 34–43, Heidelberg, Germany, September 1994. Springer-Verlag.

[kappner1995a] Thomas Käppner and Ralf Steinmetz. Computational models for multimedia applications. In Jan van Leeuwen, editor, Computer Science Today — Recent Trends and Developments, volume 1000 of Lecture Notes in Computer Science, pages 592– 607. Springer-Verlag, 1995.

[karlsen2003b] Randi Karlsen. An adaptive transactional system – framework and service synchronization. In The International Symposium on Distributed Objects and Applications, DOA, Catania, Sicily, November 2003.

[karlsson1991a] Even-André Karlsson, Sivert Sørumgård, and Eirik Tryggeseth. Classification of object-oriented components for reuse. Technical report, Division of Computer Systems and Telematics, Norwegian Institute of Technology, October 1991.

[katiyar1994a] Dinesh Katiyar, David Luckham, John Mitchell, and Sigurd Meldal. Polymorphism and subtyping in interfaces. In Jeannette M. Wing, editor, Proceeding of the Workshop on Interfaces Definition Languages, volume 29/8 of Sigplan Notices, pages 22–34. ACM Press, August 1994.

[katz1994a] Randy H. Katz. Adaptation and mobility in wireless information systems. IEEE Personal Communications, 1(1):6–17, First Quarter 1994.

[kawachiya1995a] Kiyokuni Kawachiya, Masanobu Ogata, Nobuhiko Nishio, and Hideyuki Tokuda. Evaluation of QOS-control servers on Real-Time Mach. Technical report, IBM Research, Tokyo Research Laboratory, 1995.

[keller1994a] Ralf Keller, Stefan Fischer, and Wolfgang Effelsberg. Implementing movie control, access and managment — from a formal description to a working multimedia system. In Proceedings of the 14th International Conference on Distributed Computing Systems, pages 276–283, Poznan, Poland, June 1994. IEEE Computer Society Press.

[kemp1995a] Zarine Kemp. Multimedia and spatial information systems. IEEE Multimedia, 2(1):68–76, Spring 1995.

[kerherve1995a] B. Kerherve, A. Vogel, G. von Bochmann, R. Dssouli, J. Gecsei, and A. Hafid. On distributed multimedia presentational applications: Functional and computational architecture and QoS negotiation. In Gerald Neufield and Mabo Ito, editors, Protocols for High Speed Networks IV, pages 21–37. IFIP, Chapman & Hall, 1995.

[kermarrec1998a] Anne-Marie Kermarrec, Ihor Kuz, and Maarten van Steen. A framework for consistent, replicated web objects. In Proceeding of the 18th International Conference on Distributed Computing Systems, Amsterdam, The Netherlands, May 1998.

[keshav1994a] S. Keshav and H. Saran. Semantics and implementation of a native-mode ATM protocol stack. Technical memorandum, AT&T Bell Laboratories, February 1994.

[khalaila1996a] Ahmad H. Khalaila, Frank Eliassen, and Catriel Beeri. Efficient bill-of-materials algorithms. Technical report, Department of Computer Science, University of Tromsø, September 1996.

[khalaila1997a] Ahmad H. Khalaila and Frank Eliassen. A very efficient bill-of-materials algorithms. Technical report, Department of Computer Science, University of Tromsø, June 1997.

[kiczales1992a] Gregor Kiczales. Towards a new model of abstraction in the engineering of software. In Workshop on Reflection and Meta-level Architectures, IMSA’92, 1992.

[kiczales1993a] Gregor Kiczales, John Lamping, Chris Maeda, David Keppel, and Dylan Mc-Namee. The need for customizable operating systems. In Proceedings of the Fourth Workshop on Workstation Operating Systems, 1993.

[kiczales1993b] Gregor Kiczales, J. Michael Ashley, Luis Rodriguez, Amin Vahdat, and Daniel G. Bobrow. Metaobject protocols: Why we want them and what else they can do. In A. Paepcke, editor, Object-Oriented Programming: The CLOS Perpective, pages 101–118. The MIT Press, 1993.

[kiczales1996a] Gregor Kiczales. Beyond the black box: Open implementation. IEEE Software, January 1996.

[kiczales1997a] Gregor Kiczales, John Lamping, Anurag Mendhekar, Chris Maeda, Cristina Lopes, Jean-Marc Loingtier, and John Irwin. Aspect-oriented programming. In Mehmet Aksit and Satoshi Matsuoka, editors, ECOOP’97 — The 11th European Conference on Object-Oriented Programming, volume 1241 of Lecture Notes in Computer Science, pages 220–242, Jyväskylä, Finland, June 1997. Springer-Verlag.

[kiczales1997b] Gregor Kiczales, John Lamping, Christina Videira Lopes, Chris Maeda, Anurag Mendhekar, and Gail Murphy. Open implementation design guidelines. In Proceedings of the 1997 international conference on Software engineering, pages 481–490. ACM, 1997.

[kileng1994a] Frode Kileng. Mira-III redesign. Rapport IT3032/4-94, Norut IT, January 1994.

[kileng1994b] Frode Kileng. Mira-III, en itegrert telemedisinsk arbeidsstasjon. Prosjektbeskrivelse IT234/1-94, Norut IT, November 1994.

[killijian1998a] Marc-Olivier Killijian, Jean-Charles Fabre, and Juan-Carlos Ruiz-Garcia. A metaobject protocol for fault-tolerant CORBA applications. In Proceedings of 17th IEEE Symposium on Reliable Distributed Systems, pages 127–134, West Lafayette, USA, October 1998.

[king1992a] Steven S. King. Middleware! Data Communications, 21(4):58–67, March 1992.

[king1992b] Steven S. King. Message-delivery APIs: The message is the medium. Data Communications, 21(6):85–97, April 1992.

[kirtland1997a] Mary Kirtland. The COM+ programming model makes it easy to write components in any language. Technical report, Microsoft Corporation, November 1997.

[kistler1991a] James J. Kistler and M. Satyanarayanan. Disconnected operation in the Coda file system. In Proceedings of the 13th ACM Symposium on Operating System Principles, volume 10 of ACM Transactions on Computer Systems, October 1991.

[kitayama199?a] Takuro Kitayama, Tatsuo Nakajima, and Hideyuki Tokuda. RT-IPC: An IPC extension for Real-Time mach. Technical report, School of Computer Science, Carnegie Mellon University / Japan Advanced Institute of Science and Technology, 199?

[klas1995a] Wolfgang Klas and Karl Aberer. Multimedia applications and their implications on database architectures. In Advanced Course: Multimedia Databases in Perspective, pages 19–72, June 1995.

[kleiman1986a] S. R. Kleiman. Vnodes: An architecture for multiple file system types in Sun UNIX. In Proceedings of the USENIX 1986 Summer Conference, Atlanta, June 1986.

[kleinoder1997a] Jürgen Kleinöder and Michael Golm. MetaJava — a platform for adaptable operating-system mechanisms. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Object-Orientation and Operating Systems, volume 1357 of Lecture Notes in Computer Science, pages 507–514, Jyväskylä, Finland, June 1997. Springer-Verlag.

[klunder1988a] Doug Klunder. Hungarian naming conventions, January 1988.

[knightly1999a] Edward W. Knightly and Ness B. Shroff. Admission control for ststistical QoS: Theory and practice. IEEE Network, 13(2):20–29, March/April 1999.

[knuth1964a] Donald E. Knuth. Backus Normal Form vs. Backus Naur Form. Communications of the ACM, 7(12):735–736, December 1964.

[koblas1993a] Michelle R. Koblas. Support for the Eldorado transport protocol. Internal Report IT3032/2-93, Norut IT, December 1993.

[koegel1992a] John F. Koegel. On the design of multimedia interchange formats. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 262–271, La Jolla, California, USA, November 1992. Springer-Verlag.

[kohn1987a] Alfie Kohn. Studies find reward often no motivator — creativity and intrinsic interest diminish if task is done for gain. Boston Globe, January 1987.

[kon1998a] Fabio Kon, Ashish Singhai, Roy H. Campbell, Dulcineia Carvalho, Robert Moore, and Franscisco J. Ballesteros. 2K: A reflective, component-based operating system for rapidly changing environments. In ECOOP’98 Workshop on Reflective Object-Oriented Programming and Systems, Brussels, Belgium, July 1998.

[kon1999a] Fabio Kon and Roy H. Campbell. Supporting automatic configuration of component-based distributed systems. In 5th USENIX Conference on Object-Oriented Technologies and Systems, San Diego, CA, May 1999.

[kon1999b] Fabio Kon, Dulcineia Carvalho, and Roy H. Campbell. Automatic configuration in the 2K operating system. In 2nd ECOOP Workshop on Object-Orientation and Operating Systems, Lisbon, Portugal, June 1999.

[kon1999c] Fabio Kon and Roy H. Campbell. A framework for dynamically configurable multimedia distribution. In 9th ECOOP Workshop for PhD Students in Object Oriented Systems, Lisbon, Portugal, June 1999.

[kongsli1990a] Trond Arne Kongsli, Jan Børre Rydningen, and Svein Arne Solbakk. High speed data communication — an overview. Report R 0413, Forut, IT, September 1990.

[kongsli1993a] Trond Arne Kongsli. Telemedisinsk multimedia konferansesystem. Prosjektbeskrivelse IT3032/8-93, Norut IT, December 1993.

[koo1987a] Richard Koo and Sam Toueg. Checkpointing and rollback-recovery for distributed systems. IEEE Transactions on Software Engineering, SE-13(1):23–31, January 1987.

[kosaraju1973a] S. Rao Kosaraju. Limitations of Dijkstra’s semaphore primitives and Petri nets. Operating Systems Review, 7(4):122–126, 1973.

[koymans1991a] Ron Koymans. (Real) time: A philosophical perspective. In J. W. de Bakker, C. Huizing, W. P. de Roever, and G. Rozenberg, editors, Proceedings of the REX Workshop 1991, Real-Time: Theory in Practice, volume 600 of Lecture Notes in Computer Science, pages 353–370, Mook, The Netherlands, June 1991. Springer-Verlag.

[krakowiak1989a] S. Krakowiak, M. Meysembourg, H. Nguyen Van, M. Riveill, and C. Roisin. Design and implementation of an object-oriented, strongly typed language for distributed applications. Technical Report Guide-R8, Laboratoire de Génie Informatique, IMAG, and Centre de Recherche Bull, Grenoble, France, June 1989.

[kramp1999a] Thorsten Kramp and Rainer Koster. Flexible event-based threading for qossupporting middleware. In Proceedings of the Second IFIP WG 6.1 International Working Conference on Distributed Applications and Interoperable Systems, June/July 1999. (accepted).

[krishnamurthy1994a] A. Krishnamurthy and T. D. C. Little. Connection-oriented service renegotiation for scalable video delivery. In Proceedings of the 1st IEEE International Conference on Multimedia Computing and Systems, ICMCS’94, pages 502–507, Address, May 1994.

[kristensen1999a] Tom Kristensen and Thomas Plagemann. Extending the object request broker COOL with flexible QoS support. Technical report, Center for Technology at Kjeller, University of Oslo, 1999. Submitted IDMS’99.

[kuchling2000a] Andrew M. Kuchling. Python conference report. Linux Journal – Strictly On-Line, (73), May 2000.

[kupfer1993a] Michael D. Kupfer. Sprite on Mach. In Proceedings of the USENIX Mach III Symposium, pages 307–322, Santa Fe, 1993.

[kutvonen1998a] Lea Kutvonen. Architectures for distributed systems: Open distributed processing reference model. In HeCSE Workshop on Emerging Technologies in Distributed Systems, Lammi, Finland, January 1998.

[kvarnrud1994a] Pål Kværnrud. Autentisering av betaling. Technical report, TF, 1994.

[lakas1996a] A. Lakas, G. S. Blair, and A. Chetwynd. A formal approach to the design of QoS parameters in multimedia systems. In 4th International Workshop on Quality of Service, Paris, France, March 1996.

[lamming1992a] Mik Lamming and Sape J. Mullender. Security in distributed multimedia systems. (Draft), September 1992.

[lamport1974a] Leslie Lamport. A new solution of Dijkstra’s concurrent programming problem. Communications of the ACM, 17(8):453–455, August 1974.

[lamport1978a] Leslie Lamport. Time, clocks, and the ordering of events in a distributed system. Communications of the ACM, 21(7):558–565, July 1978.

[lamport1980a] Leslie Lamport. The ‘Hoare logic’ of concurrent programs. Acta Informatica, 14:21–37, 1980.

[lamport1982a] Leslie Lamport, Robert Shostak, and Marshall Pease. The byzantine general problem. ACM Transactions on Programming Languages and Systems, 4(3):382– 401, July 1982.

[lamport1983a] Leslie Lamport. Specifying concurrent program modules. ACM Transactions on Programming Languages and Systems, 5(2):190–222, April 1983.

[lamport1984a] Leslie Lamport. Solved problems, unsolved problems and non-problems in concurrency. In Proceedings of the 3rd PODC Conference. ACM, 1984.

[lamport1984b] Leslie Lamport. Using time instead of timeout for fault-tolerant distributed systems. ACM Transactions on Programming Languages and Systems, 6(2):254–280, April 1984.

[lamport1985a] Leslie Lamport and Fred B. Schneider. Formal foundation for specification and verification. In M. Paul and H. J. Siegert, editors, Distributed Systems, Methods and Tools for Specification — An Advanced Course, volume 190 of Lecture Notes in Computer Science, chapter 5, pages 103–285. Springer-Verlag, 1985.

[lamport1986a] Leslie Lamport. On interprocess communication — part i: Basic formalism. Distributed Computing, 1:77–85, 1986.

[lamport1986b] Leslie Lamport. On interprocess communication — part ii: Algorithms. Distributed Computing, 1:86–101, 1986.

[lamport1986c] Leslie Lamport. A simple approach to specifying concurrent systems. SRC Research Reports 15, DEC’s System Research Center, December 1986.

[lamport1986d] Leslie Lamport. The mutual exclusion problem: Part i—a theory of interprocess communication. Journal of the Association for Computing Machinery, 33(2):313– 326, April 1986.

[lamport1987a] Leslie Lamport. A formal basis for the specification of concurrent systems. In Y. Paker et al., editors, Distributed Operating Systems — Theory and Practice, volume F28, pages 4–46, Berlin, July 1987. NATO Advanced Study Institute, Springer-Verlag.

[lamport1988a] Leslie Lamport. Concurrent reading and writing of clocks. SRC Research Reports 27, DEC’s System Research Center, April 1988. (revised November 20, 1990).

[lamport1990a] Leslie Lamport. A theorem on atomicity in distributed algorithms. Distributed Computing, 4:59–68, 1990.

[lamport1991a] Leslie Lamport. The mutual exclusion problem has been solved. Communications of the ACM, 34(1), January 1991.

[lamport1991b] Leslie Lamport. The temporal logic of actions. SRC Research Reports 79, DEC’s System Research Center, December 1991.

[lamport1994b] Leslie Lamport. Introduction to TLA. Technical report, Digital Equipment Corporation, Systems Research Center, December 1994.

[lamport1995a] Leslie Lamport. Types are not harmless. Technical report, Digital Equipment Corporation, Systems Research Center, July 1995.

[lampson1983a] Butler W. Lampson. Hints for computer system design. In Proceedings of the 9th Symposium on Operating System Principles, volume 17 of System Review, pages 33–48, October 1983.

[lampson1984a] Butler W. Lampson. Hints for computer system design. IEEE Software, 1:11–28, January 1984.

[lampson1990a] Butler W. Lampson. Requirements and technology for computer science, July 1990.

[lampson1992a] Butler Lampson, Martín Abadi, Michael Burrows, and Edward Wobber. Authentication in distributed systems: Theory and practice. SRC Research Reports 83, DEC’s System Research Center, February 1992.

[landwehr1981a] Carl E. Landwehr. Formal models for computer security. ACM Computing Surveys, 13(3):247–278, September 1981.

[landwehr1997a] Carl E. Landwehr and David M. Goldschlag. Security issues in networks with internet access. Proceedings of the IEEE, 85(12), December 1997.

[langsford1994a] Alwyn Langsford. OSI management model and standards. In Morris Sloman, editor, Network and Distributed Systems Management, chapter 4, pages 69–93. Addison-Wesley, 1994.

[lann1977a] Gérard Le Lann. Distributed systems — towards a formal approach. In Bruce Gilchrists, editor, Information Processing 77 — Proceedings of IFIP Congress 77, volume 7 of IFIP Congress Series, Toronto, August 1977. North-Holland Publishing Company.

[lann1995a] G. Le Lann. On real-time and non real-time distributed computing. In Jean-Michel Hélary and Michel Raynal, editors, Distributed Algorithms, 9th International Workshop, WDAG ’95, volume 972 of Lecture Notes in Computer Science, pages 51–70, Le Mont-Saint-Michel, France, September 1995. Springer-Verlag.

[lanzillo1989a] Leo Lanzillo and Craig Partridge. Implementation of dial-up IP for UNIX systems. In Proceedings of the 1989 Winter USENIX Conference, pages 201–207, 1989.

[laroussinie1995a] François Laroussinie and Kim. G. Larsen. Compositional model checking of real-time systems. In Insup Lee and Scott A. Smolka, editors, CONCUR ’95: Concurrency Theory, 6th International Conference, volume 962 of Lecture Notes in Computer Science, pages 27–41, Philadelphia, PA, USA, August 1995. Springer-Verlag.

[lazar1992a] Aurel A. Lazar. A real-time management, control and information transport architecture for broadband networks. In Proceedings of the 1992 International Zurich Seminar on Digital Communications, Zurich, Switzerland, March 1992.

[lazar1994a] Aurel A. Lazar. Challenges in multimedia networking. In Proceedings of the International Hi-Tech Forum, Osaka’94, Osaka, Japan, February 1994.

[lazar1994b] Aurel A. Lazar, Giovanni Pacifici, and Dimitrios E. Pendarakis. Modeling video sources for real time scheduling. Multimedia Systems, 1(6):253–266, April 1994.

[lazar1994c] Aurel A. Lazar, Shailendra K. Bhonsle, and Koon Seng Lim. A binding architecture for multimedia networks. In Proceedings of the International COST 237 Workshop Multimedia Transport and Teleservices, pages 24–33, Vienna, Austria, November 1994.

[lazar1994d] A. A. Lazar. A research agenda for multimedia networking. In Workshop on Fundamentals and Perspectives on Multimedia Systems, International Conference Center for Computer Science, Dagstuhl Castle, Germany, July 1994.

[lazar1995a] A. A. Lazar, L. H. Ngoh, and A. Sahai. Multimedia networking abstractions with quality of service guarantees. In Proceedings of the SPIE Conference on Multimedia Computing and Networking, pages 24–33, San Jose, CA, February 1995.

[lea1993a] Rodger Lea, Christian Jacquemot, and Eric Pillevesse. COOL: system support for distributed object-oriented programming. Communications of the ACM, 36, September 1993.

[leboucher1995a] Laurent Leboucher and Jean-Bernard Stefani. Admission control for end-to-end distributed bindings. In D. Hutchison, H. Christiansen, G. Coulson, and A. Danthine, editors, Teleservices and Multimedia Communications — Second International COST 237 Workshop, volume 1052 of Lecture Notes in Computer Science, pages 192–208, Copenhagen, Denmark, November 1995. Springer-Verlag.

[lee1996a] Chen Lee, Ragunathan Rajkumar, and Cliff Mercer. Experiences with processor reservation and dynamic QOS in Real-Time Mach. In Proceedings of Multimedia Japan 96, March 1996.

[lee1996b] Chen Lee, Katsuhiko Yoshida, Cliff Mercer, and Ragunathan Rajkumar. Predictable communication protocol processing in Real-Time Mach. In Proceedings of IEEE Real-time Technology and Applications Symposium, June 1996.

[lee1999a] Elliot Lee. CORBA applications in GNOME. Technical report, Red Hat Software, July 1999.

[lee1999b] Elliot Lee. GNOME multimedia framework. Technical report, Red Hat Software, 1999.

[leopold1992a] Helmut Leopold, Andrew Campbell, David Hutchison, and N. Singer. Towards a integrated quality of service architecture (QOS-A) for distributed multimedia communications. In A. Danthine and O. Spaniol, editors, High Performance Networking, IV, Proceedings of the IFIP TC6/WG6.4 Fourth International Conference on High Performance Networking, IFIP Transactions, volume C-14, pages 169–182, Liège, Belgium, December 1992. Elsevier Science B. V.

[leopold1994a] Helmut Leopold, Kwaku Frimpong-Ansah, and Nikolaus Singer. From broadband network service to a distributed multimedia support environment. In D. Hutchison, A. Danthine, H. Leopold, and G. Coulson, editors, Multimedia Transport and Teleservice, International COST 237 Workshop, volume 882 of Lecture Notes in Computer Science, pages 47–68, Vienna, Austria, November 1994. Springer-Verlag.

[lerner2000a] Josh Lerner and Jean Tirole. The simple economics of open source. Technical report, Harvard Business School, February 2000.

[leslie1994a] Ian M. Leslie. Second year progress report. Pegasus paper 94-15, University of Cambridge Computer Laboratory, University of Twente Faculty of Computer Science, October 1994.

[leslie1996a] Ian Leslie, Derek McAuley, Richard Black, Timothy Roscoe, Paul Barham, David Evers, Robin Fairbairns, and Eoin Hyden. The design and implementation of an operating system to support distributed multimedia applications. Draft of JSAC paper, 1996.

[levin1983a] Roy Levin and David D. Redell. An evaluation of the ninth SOSP submissions or how (and how not) to write a good system paper. Operating Systems Review, 17(3):35–40, July 1983.

[levy1982a] Henry M. Levy and Peter H. Lipman. Virtual memory managment in the VAX/VMS operating system. Computer, pages 35–41, March 1982.

[leydekkers1995a] Peter Leydekkers, Valérie Gay, and Leonard Franken. A computational and engineering view on open distributed real-time multimedia exchange. Technical report, TINA, Core Team / Université Paris / PTT Research, 1995.

[li1994a] L. Li, A. Karmouch, and N. D. Georganas. Multimedia teleorchestra with independent sources: Part 1 — temporal modeling of collaborative multimedia senarios. Multimedia Systems, 1(4):143–153, February 1994.

[li1994b] L. Li, A. Karmouch, and N. D. Georganas. Multimedia teleorchestra with independent sources: Part 2 — synchronization algorithms. Multimedia Systems, 1(4):154–165, February 1994.

[li1994c] Guangxing Li. Distributed real-time objects: Some early experiences. External Paper APM.1231.01, ANSA, Architecture Projects Managment Limited, October 1994.

[li1994d] Guangxing Li. Some engineering aspects of real-time. Technical Report APM.1222.01, ANSA, Architecture Projects Managment Limited, October 1994.

[li1994e] Guangxing Li and Dave Otway. An open architecture for real-time processing. External Paper APM.1270.02, ANSA, Architecture Projects Managment Limited, October 1994.

[li1997a] Victor O. K. Li and Wanjiun Liao. Distributed multimedia systems. Proceedings of the IEEE, 85(7):1063–1108, July 1997.

[li1999a] Baochun Li and Klara Nahrstedt. A control-based middleware framework for quality of service adaptations. ieee-jsac, 1999.

[li1999b] Baochun Li and Klara Nahrstedt. Dynamic reconfiguration for complex multimedia applications. In Proceedings of IEEE International Conference on Multimedia Computing and Systems, Florence, Italy, June 1999.

[lima1997a] L. Paula Lima, Jr. and A. R. Cavalli. Test execution of telecommunications services using CORBA. In Howard Bowman and John Derrick, editors, IFIP TC6 WG6.1 International Workshop on Formal Methods for Open Object-based Distributed Systems (FMOODS ’97), volume 2, pages 409–419, Canterbury, Kent, UK, July 1997. International Federation for Information Processing (IFIP), Chapman & Hall.

[lincoln1990a] Patrick Lincoln, John Mitchell, Andre Scedrov, and Natarajan Shankar. Decision problems for propositional linear logic. In Proceedings of 31st Annual IEEE Symposium on Foundations of Computer Science, St. Louis, Missouri, October 1990.

[linden1993a] Rob van der Linden. An overview of ANSA. Architecture Report APM.1000.01, ANSA, Architecture Projects Managment Limited, July 1993.

[linden1998a] Ted Linden. Component-based development of complex applications. In Workshop on Compositional Software Architectures, Monterey, California, January 1998. (position paper).

[lindsey1995a] Donna Lindsey and Peter F. Linington. RIVUS: A stream template language for capturing multimedia requirements. In D. Hutchison, H. Christiansen, G. Coulson, and A. Danthine, editors, Teleservices and Multimedia Communications — Second International COST 237 Workshop, volume 1052 of Lecture Notes in Computer Science, pages 259–277, Copenhagen, Denmark, November 1995. Springer-Verlag.

[liskov1974a] Barbara Liskov and Stephen Zilles. Programming with abstract data types. Sigplan Notices, 9:50–59, 1974.

[liskov1985a] Barbara Liskov. The Argus language and system. In M. Paul and H. J. Siegert, editors, Distributed Systems, Methods and Tools for Specification — An Advanced Course, volume 190 of Lecture Notes in Computer Science, chapter 7, pages 343– 430. Springer-Verlag, 1985.

[liskov1986a] Barbara Liskov and William Weihl. Specifications of distributed programs. Distributed Computing, 1:102–118, 1986.

[liskov1987a] Barbara Liskov, Dorothy Curtis, Paul Johnson, and Robert Scheifler. Implementation of Argus. In Proceedings of the Eleventh ACM Symposium on Operating Systems Principles, pages 111–122, Austin, 1987.

[liskov1988a] Barbara Liskov. Distributed programming in Argus. Communications of the ACM, 31(3):300–312, March 1988.

[liskov1991a] Barbara Liskov. Practical use of synchronized clocks in distributed systems. In Proceedings of the Tenth Annual ACM Symposium on Principles of Distributed Computing, pages 1–9, Montreal, Canada, August 1991.

[little1991a] T. D. C. Little, A. Ghafoor, C. Y. R. Chen, C. S. Chang, and P. B. Berra. Multimedia synchronization. IEEE Data Engineering Bulletin, 14(3):26–35, September 1991.

[little1992a] T. D. C. Little and F. Kao. An intermedia skew control system for multimedia data presentation. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 130–141, La Jolla, California, USA, November 1992. Springer-Verlag.

[little1993a] T. D. C. Little, G. Ahanger, R. J. Folz, J. F. Gibbon, F. W. Reeve, D. H. Schelleng, and D. Venkatesh. A digital video-on-demand service supporting content-based queries. In Proceedings of ACM Multimedia’93, Anaheim CA, August 1993.

[little1993b] T. D. C. Little and D. Venkatesh. Probabilistic assignment of movies to storage devices in a video-on-demand system. In Proceedings of the 4th International Workshop on Network and Operating System Support for Digital Audio and Video, pages 213–224, Lancaster, UK, November 1993.

[little1993c] T. D. C. Little. A framework for synchronous delivery of time-dependent multimedia data. Multimedia Systems, 1(2):87–94, 1993.

[little1993d] T. D. C. Little and A. Ghafoor. Interval-based conceptual models for timedependent multimedia data. IEEE Transactions on Knowledge and Data Engineering, 5(4):551–563, August 1993.

[little1994a] T. D. C. Little and D. Venkatesh. Client-server metadata management for the delivery of movies in a video-on-demand system. In Proceedings of the 1st International Workshop on Services in Distributed and Networked Environments, SDNE, pages 11–18, Prague, Czech Republic, June 1994.

[little1994b] T. D. C. Little and D. Venkatesh. Prospects for interactive video-on-demand. IEEE Multimedia, 1(3):14–24, Fall 1994.

[little1994c] T. D. C. Little and D. Venkatesh. Popularity-based assignment of movies to storage devices in a video-on-demand system. Multimedia Systems, 1994.

[little1994d] T. D. C. Little, M. Ramanathan, and D. Venkatesh. Distributed storage and access of motion pictures via the virtual video browser. Technical report, Multimedia Communications Laboratory, Department of Electrical, Computer and Systems Engineering, Boston University, 1994.

[little1995a] T. D. C. Little. Multimedia synchronization is a process. Technical report, Multimedia Communications Laboratory, Department of Electrical, Computer and Systems Engineering, Boston University, May 1995.

[little1995b] T. D. C. Little and D. Venkatesh. The use of multimedia technology in distance learning. In Proceedings of IEEE International Conference on Multimedia and Networking, pages 3–17, Aizu-Wakamatsu, Japan, September 1995.

[lo1987a] Shau-Ping Lo and Virgil D. Gilgor. A comparative analysis of multiprocessor scheduling algorithms. In Proceedings of the 7th International Conference on Distributed Computing Systems, pages 356–363, Berlin, 1987.

[loepere1992a] Keith Loepere. Mach 3 Kernel Principles. Open Software Foundation Mach 3 Series. Open Software Foundation and Carnegie Mellon University, January 1992.

[long19??a] D. D. E. Long, J. L. Carroll, and C. J. Park. Estimating the reliability of hosts using the internet. Computer and Information Sciences, University of California and Mathematical Sciences, San Diego State University.

[lopes1997a] Cristina Videira Lopes and Gregor Kiczales. D: A language framework for distributed programming. Technical Report SPL97-010, P9710047, Xerox Palo Alto Research Center, February 1997.

[lorcy1997a] Stéphane Lorcy and Noël Plouzeau. A scheme for quality of service management in distributed applications. Publication interne 1134, IRISA, October 1997.

[lougher1992a] Phillip Lougher and Doug Shepherd. The design and implementation of a continuous media storage server. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 69–80, La Jolla, California, USA, November 1992. Springer-Verlag.

[lovett1990a] Hilde Lovett, Katrine Weisteen, Kjell Åge Bringsrud, Siri Jensen, Hanne Larsen, Arne Litlerè, Gunn Skogseth, and Jon Ølnes. En innføring i applikasjonslaget i OSI-modellen. Technical Report DTEK/05/90, Norsk Regnesentral, March 1990.

[loyall1999a] Joseph P. Loyall, Alia K. Atlas, Richard Schantz, Christopher D. Gill, David L. Levine, Carlos O’Ryan, and Douglas C. Schmidt. Flexible and adaptive control of real-time distributed object middleware. Technical report, Submitted to The International Journal of Time-Critical Computing Systems, Kluwer Academic Publishers, 1999.

[luckham1993a] David C. Luckham, James Vera, Dough Bryan, and Larry Augustin. Partial orderings of event sets and their application to prototyping concurrent, timed systems. The Journal of System and Software, 21:253–265, 1993.

[luckham1995a] David C. Luckham, James Vera, and Sigurd Meldal. Three concepts of system architecture. Technical Report CSL-TR-95-674, Computer Systems Laboratory, Stanford University, July 1995.

[luckham1995b] David C. Luckham, John L. Kenney, Larry M. Augustin, James Vera, Doug Bryan, and Walter Mann. Specification and analysis of system architecture using Rapide. IEEE Transactions on Software Engineering, 21(4):336–355, April 1995.

[luckham1995c] David C. Luckham and James Vera. An event-based architecture definition language. IEEE Transactions on Software Engineering, 21(9):717–734, September 1995.

[luckham1998a] David C. Luckham. Rapide: A language and toolset for causual event modelling of distributed system architectures. In Yoshifumi Masunaga, Takuya Katayama, and Michiharu Tsukamoto, editors, Worldwide Computing and Its Applications, volume 1368 of Lecture Notes in Computer Science, pages 88–96, Tsukuba, Japan, March 1998. Springer-Verlag.

[lunau1997a] Charlotte Pii Lunau. A reflective architecture for process control applications. In Mehmet Aksit and Satoshi Matsuoka, editors, ECOOP’97 — The 11th European Conference on Object-Oriented Programming, volume 1241 of Lecture Notes in Computer Science, pages 170–189, Jyväskylä, Finland, June 1997. Springer-Verlag.

[lynch1989a] Nancy A. Lynch and Mark R. Tuttle. An introduction to input/output automata. CWI Quarterly, 2(3):219–246, 1989.

[madelaine1993a] E. Madelaine and R. de Simone. FC2: Reference manual, version 1.1. Technical report, INRIA Sophia-Antipolis, 1993.

[maeda1997a] Chris Maeda, Arthur Lee, Gail Murphy, and Gregor Kiczales. Open implementation analysis and design. Technical report, Xerox PARC, 1997.

[maekawa1977a] Mamoru Maekawa. Interprocess communications in a highly diversified distributed system. In Bruce Gilchrists, editor, Information Processing 77 — Proceedings of IFIP Congress 77, volume 7 of IFIP Congress Series, Toronto, August 1977. North-Holland Publishing Company.

[maes1987a] Pattie Maes. Concepts and experiments in computational reflection. In Proceedings of OOPSLA ’87, volume 22 of Sigplan Notices, pages 147–155. ACM Press, October 1987.

[maes1988a] Pattie Maes. Issues in computational reflection. In Pattie Maes and Daniel Nardi, editors, Meta-Level Architectures and Reflection. North-Holland Publishing Company, 1988.

[malan1997a] G. Robert Malan, Farnam Jahanian, and Peter Knoop. Comparison of two middleware data dissemination services in a wide-area distributed system. In Proceedings of the 17th International Conference on Distributed Systems, Baltimore, Maryland, USA, May 1997. IEEE Computer Society Press.

[malenfant1996a] J. Malenfant, M. Jacques, and F.-N. Demers. A tutorial on behavioral reflection and its implementation. In Gregor Kiczales, editor, Proceedings of the Reflection’96 Conference, pages 1–20, April 1996.

[marcos1994a] Adérito F. Marcos. A distributed environment to support cooperative software development. In S. Fdida, editor, High Performance Networking, V — Proceedings of the IFIP TC6/WG6.4 Fifth International Conference on High Performance Networking, volume C-26 of IFIP Transactions, pages 375–390, Grenoble, France, 1994. North-Holland Publishing Company.

[martinez1998a] Lourdes Tajes Martínez, Fernando Álvarez García, Marián Díaz Fondón, Darío Álvarez Gutiérrez, and Juan Manuel Cueva Lovelle. A computational model for a distributed object-oriented operating system based on a reflective abstract machine. In ECOOP’98 Workshop on Reflective Object-Oriented Programming Systems, Brussels, Belgium, July 1998.

[marvie2000a] Raphaël Marvie, Philippe Merle, and Jean-Marc Geib. Towards a dynamic CORBA component platform. In DOA 2000 – 2nd International Symposium on Distributed Object Applications, Antwerp, Belgium, September 2000.

[mason1999a] David C. Mason. GNOME technologies: A brief description of the technology behind GNOME. Technical report, Red Hat Software, Inc., July 1999.

[masuhara1992a] Hidehiko Masuhara, Satoshi Matsuoka, Takuo Watanabe, and Akinori Yonezawa. Object-oriented concurrent reflective languages can be implemented efficiently. In OOPSLA ’92 Proceeding, volume 27(10) of Sigplan Notices, pages 127–144. ACM, ACM Press, October 1992.

[masuhara1996a] Hidehiko Masuhara, Satoshi Matsuoka, and Akinori Yonezawa. Implementing parallel language constructs using a reflective object-oriented language. In Gregor Kiczales, editor, Proceedings of the Reflection’96 Conference, pages 79–91, April 1996.

[masuhara1998a] Hidehiko Masuhara. Efficient implementation techniques of distributed/mobile reflective objects. In ECOOP’98 Workshop on Reflective Object-Oriented Programming Systems, Brussels, Belgium, July 1998.

[masuhara1998b] Hidehiko Masuhara and Akinori Yonezawa. Design and partial evaluation of meta-objects for a concurrent reflective language. In Eric Jul, editor, ECOOP’98 – Object-Oriented Programming, volume 1445 of Lecture Notes in Computer Science, pages 418–439, Brussels, Belgium, July 1998. Springer-Verlag.

[mathur199?a] Amit G. Mathur and Atul Prakash. Protocols for integrated audio and shared windows in collaborative systems. Technical report, Department of Electricla Enginnering and Computer Science, University of Michigan, 199?

[mathy1994a] Laurent Mathy and Olivier Bonaventure. QoS negotiation for multicast communication. In D. Hutchison, A. Danthine, H. Leopold, and G. Coulson, editors, Multimedia Transport and Teleservice, International COST 237 Workshop, volume 882 of Lecture Notes in Computer Science, pages 199–218, Vienna, Austria, November 1994. Springer-Verlag.

[mathy1995a] Laurent Mathy. On the design of a transport service to support multimedia applications. In F. Williams, H. Brä, and J. Nolan, editors, Proceedings of the Fourth International Conference Broadband Island ’95, Global Broadband & Beyond, pages 67–75, Dublin, Ireland, September 1995.

[mato-mira1993a] Fernando D. Mato Mira. ECLOS: An extended CLOS. In ECOOP 93 Workshop, Kaiserslautern, Germany, September 1993.

[mato-mira1995a] Fernando D. Mato Mira. Why the CLOS MOP should the mopped: The ECLOS metaclass library. In META’95 Workshop on Advances in Metaobject Protocols and Reflection (at ECOOP’95), Aarhus, Denmark, August 1995.

[matsuoka1991a] Satoshi Matsuoka, Takuo Watanabe, and Akinori Yonezawa. Hybrid group reflection architecture for object-oriented concurrent reflective programming. In Pierre America, editor, Proceedings of the European Conference on Object-Oriented Programming, ECOOP ’91, volume 512 of Lecture Notes in Computer Science, Geneva, Switzerland, July 1991. Springer-Verlag.

[mauthe1995a] Andreas Mauthe and Geoff Coulson. Scheduling and admission testing of jitter constrained periodic threads. In Proceedings of the 5th Internal Workshop on Network and Operating System Support for Digital Audio and Video, NOSSDAV 95, number MPG-95-04 in Lancaster University Report, Boston, 1995.

[mcaffer1996a] J. McAffer. Engineering the meta-level. In Gregor Kiczales, editor, Proceedings of the Reflection’96 Conference, pages 39–61, April 1996.

[mcaffer1996b] J. McAffer. Meta-level architecture support for distributed objects. In Gregor Kiczales, editor, Proceedings of Reflection’96, pages 39–62, San Francisco, 1996.

[mcauley1993a] Derek McAuley. Communication work package. Summary report, Esprit BRA 6865 — Pegasus, December 1993.

[mcauley1993b] D. R. McAuley. Operating system support for the desk area network. In D. Shepherd, G. Blair, G. Coulson, N. Davies, and F. Garcia, editors, Proceedings of the 4th International Workshop on Network and Operating System Support for Digital Audio and Video, volume 846 of Lecture Notes in Computer Science, pages 13–20, Lancaster, U.K., November 1993. Springer-Verlag.

[mccarthy1960a] John McCarthy. Recursive functions of symbolic expressions and their computation by machine, part i. Communications of the ACM, April 1960.

[mccarthy1980a] John McCarthy. Lisp—notes on its past and future—1980. In Lisp Conference, Stanford, 1980.

[mcclure1989a] Carma McClure. The CASE experience. Byte, pages 235–246, April 1989.

[mccluskey1997a] Glen McCluskey. Remote method invocation: Creating distributed java-to-java applications. Technical report, Sun Microsystems, Inc., October 1997.

[mckusick1984a] Marshall Kirk McKusick, William N. Joy, Samuel J. Leffler, and Robert S. Fabry. A fast file system for UNIX. ACM Transactions on Computer Systems, 2(3), August 1984.

[mcvoy1993a] Larry McVoy. The sourceware operating system proposal. Technical report, Sun Microsystems Computer Corporation, November 1993.

[meer1995a] Jan de Meer. Guest editorial: The ISO reference model for opend distributed processing. Computer Networks and ISDN Systems, 27(8):1211–1214, July 1995.

[meer1995b] J. de Meer and A. Hafid. QoS modelling of distributed teleoperating services. In Anne Clarke, Mario Campolargo, and Nikos Karatzas, editors, Bringing telecommunication services to the people — IS and N ’95: Third International Conference on Intelligence in Broadband Services and Networks, volume 998 of Lecture Notes in Computer Science, Heraklion, Crete, Greece, October 1995. Springer-Verlag.

[meland1988a] Truls Meland. Chaos i Tromsø: En forvreng hacker-historie fra virkeligheten. Datatid, (11):64–65, 1988.

[meldal1991a] Sigurd Meldal. A complete axiomatic semantics of spawning. Distributed Computing, 5:159–174, 1991.

[meldal1997a] Sigurd Meldal and David C. Luckham. Defining a security reference architecture. Technical Report CSL-TR-97-728, Computer Systems Laboratory, Stanford University, June 1997.

[meldal1997b] Sigurd Meldal and David C. Luckham. NSA’s MISSI reference architecture — moving from prose to precise specifications. In Manfred Broy and Bernhard Rumpe, editors, Requirements Targeting Software and System Engineering, volume 1526 of Lecture Notes in Computer Science, pages 293–329, Bernried, Germany, October 1997. Springer-Verlag.

[mendhekar1993a] Anurag Mendhekar and Daniel P. Friedman. Towards a theory of reflective programming languages. In Proceedings of the 1993 OOPSLA Workshop on Reflection and Meta-level Architectures, 1993.

[mendhekar1996a] Anurag Mendhekar and Daniel P. Friedman. An exploration of relationships between reflective theories. In Gregor Kiczales, editor, Proceedings of the Reflection’96 Conference, pages 245–262, April 1996.

[mens1997a] Kim Mens, Cristina Lopes, Bedir Tekinerdogan, and Gregor Kiczales. Aspectoriented programming workshop report. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Aspect-Oriented Programming, volume 1357 of Lecture Notes in Computer Science, pages 483–496, Jyväskylä, Finland, June 1997. Springer-Verlag.

[mercer1993a] Clifford W. Mercer, Stefan Savage, and Hideyuki Tokuda. Processor capacity reservers for multimedia operating systems. Technical Report CMU-CS-93-157, Carnegie Mellon University, School of Computer Science, May 1993.

[mercer1993b] Clifford W. Mercer, Ragunathan Rajkumar, and Hideyuki Tokuda. Applying hard real-time technology to multimedia systems. In Proceedings of the Workshop on the Role of Real-Time in Multimedia/Interactive Computing System, Raleigh-Durham, NC, November 1993.

[mercer1993c] Clifford W. Mercer, Stefan Savage, and Hideyuki Tokuda. Processor capacity reserves: An abstraction for managing processor usage. In Proceedings of the Fourth Workshop on Workstation Operating Systems, WWOS-IV, October 1993.

[mercer1994a] Clifford W. Mercer, Stefan Savage, and Hideyuki Tokuda. Processor capacity reservers: Operating system support for multimedia applications. In Proceedings of the IEEE International Conference on Multimedia Computing and Systems, May 1994.

[meyer-boudnik1995a] Thomas Meyer-Boudnik and Wolfgang Effelsberg. MHEG explained. IEEE Multimedia, 2(1):26–38, Spring 1995.

[meyer1999a] Bertrand Meyer. The significance of components. Software Development Online, November 1999.

[michael1993a] Maged M. Michael and Michael L. Scott. Fast mutual exclusion, even with contention. Technical report, Computer Science Department, University of Rochester, June 1993.

[michelitsch1997a] Georg Michelitsch, Max Ott, Daniel Reininger, and Girish Welling. QoS aware browsing in distributed multimedia systems. In Ralf Steinmetz and Lars C. Wolf, editors, IDMS’97 — 4th International Workshop on Interactive Distributed Multimedia Systems and Telecommunication Services, volume 1309 of Lecture Notes in Computer Science, Darmstadtm Germany, September 1997. Springer-Verlag.

[miller1997a] Robert Miller and Anand Tripathi. Issues with exception handling in object oriented systems. In Mehmet Aksit and Satoshi Matsuoka, editors, ECOOP’97 — The 11th European Conference on Object-Oriented Programming, volume 1241 of Lecture Notes in Computer Science, pages 85–103, Jyväskylä, Finland, June 1997. Springer-Verlag.

[millstein1999a] Todd Millstein and Craig Chambers. Modular statically typed multimethods. In Proceedings of the Thirteenth European Conference on Object-Oriented Programming (ECOOP 99), volume 1628 of Lecture Notes in Computer Science, pages 279–303, Lisbon, Portugal, June 1999. Springer-Verlag.

[milner1978a] Robin Milner. A theory of type polymorphism in programming. Journal of Computer and System Science, 17(3):348–375, 1978.

[milner1997a] Robin Milner. The Pi calculus and its applications. In Howard Bowman and John Derrick, editors, IFIP TC6 WG6.1 International Workshop on Formal Methods for Open Object-based Distributed Systems (FMOODS ’97), volume 2, pages 3–4, Canterbury, Kent, UK, July 1997. International Federation for Information Processing (IFIP), Chapman & Hall.

[milojicic1999a] Dejan Milojicic, David Black, Bill Bolosky, Frans Kaashoek, JochenLiedtke, Jeff Mogul, and John Wilkes. Trend wars: Operating systems—now and in the future. IEEE Concurrency, 7(1):12–21, January–March 1999.

[mitchell1996a] Scott Mitchell. Provision of support for high-level specification and dynamic configuration of multimedia groupware applications: A literature survey. Technical report, Department of Computer Science, Queen Mary & Westfield College, University of London, November 1996.

[mitchell1997a] S. E. Mitchell and R. J. Stroud. Reflective real-time object-oriented programming and systems, introduction to ECOOP’97 workshop #3. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Reflective Real-Time Object-Oriented Programming and Systems, volume 1357 of Lecture Notes in Computer Science, pages 51–54, Jyväskylä, Finland, June 1997. Springer-Verlag.

[mitchell1997b] S. E. Mitchell, A. Burns, and A. J. Wellings. Adaptive scheduling using reflection. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Reflective Real-Time Object-Oriented Programming and Systems, volume 1357 of Lecture Notes in Computer Science, pages 74–79, Jyväskylä, Finland, June 1997. Springer-Verlag.

[mitchell1998a] Scott Mitchell, Hani Naguib, George Coulouris, and Tim Kindberg. Dynamically reconfiguring multimedia components: A model-based approach. In Proceedings of SIGOPS EW98, the Eighth ACM SIGOPS European Workshop, pages 7–10, Sintra, Portugal, September 1998.

[moggi1991a] Eugenio Moggi. A category-theoretic account of program modules. Mathematical Structures in Computer Science, 1, 1991.

[monien1995a] Bukhard Monien, Ralf Diekmann, Rainer Feldmann, Ralf Klasing, Reinhard Lüling, Knut Menzel, Thomas Römke, and Ulf-Peter Schroeder. Efficient use of parallel & distributed systems: From theory to practice. In Jan van Leeuwen, editor, Computer Science Today — Recent Trends and Developments, volume 1000 of Lecture Notes in Computer Science, pages 62–77. Springer-Verlag, 1995.

[moran1992a] Mark Moran and Riccardo Gusella. System support for efficient dynamically configurable multi-party interactive multimedia applications. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 143–156, La Jolla, California, USA, November 1992. Springer-Verlag.

[moreira1994a] Ana M. D. Moreira and Robert G. Clark. Combining object-oriented analysis and formal description techniques. In Mario Tokoro and Remo Pareschi, editors, Proceedings of the Eighth European Conference on Object-Oriented Programming, volume 821 of Lecture Notes in Computer Science, pages 344–364, Bologna, Italy, July 1994. Springer-Verlag.

[morishima2000a] Naoto Morishima, Akimichi Ogawa, Keijiro Ehara, and Youki Kadobayashi. Field-trial of dynamic SLA in diffserv-capable network. In 1st International workshop on Quality of future Internet Services, Berlin, Germany, 2000.

[morris1977a] D. Morris, G. R. Frank, and C. J. Theaker. Machine-independent operating systems. In Bruce Gilchrists, editor, Information Processing 77 — Proceedings of IFIP Congress 77, volume 7 of IFIP Congress Series, Toronto, August 1977. North-Holland Publishing Company.

[morris1979a] Rober Morris and Ken Thompson. Password security: A case history. Communications of the ACM, 22(11):594–597, November 1979.

[morris1986a] James H. Morris, Mahadev Satyanarayanan, Micheal H. Conner, John H. Howard, David S. H. Rosenthal, and F. Donelson Smith. Andrew: A distributed personal computing environment. Communications of the ACM, 29(3):184–201, March 1986.

[moser1998a] L.E. Moser, P.M. Melliar-Smith, P. Narasimhan, V. Kalogeraki, and L. Tewksbury. The Eternal system. In Workshop on Compositional Software Architectures, Monterey, California, January 1998. (position paper).

[mowbray1998a] Thomas J. Mowbray. The trouble with dcom, complus, and corba — part 1 of 2. Object News, May 1998.

[mowbray1998b] Thomas J. Mowbray. The trouble with dcom, complus, and corba — part 2 of 2. Object News, May 1998.

[muhlhauser1996a] Max Mühlhäuser. Issues in multimedia software development. In Proceedings of the International Workshop on Multimedia Software Development, pages 2–7, Berlin, Germany, March 1996. IEEE Computer Society Press.

[mullender1991a] Sape J. Mullender. Experiences with Amoeba. In Distributed Open Systems in Perspective — Proceedings of the Spring 1991 EurOpen Conference, pages 1–11, Tromsø, Norway, May 1991. EurOpen.

[mullender1994a] Sape J. Mullender, Ian M. Leslie, and Derek McAuley. Operating-system support for distributed multimedia. Pegasus paper 94-6, University of Cambridge Computer Laboratory/University of Twente Faculty of Computer Science, June 1994.

[munch-ellingsen1990a] Arne Munch-Ellingsen. A functional description of the Basic Virtual Engine operating system. Technical report, Department of Computer Science, University of Tromsø, March 1990.

[munch-ellingsen1990b] Arne Munch-Ellingsen. The BVE MVME147 design and implementation. Technical report, Department of Computer Science, University of Tromsø, March 1990.

[murata199?a] Kenichi Murata, R. Nigel Horspool, Eric G. Manning, Yasuhiko Yokote, and Mario Tokoro. Unification of active and passive objects in an object-oriented operating system. Technical report, Keio University, Japan, University of Victoria, Canada, Sony Computer Science Laboratory, Japan, 199?

[murata199?b] Kenichi Murata, R. Nigel Horspool, Eric G. Manning, Yasuhiko Yokote, and Mario Tokoro. Unification of compile-time and run-time metaobject protocols. Technical report, Keio University, Japan, University of Victoria, Canada, Sony Computer Science Laboratory, Japan, 199?

[murphy1994a] David Murphy. Two papers concerning categories in concurrency theory. Report 5/94, School of Cognitive and Computing Sciences, University of Sussex, April 1994.

[myrvang2002a] Per Harald Myrvang and Tage Stabell-Kulø. Real time cryptographic protocol analysis. Technical report, Department of Computer Science, University of Tromsø, February 2002. (draft).

[naghshineh1997a] Mahmoud Naghshineh and Marc Willebeek-LeMair. End-to-end QoS provisioning in multimedia wireless/mobile networks using an adaptive framework. IEEE Communications Magazine, 35(11):72–81, November 1997.

[naguib1998a] Hani Naguib, Scott Mitchell, Tim Kindberg, and George Coulouris. QoS support in a model-based framework for distributed applications. In Poster presentation at Middleware 98, The Lake District, England, September 1998.

[nahrstedt1994a] Klara Nahrstedt and Jonathan Smith. A service kernel for multimedia endstations. In Ralf Steinmetz, editor, Multimedia: Advanced Teleservices and High-Speed Communication Architectures, Second International Workshop, IWACA’94, volume 868 of Lecture Notes in Computer Science, pages 8–22, Heidelberg, Germany, September 1994. Springer-Verlag.

[nahrstedt1995a] Klara Nahrstedt and Jonathan M. Smith. The QOS broker. IEEE Multimedia, 2(1):53–67, Spring 1995.

[nahrstedt1999a] Klara Nahrstedt, Hao hua Chu, and Srinivas Narayan. QoS-aware resource management for distributed multimedia applications. Journal of High-Speed Networking, 1999.

[najm1997a] Elie Najm and Jean-Bernard Stefani. Computational models for open distributed systems. In Howard Bowman and John Derrick, editors, IFIP TC6 WG6.1 International Workshop on Formal Methods for Open Object-based Distributed Systems (FMOODS ’97), pages 157–176, Canterbury, Kent, UK, July 1997. International Federation for Information Processing (IFIP), Chapman & Hall.

[najm1997b] E. Najm and A. Nimour. A calculus of object bindings. In Howard Bowman and John Derrick, editors, IFIP TC6 WG6.1 International Workshop on Formal Methods for Open Object-based Distributed Systems (FMOODS ’97), volume 2, pages 5–20, Canterbury, Kent, UK, July 1997. International Federation for Information Processing (IFIP), Chapman & Hall.

[nakajima1993a] Tatsuo Nakajima, Takuro Kitayama, and Hideyuki Tokuda. Experiments with real-time servers in real-time servers in real-time Mach. In MACH III Symposium, Sante Fe, NM, April 1993. USENIX.

[nakajima1998a] Tatsuo Nakajima. Resource reservation fo adaptive QOS mapping in Real-Time Mach. In José Rolim, editor, Proceeding of the 10 IPPS/SPDP’98 Workshops: Workshop on Parallel and Distributed Real-Time Systems, volume 1388 of Lecture Notes in Computer Science, pages 1047–1056, Orlando, Florida, March/April 1998. Springer-Verlag.

[nakajima1998b] Tatsuo Nakajima, Hiroyuki Aizu, Masaru Kobayashi, and Kenji Shimamoto. Environment server: A system support for adaptive distributed applications. In Yoshifumi Masunaga, Takuya Katayama, and Michiharu Tsukamoto, editors, Worldwide Computing and Its Applications, volume 1368 of Lecture Notes in Computer Science, pages 142–157, Tsukuba, Japan, March 1998. Springer-Verlag.

[nakajima199?a] Tatsuo Nakajima and Hiroshi Tezuka. A continuos media application supporting dynamic QoS control on Real-Time Mach. Technical report, Japan Advanced Institute of Science and Technlogy, 199?

[namuye1996a] Silvester Namuye, David Hutchison, Andreas Mauthe, and Geoff Coulson. Group managment services for distributed multimedia applications. Lancaster University Report MPG-96-04, Computing Department, Lancaster University, 1996.

[naur1960a] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Communications of the ACM, 3:299–314, May 1960.

[nebb1998a] Robb D. Nebb. Coordination and composition: The two paradigms underlaying AOP? In ECOOP ’98 Workshop on Aspect-Oriented Programming, July 1998. Position Paper.

[needham1972a] R. M. Needham. Tuning the Titan operating system. In C. A. R. Hoare and R. H. Perrott, editors, Operating Systems Techniques, Proceeding of a Seminar Held at Queen’s University, pages 277–281, Belfast, 1972. Academic Press.

[needham1992a] Roger Needham and Akira Nakamura. An approach to real-time scheduling — but i it really a problem for multimedia? In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 32–39, La Jolla, California, USA, November 1992. Springer-Verlag.

[neiger19??a] Gil Neiger and Sam Toueg. Providing design abstractions in distributed systems. School of Information and Computer Science, Georgia Institute of Technology and Department of Computer Science, Cornell University.

[nelson1990a] Victor P. Nelson. Fault-tolerant computing: Fundamental concepts. Computer, pages 19–25, July 1990.

[nelson1997a] Jeff Nelson. CorbScript: Using JavaScript with CORBA to clue your large components together. Java Developer’s Journal, 3(3):64–67, 1997.

[neuman19??a] B. Clifford Neuman and Jennifer G. Steiner. Authentication of unknown entities on an insecure network of untrusted workstations. Project Athena, Massachusetts Institute of Technology.

[neumann1997a] Luc Neumann and Alberto B. Raposo. An approach for an adaptive vizualization in a mobile environment. In Ralf Steinmetz and Lars C. Wolf, editors, IDMS’97 — 4th International Workshop on Interactive Distributed Multimedia Systems and Telecommunication Services, volume 1309 of Lecture Notes in Computer Science, Darmstadtm Germany, September 1997. Springer-Verlag.

[new1992a] Darren New, Jonathan Rosenberg, Gil Cruz, and Tom Judd. Requirements for network delivery of stored interactive multimedia. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 157–163, La Jolla, California, USA, November 1992. Springer-Verlag.

[newcomb1995a] Steven R. Newcomb. Multimedia interchange using SGML/HyTime, part i: Structures. IEEE Multimedia, 2(2):86–89, Summer 1995.

[newcomb1995b] Steven R. Newcomb. Multimedia interchange using SGML/HyTime, part ii: Principles and applications. IEEE Multimedia, 2(3):60–64, Fall 1995.

[ni1993a] Y. Ni and A. Goscinski. Resource and service trading in a heterogeneous large distributed system. In Proceedings of the IEEE Workshop on Advances in Parallel and Distributed Systems, Princeton, New Jersey, October 1993.

[nickerson1999a] Doug Nickerson. Components. Computer Bits, 9(10), October 1999.

[nicol1989a] John R. Nicol, Gordon S. Blair, Javad Malik, and Jonathan Walpole. Cosmos: An architecture for a distributed programming environment. Computer Communications, 12(3):147–157, June 1989.

[nicol1993a] John R. Nicol, C. Thomas Wilkes, and Frank A. Manola. Object orientation in hetrogeneous distributed computing system. Computer, 26(6):57–67, June 1993.

[nieh1993a] Jason Nieh, James G. Hanko, J. Duane Northcutt, and Gerard A. Wall. SVR4 UNIX scheduler unacceptable for multimedia applications. In D. Shepherd, G. Blair, G. Coulson, N. Davies, and F. Garcia, editors, Proceedings of the 4th International Workshop on Network and Operating System Support for Digital Audio and Video, volume 846 of Lecture Notes in Computer Science, pages 41–53, Lancaster, U.K., November 1993. Springer-Verlag.

[nishimura1977a] Toshio Nishimura. Formalization of concurrent processes. In Bruce Gilchrists, editor, Information Processing 77 — Proceedings of IFIP Congress 77, volume 7 of IFIP Congress Series, Toronto, August 1977. North-Holland Publishing Company.

[nitto1999a] Elisabetta Di Nitto and David S. Rosenblum. On the role of style in selecting middleware and underwear. In Engineering Distributed Objects Workshop, ICSE’99, Los Angeles, May 1999. Position Paper.

[nygaard1997a] Kristen Nygaard. GOODS to appear on the stage. In Mehmet Aksit and Satoshi Matsuoka, editors, ECOOP’97 — The 11th European Conference on Object-Oriented Programming, volume 1241 of Lecture Notes in Computer Science, pages 1–31, Jyväskylä, Finland, June 1997. Springer-Verlag.

[oberon1998a] Oberon microsystems. JBed whitepaper: Component software and real-time computing. Whitepaper, Oberon microsystems, September 1998.

[okamura1992a] H. Okamura, Y. Ishikawa, and M. Tokoro. AL-1/D: A distributed programming system with multi-model reflection framework. In Proceedings of the Workshop on New Models for Software Architecture, November 1992.

[oliver1997a] Huw Oliver, Mark Banfield, and David Hutchison. Distributed connection management for real-time multimedia services. In André Danthine and Christophe Diot, editors, From Multimedia Services to Network Services, Proceedings of the 4th International COST 237 Workshop, volume 1356 of Lecture Notes in Computer Science, pages 59–74, Lisboa, Portugal, December 1997. Springer-Verlag.

[oliver1998a] Huw Oliver, Christopher Edwards, Frederic Dang Tran, Jean-Bernard Stefani, and David Hutchison. Supporting real-time multimedia applications with distributed object controlled networks. In Proceedings of the 1st IEEE International Symposium on Object-Oriented Real-Time Distributed Computing, Kyoto, Japan, April 1998.

[omg2001a] Object Managment Group. The common object request broker: Architecture and specification. Technical report, Object Managment Group, February 2001. (revision 2.4.2).

[orgun19??a] Mehmet A. Orgun. An overview of temporal and modal logic programming. Department of Computing, Macquarie University and Computer Sciences Laboratory, RSISE, The Australian National University.

[orr1989a] Ken Orr, Chris Gane, Edward Yourdon, Peter P. Chen, and Larry L. Constantine. Methodology: The experts speak. Byte, pages 221–233, April 1989.

[orwell1946a] George Orwell. Politics and the english language, 1946.

[osf1991a] Open Software Foundation. The OSF/1 operating system. In Distributed Open Systems in Perspective — Proceedings of the Spring 1991 EurOpen Conference, pages 33–41, Tromsø, Norway, May 1991. EurOpen.

[ossher1999a] H. Ossher and P. Tarr. Multi-dimensional separation of concerns using hyperspaces. IBM Research Report 21452, IBM, April 1999.

[ott1996a] M. Ott, D. Reininger, G. Michelitsch, V. Bansal, R. J. Siracusa, and D. Raychaudhuri. Heidi-II: A software architecture for ATM network based on distributed multimedia systems. In Berthold Butscher, Eckhard Moeller, and Herwart Pusch, editors, Interactive Distributed Multimedia Systems and Services — European Workshop IDMS’96, volume 1045 of Lecture Notes in Computer Science, pages 1–14, Berlin, Germany, March 1996. Springer-Verlag.

[otway1995a] Dave Otway. Streams and signals. Architecture Report APM.1393.01, ANSA, Architecture Projects Managment Limited, January 1995.

[ousterhout1990a] John K. Ousterhout. Tcl: An embeddable command language. In Proceedings of the 1990 Winter USENIX Conference, 1990.

[ousterhout1991a] John K. Ousterhout. An X11 toolkit based on the Tcl language. In Proceedings of the 1991 Winter USENIX Conference, 1991.

[oxford1991a] Oxford University Timed CSP Group. Timed CSP: Theory and practice. In J. W. de Bakker, C. Huizing, W. P. de Roever, and G. Rozenberg, editors, Proceedings of the REX Workshop 1991, Real-Time: Theory in Practice, volume 600 of Lecture Notes in Computer Science, pages 640–675, Mook, The Netherlands, June 1991. Springer-Verlag.

[page1988a] Bob Page. A report in the Internet worm. Report, Computer Science Department, University of Lowell, November 1988.

[pan1995a] Davis Pan. A totorial on MPEG/audio compression. IEEE Multimedia, 2(2):60–74, Summer 1995.

[pancake1990a] Cherri M. Pancake and Donna Bergmark. Do parallel languages respond to the needs of scientific programmers? Computer, pages 13–23, December 1990.

[panzieri1994a] Fabio Panzieri and Marco Roccetti. A scalable architecture for reliable distributed multimedia applications. In Proceedings of the 14th International Conference on Distributed Computing Systems, pages 284–293, Poznan, Poland, June 1994. IEEE Computer Society Press.

[papathomas1994a] M. Papathomas, G. S. Blair, and G. Coulson. A model for active object coordination and its use for distributed multimedia applications. In Proceedings of ECOOP’94, Workshop on Coordination Models and Languages for Distribution and Parallelism, volume 924 of Lecture Notes in Computer Science. Springer-Verlag, 1994.

[papathomas1995a] M. Papathomas, G. S. Blair, G. Coulson, and P. Robin. Addressing the realtime synchronisation requirements of multimedia in an object-oriented framework. In Multimedia Computing and Networking, volume 2417 of IST&T/SPIE Proceedings, San Jose, USA, 1995.

[papathomas1996a] Michael Papathomas. ATOM: An active object model for enhancing reuse in the development of concurrent software. Research Report RR 963-I-LSR-2, LSR-IMAG, Grenoble, France, November 1996.

[papathomas1996b] Michael Papathomas. An active object model and its implementation in an open prototyping framework. Technical report, Bull-IMAG Systèmes, 1996.

[papathomas1996c] Michael Papathomas. Inheritance in concurrent object-oriented programming: Doing more with ATOM. Technical report, LSR-IMAG, Grenoble, France, 1996.

[papathomas1997a] Michael Papathomas. Using reflection in middleware platforms. Technical report, Computing Department, Lancaster University, June 1997. Draft.

[papathomas1997c] Michael Papathomas, Juan Hernandez, Juan Manual Murillo, and Fernando Sanchez. Inheritance and expressive power in concurrent object-oriented programming. In Roland Ducournau and Serge Garlatti, editors, Proceedings of Langages et Mod\les, Objects ’97, pages 45–60, Roscoff, France, 1997.

[park1997a] Jun-Cheol Park and Raymond E. Miller. Synthesizeing protocol specifications from service specifications in timed extended finite state machines. In Proceedings of the 17th International Conference on Distributed Systems, Baltimore, Maryland, USA, May 1997. IEEE Computer Society Press.

[parlavantzas2000a] Nikos Parlavantzas, Geoff Coulson, Mike Clarke, and Gordon Blair. Towards a reflective component based middleware architecture. In ECOOP’2000 Workshop on Reflection and Metalevel Architectures, Sophia Antipolis and Cannes, France, June 2000.

[parnas1977a] Davis L. Parnas. The use of precise specifications in the development of software. In Bruce Gilchrists, editor, Information Processing 77 — Proceedings of IFIP Congress 77, volume 7 of IFIP Congress Series, Toronto, August 1977. North-Holland Publishing Company.

[parnes1997a] Peter Parnes, Kåre Synnes, and Dick Schefstroöm. mTunnel: A multicast tunneling system with a user-based quality-of-service model. In Ralf Steinmetz and Lars C. Wolf, editors, IDMS’97 — 4th International Workshop on Interactive Distributed Multimedia Systems and Telecommunication Services, volume 1309 of Lecture Notes in Computer Science, Darmstadtm Germany, September 1997. Springer-Verlag.

[parris1994a] Colin Parris, Hui Zhang, and Domenico Ferrari. Dynamic managment of guaranteed-performance multimedia connections. Multimedia Systems, 1(6):267– 283, April 1994.

[pasquale1992a] Joseph C. Pasquale, Georg C. Polyzos, Eric W. Anderson, and Vachaspathi P. Kompella. The multimedia multicast channel. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 197–208, La Jolla, California, USA, November 1992. Springer-Verlag.

[pawlak1998a] R. Pawlak, L. Duchien, L. Seinturier, P. Champagnoux, D. Enselme, and G. Florin. A reflective implementation of a distributed programming model. In ECOOP’98 Workshop on Reflective Object-Oriented Programming Systems, Brussels, Belgium, July 1998.

[pawlan2001a] Monica Pawlan. Introduction to the J2EE platform. Technical report, Sun Microsystems, Inc., March 2001.

[payette1998a] Sandra Payette and Carl Lagoze. Flexible and extensible digital object and repository architecture (FEDORA). In Second European Conference on Research and Advanced Technology for Digital Libraries, September 1998.

[peleska1991a] Jan Peleska. Design and verification af fault tolerant systems with CSP. Distributed Computing, 5:95–106, 1991.

[pereira1994a] Carlos E. Pereira. Real-time active objects in C++/Real-Time UNIX. In Proceedings of the 1994 ACM SIGPLAN Workshop on Languages, Compilers and Tools for Real-Time Systems, June 1994.

[perez-luque1995a] M. J. Pérez-Luque and T. D. C. Little. A temporal reference framework for multimedia synchronization. In Proceedings of the International Workshop Multimedia Synchronization, Tysons Corner, Virginia, May 1995.

[peschanski1999a] Frédéric Peschanski. A component-based reflective architecture for concurrent and distributed programming. In Walter Cazzola, Robert J. Stroud, and Francesco Tisato, editors, OOPSLA’99 Workshop on Object Oriented Reflection and Software Engineering, pages 77–84, Denver, Colorado, US, November 1999.

[peterson1990a] Larry Peterson, Norman Hutchinson, Sean O’Malley, and Herman Rao. The xkernel: A platform for accessing internet resources. Computer, pages 23–33, May 1990.

[pfeifer1993a] Tom Pfeifer. Micronet machines — new architectural approaches for multimedia end-systems. In D. Shepherd, G. Blair, G. Coulson, N. Davies, and F. Garcia, editors, Proceedings of the 4th International Workshop on Network and Operating System Support for Digital Audio and Video, volume 846 of Lecture Notes in Computer Science, pages 29–40, Lancaster, U.K., November 1993. Springer-Verlag.

[pike1990a] Rob Pike, Dave Presotto, Ken Thompson, and Howard Trickey. Plan 9 from Bell Labs. In Proceedings of the UKUUG Summer ’90 Conference, pages 1–9, London, July 1990. UKUUG.

[pike1991a] Rob Pike, Dave Presotto, Ken Thompson, and Howard Trickey. Designing Plan 9. Dr. Dobb’s Journal, pages 49–60, January 1991.

[pike1991b] Rob Pike, Dave Presotto, Ken Thompson, Howard Trickey, Tom Duff, and Gerard Holzmann. Plan 9: the early papers. Technical report, AT&T Bell Laboratories, Murray Hill, New Jersey 07974, 1991.

[pinto1993a] Paulo F. Pinto and Peter F. Linington. A language for the specification of interactive and distributed multimedia applications. In J. De Meer, B. Mahr, and S. Storp, editors, Open Distributed Processing II, Proceedings of the IFIP TC6/WG6.1 International Conference on Distributed Processing, IFIP Transactions, volume C-20, pages 247–264, Berlin, Germany, September 1993. Elsevier Science B. V.

[plagemann1993a] Thomas Plagemann, Bernhard Plattner, Martin Vogt, and Thomas Walter. Modules as building blocks for protocol configuration. In Proceedings International Conference on Network Protocols, ICNP’93, pages 106–115, San Francisco, USA, October 1993.

[plagemann1994a] Thomas Plagemann, Janusz Waclawczyk, and Bernhard Plattner. Management of configurable protocols for multimedia. In Proceedings of ISMM International Conference Distributed Multimedia Systems and Applications, pages 78–81, Honolulu, USA, August 1994.

[plagemann1994b] Thomas Plagemann and Bernhard Plattner. Evaluating crucial performance issues of protocol configuration in Da CaPo. In Proceedings of First Workshop on High Performance Protocol Architecture, HIPPARCH’94, Sophia Antipolis, France, December 1994.

[plagemann1994c] Thomas Plagemann, Andreas Gotti, and Bernhard Plattner. CoRA — a heuristic for protocol configuration and resource allocation. In Gerald Neufield and Mabo Ito, editors, Workshop on Protocols for High Speed Networks IV, pages 85–102, Vancouver, Canada, August 1994. IFIP, Chapman & Hall.

[plagemann1995a] Thomas Plagemann, Knut A. Sæthre, and Vera Goebel. Application requirements and qos negotiation in multimedia systems. In Proceedings of Second Workshop on Protocols for Multimedia Systems, Salzburg Austria, October 1995.

[plagemann1999a] Thomas Plagemann, Frank Eliassen, Vera Goebel, Tom Kristensen, and Hans Ole Rafaelsen. Adaptive qos aware binding of persistent multimedia objects. Technical report, Center for Technology at Kjeller, University of Oslo, Department of Informatics, University of Oslo and Department of Computer Science, Univeristy of Tromsø, 1999. Submitted DOA’99.

[postel1980a] Johnathan B. Postel. Internetwork protocol approaches. IEEE Transactions on Communications, COM-28(4):604–611, April 1980.

[pountain1989a] Dick Pountain. The X window system. Byte, pages 353–360, January 1989.

[pountain1994a] Dick Pountain. The Chorus microkernel. Byte, pages 131–136, January 1994.

[presotto1991a] Dave Presotto, Rob Pike, Ken Thompson, and Howard Trickey. Plan 9, a distributed system. In Distributed Open Systems in Perspective — Proceedings of the Spring 1991 EurOpen Conference, pages 43–50, Tromsø, Norway, May 1991. EurOpen.

[puder1995a] A. Puder, S. Markwitz, F. Gudermann, and K. Geihs. AI-based trading in open distributed environments. In Proceedings of International Conference on Open Distributed Processing (ICODP’95). Chapman & Hall, 1995.

[puder1995b] A. Puder, S. Markwitz, and F. Gudermann. Service trading using conceptual structures. In Gerard Ellis, Robert Levinson, William Rich, and John F. Sowa, editors, Proceedings of the 3rd International Conference on Conceptual Structures (ICCS’95): Applications, Implementation and Theory, volume 954 of Lecture Notes in AI, pages 59–73, Berlin, Germany, 1995. Springer-Verlag.

[puder1996a] A. Puder and K. Geihs. System support for knowledge-based trading in open service markets. In Proceedings of 7th European SIGOPS Workshop, Connemara, Ireland, 1996.

[puder1996b] A. Puder and C. Burger. New concepts for qualitative trader cooperation. In International Conference on Distributed Platforms (ICDP’96), 1996.

[pugh1990a] William Pugh. Skip lists: A probabilistic alternativ to balanced trees. Communications of the ACM, 33(6), June 1990.

[puntigam1997a] Franz Puntigam. Coordinating requirements expressed in types for active objects. In Mehmet Aksit and Satoshi Matsuoka, editors, ECOOP’97 — The 11th European Conference on Object-Oriented Programming, volume 1241 of Lecture Notes in Computer Science, pages 367–388, Jyväskylä, Finland, June 1997. Springer-Verlag.

[pyarali1999a] Irfan Pyarali, Carlos O’Ryan, Douglas Schmidt, Nanbor Wang, Vishal Kachroo, and Aniruddha Gokhale. Applying optimization principle patterns to design realtime ORBs. In USENIX Conference on OO Technologies and Systems, San Diego, CA, US, May 1999.

[qian1996a] Tin Qian, See-Mong Tan, and Roy Campbell. An integrated architecture for open distributed multimedia computing. In Proceedings of the International Workshop on Multimedia Software Development, pages 24–30, Berlin, Germany, March 1996. IEEE Computer Society Press.

[raghavan1994a] S. V. Raghavan, B. Prabhakaran, and Satish K. Tripathi. Quality of service negotiation for orchestrated, distributed multimedia presentations. In S. Fdida, editor, High Performance Networking, V — Proceedings of the IFIP TC6/WG6.4 Fifth International Conference on High Performance Networking, volume C-26 of IFIP Transactions, pages 219–240, Grenoble, France, 1994. North-Holland Publishing Company.

[rajan1995a] Sreeranga Rajan, P. Venkat Rangan, and Harrick M. Vin. A formal basis for structured multimedia collaborations. In Proceedings of the International Conference on Multimedia Computing and Systems, pages 194–201, Washington, D.C., May 1995. IEEE Computer Society Press.

[rajkumar1997a] Raj Rajkumar, Chen Lee, John Lehoczky, and Dan Siewiorek. A resource allocation model for QoS management. In Proceedings of the IEEE Real-Time Systems Symposium, December 1997.

[rajkumar1998a] Ragunathan (Raj) Rajkumar. System software for multimedia computing. In Yoshifumi Masunaga, Takuya Katayama, and Michiharu Tsukamoto, editors, Worldwide Computing and Its Applications, volume 1368 of Lecture Notes in Computer Science, pages 120–141, Tsukuba, Japan, March 1998. Springer-Verlag.

[ramakrishnan1995a] K. K. Ramakrishnan, Lev Vaitzblit, Cary Gray, Uresh Vahalia, Dennis Ting, Percy Tzelnic, Steve Glaser, and Wayne Duso. Operating system support for a video-on-demand file service. Multimedia Systems, 3(2):53–65, May 1995.

[ramsey1994a] Norman Ramsey. Literate programming simplified. IEEE Software, pages 97–105, September 1994.

[ranum1992a] Marcus J. Ranum. A network firewall. Technical report, Digital Equipment Corporation, Washington Open System Resource Center, June 1992.

[ranum19??a] Marcus J. Ranum. Thinking about firewalls. Trusted Information Systems, Inc.

[rao1991a] Ramana Rao. Implementational reflection in Silica. In Pierre America, editor, ECOOP ’91, European Conference on Object-Oriented Programming, volume 512 of Lecture Notes in Computer Science, Geneva, Switzerland, July 1991. Springer-Verlag.

[rashid1986a] Richard F. Rashid. From RIG to Accent to Mach: The evolution of a network operating system. In Proceedings of Fall Joint Computer Conference, pages 1128– 1137. ACM/IEEE Computer Society, 1986.

[ravindran1993a] K. Ravindran. Transport-oriented service models for synchronization of multimedia data streams. Technical Report TR:93-8, Department oc Computing and Information Sciences, Kansas State University, January 1993. Revised August 1995.

[raymond1993a] K. A. Raymond. Reference model of open distributed processing: A tutorial. In J. De Meer, B. Mahr, and S. Storp, editors, Open Distributed Processing II, Proceedings of the IFIP TC6/WG6.1 International Conference on Distributed Processing, IFIP Transactions, volume C-20, pages 3–14, Berlin, Germany, September 1993. Elsevier Science B. V.

[raymond1993b] Kerry Raymond. Reference model of open distributed processing (RM-ODP): Introduction. Technical report, CRC for Distributed System Technology, Centre for Information Technology Research, University of Queensland, 1993.

[raymond1998a] Eric S. Raymond. The cathedral and the bazar, May 1998.

[reeds1984a] J. A. Reeds and P. J. Weinberger. File security and the UNIX system crypt command. AT&T Bell Laboratories Technical Journal, 63(8):1673–1683, October 1984.

[rees1986a] Jim Rees, Paul H. Levine, Nathaniel Mishkin, and Paul J. Leach. An extensible i/o system. Technical report, Apollo Computer inc, 1986. (Proceedings of the Summer 1986 Usenix Conference, Atlanta, Georgia).

[rees1996a] Jonathan A. Rees. A security kernel based on the lambda-calculus. A.I. Memo No. 1564, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, March 1996.

[reid1986a] Brian Reid. Lessons from the UNIX breakins at Standford. Software Engineering Notes, 11(5):29–35, October 1986.

[reinhardt1994a] Wilko Reinhardt. Advance reservation of network resources for multimedia applications. In Ralf Steinmetz, editor, Multimedia: Advanced Teleservices and High-Speed Communication Architectures, Second International Workshop, IWACA ’94, volume 868 of Lecture Notes in Computer Science, pages 23–33, Heidelberg, Germany, September 1994. Springer-Verlag.

[reinhardt1995a] Wilko Reinhardt and Dirk Trossen. An enhanced admission control scheme for deterministic and predictive services. In D. Hutchison, H. Christiansen, G. Coulson, and A. Danthine, editors, Teleservices and Multimedia Communications — Second International COST 237 Workshop, volume 1052 of Lecture Notes in Computer Science, pages 209–228, Copenhagen, Denmark, November 1995. Springer-Verlag.

[reiter1992a] Michael Reiter, Kenneth Birman, and Li Gong. Integrating security in a group oriented distributed system. In Proceedings of the 1992 IEEE Symposium on Research in Security and Privacy, 1992.

[ren1995a] Shangping Ren and Gul A. Agha. RTsynchronizer: Language support for real-time specifications in distributed systems. In Proceedings of the 1995 ACM SIGPLAN Workshop on Languages, Compilers and Tools for Real-Time Systems, June 1995.

[ren1997a] Shangping Ren, Nalini Venkatasubramanian, and Gul A. Agha. Formalizing multimedia QoS constraints using actors. In Howard Bowman and John Derrick, editors, IFIP TC6 WG6.1 International Workshop on Formal Methods for Open Objectbased Distributed Systems (FMOODS ’97), volume 2, pages 139–153, Canterbury, Kent, UK, July 1997. International Federation for Information Processing (IFIP), Chapman & Hall.

[renesse1997a] Robbert van Renesse, Ken Birman, Mark Hayden, Alexey Vaysburd, and David Karr. Building adaptive systems using Ensemble. Technical report, Department of Computer Science, Cornell University, 1997.

[retina1999a] ReTINA. Extended DPE resource control framework specifications. ReTINA Deliverable AC048/D1.01xtn, France Telecom – CNET and GMD Fokus, January 1999. ACTS Project AC048, An Industrial-Quality TINA-Compliant Real-Time DPE.

[ricciardi1991a] Aleta M. Ricciardi and Kenneth P. Birman. Using process groups to implement failure detection in asynchronous environments. Technical report, Departement of Computer Science, Cornell University, February 1991.

[richard1993a] William D. Richard, Jerome R. Cox, Brian L. Gottlieb, and Ken Krieger. The Washington university multimedia system. Multimedia Systems, 1(3):120–131, 1993.

[ritchie1974a] Dennis M. Ritchie and Ken Thompson. The UNIX time sharing system. Communications of the ACM, 17(7):365–375, July 1974.

[ritchie1984a] Dennis M. Ritchie. The evolution of the unix time-sharing system. AT&T Bell Laboratories Technical Journal, 63(6 Part 2):1577–1593, October 1984.

[rizzo1997a] Luigi Rizzo. The FreeBSD audio driver. In André Danthine and Christophe Diot, editors, From Multimedia Services to Network Services, Proceedings of the 4th International COST 237 Workshop, volume 1356 of Lecture Notes in Computer Science, pages 90–104, Lisboa, Portugal, December 1997. Springer-Verlag.

[robin1995a] Philippe Robin, Geoff Coulson, Andrew Campbell, Gordon Blair, Michael Papathomas, and David Hutchison. Implementing a QoS controlled ATM based communication system in Chorus. In Gerald Neufield and Mabo Ito, editors, Protocols for High Speed Networks IV, pages 38–54. IFIP, Chapman & Hall, 1995.

[roman1999a] Manuel Roman, Fabio Kon, and Roy H. Campbell. Supporting dynamic reconfiguration in the dynamicTAO reflective ORB. Technical Report UIUCDCS-R-98-2085, Department of Computer Science, University of Illinois at Urbana-Champaign, February 1999.

[roman1999b] Manuel Roman, Fabio Kon, and Roy H. Campbell. Design and implementation of runtime reflection in communication middleware: the dynamicTAO case. In ICDCS’99 Workshop on Middleware, Austin, Texas, May 1999.

[roman1999d] Manuel Roman, Ashish Singhai, Dulcineia Carvalho, Christopher Hess, and Roy H. Campbell. Integrating PDAs into distributed systems: 2k and PalmORB. In International Symposium on Handheld and Ubiquitous Computing (HUC 99), Karlsruhe, Germany, September 1999.

[roman2000a] Manuel Roman, Fabio Kon, and Roy H. Campbell. LegORB and ubiquitous CORBA. In Workshop on Reflective Middleware at Middleware 2000, New York, USA, April 2000.

[roodyn1999a] Neil Roodyn. A software architecture for a real time data distributed objects system. In Engineering Distributed Objects Workshop, ICSE’99, Los Angeles, May 1999. Invited Presentation.

[rooholamini1995a] Reza Rooholamini and Vladimir Cherkassky. ATM-based multimedia servers. IEEE Multimedia, 2(1):39–52, Spring 1995.

[rose1994a] Gary P. Rose, Jeffery T. Oesterle, Joseph E. Kasper, and Robert J. Hammond. Development of a multimedia product for HP workstations. Hewlett-Packard Journal, 45(2), April 1994.

[rosenblum1991a] Mendel Rosenblum and John K. Ousterhout. The design and implementation of a log-structured file system. In Proceedings of the 13th Symposium on Operating System Principles. ACM, October 1991.

[rosenstein19??a] Mark A. Rosenstein, Daniel E. Geer, Jr., and Peter J. Levine. The Athena service managment system. Project Athena.

[rossum1997a] Guido van Rossum. Metaclasses in Python 1.5. Essay, Corporation for National Research Initiatives, 1997.

[rossum1998a] Guido van Rossum. Glue it all together with Python. In Workshop on Compositional Software Architectures, Monterey, California, January 1998. (position paper).

[rossum1998b] Guido van Rossum. Java and Python: a perfect couple. developer.com Journal, August 1998.

[roth1998a] Bill Roth. An introduction to enterprise JavaBeans technology. Technical report, Java Software, Sun Microsystems, October 1998.

[rothermel1996a] Kurt Rothermel, Gabriel Dermler, and Walter Fiederer. QoS negotiation and resource reservation for distributed multimedia applications. Fakultätsbericht 1996/09, Fakultät Informatik, Universität Stuttgart, July 1996.

[roti1994a] Steve Roti. Client/server security: A look at client/server security problems and some possible solutions. DBMS, 7(4), April 1994.

[rozier1988a] M. Rozier, V. Abrossimov, F. Armand, I. Boule, M. Gien, M. Guillemont, F. Herrmann, C. Kaiser, S. Langlois, P. Léonard, and W. Neuhauser. Chorus distributed operating system. Computing Systems, 1(4):305–370, 1988.

[rtmp1995a] Real-Time Mach Project. Real-Time Mach 3.0 user reference manual. Technical report, Department of Computer Science, Carnegie Mellon University, June 1995.

[rtmp1995b] Real-Time Mach Project. Rt-mach notes. Technical report, Department of Computer Science, Carnegie Mellon University, August 1995. (Version 1.0).

[rushby1981a] J. M. Rushby. Design and verification of secure systems. In Proceedings of the 8th Symposium on Operating System Principles. ACM, 1981.

[saikoski1999a] Katia B. Saikoski and Geoff Coulson. Adaptive groups in Open ORB. In Proceedigns of the CAiSE ’99 Doctoral Consortium, Heidelberg, Germany, June 1999.

[saltzer1984a] J. H. Saltzer, D. P. Reed, and D. D. Clark. End-to-end arguments in system design. ACM Transactions on Computer Systems, 2(4):277–288, November 1984.

[salus1994b] Peter H. Salus. Farewell to Berkley UNIX. Unix Review, pages 39–44, May 1994.

[salzmann1999a] Chris Salzmann. Managing shared business objects. In Engineering Distributed Objects Workshop, ICSE’99, Los Angeles, May 1999. Position Paper.

[sanchez1998a] Frenando Sánchez, Juan Hernández, Juan Manuel Murillo, and Enrique Pedraza. Run-time adaptability of synchronization policies in concurrent object-oriented languages. In ECOOP ’98 Workshop on Aspect-Oriented Programming, July 1998. Position Paper.

[satoh1992a] Ichiro Satoh and Mario Tokoro. A formalism for real-time concurrent objectoriented computing. In OOPSLA ’92 Proceeding, volume 27(10) of Sigplan Notices, pages 315–326. ACM, ACM Press, October 1992.

[satyanarayanan1989a] M. Satyanarayanan. Integrating security in a large distributed system. ACM Transactions on Computer Systems, 7(3):247–280, August 1989.

[satyanarayanan1990a] Mahadev Satyanarayanan and Ellen H. Siegel. Parallel communication in a large distributed environment. IEEE Transactions on Computers, 39(3):328–348, March 1990.

[satyanarayanan1990b] M. Satyanarayanan, James J. Kistler, Puneet Kumar, Maria E. Okasaki, Ellen H. Siegel, and David C. Steere. Coda: A highly available file system for a distributed workstation environment. IEEE Transactions on Computers, 39(4):447– 459, April 1990.

[satyanarayanan1990c] M. Satyanarayanan. A survey of distributed file systems. Annu. Rev. Computer Science, 4:73–104, 1990.

[satyanarayanan1990d] M. Satyanarayanan. Scalable, secure, and highly available distributed file access. Computer, pages 9–20, May 1990.

[scheifler1986a] Robert W. Scheifler. NeWS brakes. Technical report, MIT Laboratory for Computer Science, October 1986.

[schmidt1993a] Douglas C. Schmidt and Tatsuya Suda. Transport system architecture services for high-performance communications systems. IEEE Journal on Selected Areas in Communications, 11(4):489–506, May 1993.

[schneider1985a] Fred B. Schneider and Leslie Lamport. Paradigms for distributed programs. In M. Paul and H. J. Siegert, editors, Distributed Systems, Methods and Tools for Specification — An Advanced Course, volume 190 of Lecture Notes in Computer Science, chapter 8, pages 431–480. Springer-Verlag, 1985.

[schneider1991a] Fred B. Schneider Bard Bloom and Keith Marzullo. Putting time into proof outlines. In J. W. de Bakker, C. Huizing, W. P. de Roever, and G. Rozenberg, editors, Proceedings of the REX Workshop 1991, Real-Time: Theory in Practice, volume 600 of Lecture Notes in Computer Science, pages 618–639, Mook, The Netherlands, June 1991. Springer-Verlag.

[schneier1996a] Bruce Schneier. Why cryptography is harder than it looks. Technical report, Counterpane Systems, 1996.

[schonhage1999a] Bastiaan Schönhage and Anton Eliëns. From distributed object features to architectural styles. In Engineering Distributed Objects Workshop, ICSE’99, Los Angeles, May 1999. Position Paper.

[schreiber1995a] Richard Schreiber. Middleware demystified. Datamation, 41(6):41–45, April 1995.

[schreier1995a] Louis C. Schreier and Michael B. Davis. System-level resource managment for network-based multimedia applications. Technical report, SRI International, 1995.

[schroeder1984a] Michael D. Schroeder, Andrew D. Birell, and Roger M. Needham. Experience with Grapevine: The growth of a distributed system. ACM Transactions on Computer Systems, 2(1):3–23, February 1984.

[schubert1997a] Frank Schubert. A reflective architecture for an adaptable object-oriented operating system based on C++. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Object-Orientation and Operating Systems, volume 1357 of Lecture Notes in Computer Science, pages 515–522, Jyväskylä, Finland, June 1997. Springer-Verlag.

[scourias1996a] John Scourias. Overview of the global system for mobile communications. Technical report, University of Waterloo, 1996.

[seeley1988a] Donn Seeley. A tour of the worm. Technical report, Department of Computer Science, University of Utah, 1988. (Published in Proceedings of 1989 Winter Usenix Conference, San Diago, California, February 1989).

[seltzer1993a] Margo Seltzer, Keith Bostic, Marshall Kirk McKusick, and Carl Staelin. An implementation of a log-structured file system for UNIX. In Proceedings of the Winter 1993 USENIX Conference, pages 307–326, San Diego, California, January 1993.

[sha1993a] Lui Sha and Shirish S. Sathaye. A systematic approach to design distributed realtime systems. Computer, 26(9):68–78, September 1993.

[shahabi1997a] Cyrus Shahabi, Mohammad H. Alshayeji, and Shimeng Wang. A redundant hierarchical structure for a distributed continuous media server. In Ralf Steinmetz and Lars C. Wolf, editors, IDMS’97 — 4th International Workshop on Interactive Distributed Multimedia Systems and Telecommunication Services, volume 1309 of Lecture Notes in Computer Science, Darmstadtm Germany, September 1997. Springer-Verlag.

[shapiro1988a] Marc Shapiro. The design of a distributed object-oriented operating system for office applications. In ESPRIT ’88, Putting the Technology to Use, pages 1020–1027, 1988.

[shapiro1989a] Ehud Shapiro. The family of concurrent logic programming languages. ACM Computing Surveys, 21(3):413–510, September 1989.

[shapiro1990a] Marc Shapiro, David Plainfossé, and Oliver Gruber. A garbage detection protocol for a realistic distributed object-support system. Rapport de Recherche INRIA 1320, INRIA, November 1990.

[sharma1994a] R. Sharma and S. Keshav. Signaling and operating system support for nativemode ATM applications. In Proceedings of the 1994 SIGCOMM Conference, pages 149–157, London, UK, August 1994.

[shenoy1995a] Prashant J. Shenoy, Pagan Goyal, and Harrick M. Vin. Issues in multimedia server design. ACM Computing Surveys, 27(4):636–639, December 1995.

[sheu1997a] Guang-Way Sheu, Yue-Shan Chang, Deron Liang, Shyan-Ming Yuan, and Winston Lo. A fault tolerant object service on CORBA. In Proceedings of the 17th International Conference on Distributed Systems, Baltimore, Maryland, USA, May 1997. IEEE Computer Society Press.

[shrivastava1982a] S. K. Shrivastava and F. Panzieri. The design of a reliable remote procedure call mechanism. IEEE Transactions on Computers, C-31(7):692–697, July 1982.

[sikora1997a] T. Sikora. MPEG-1 and MPEG-2 digital video coding standards. In Ronald K. Jurgen, editor, Digital Consumer Electronics Handbook. McGraw-Hill Book Company, 1997.

[silberschatz1991b] A. Silberschatz, J. Peterson, and P. Galvin. The Mach operating system. In Operating System Concepts, chapter 16. Addison-Wesley, 3 edition, 1991.

[silberschatz1995a] Avi Silberschatz, Mike Stonebraker, and Jeff Ullman. Database research: Achievements and opportunities into the 21st century. In NSF Workshop on the Future of Database Systems Research, May 1995.

[silverman1983a] Jonathan M. Silverman. Reflections on the verification of the security of an operating system kernel. In Proceedings of the Ninth Symposium on Operating System Principles, October 1983. (in Operating System Review, vol. 17 no. 5, October 1993).

[simon1994a] Robert Simon, Taieb Znati, and Robert Sclabassi. Group communication in distributed multimedia systems. In Proceedings of the 14th International Conference on Distributed Computing Systems, pages 294–301, Poznan, Poland, June 1994. IEEE Computer Society Press.

[simoni1992a] Noëmie Simoni and Simon Znaty. QoS: From definition to managment. In A. Danthine and O. Spaniol, editors, High Performance Networking, IV, Proceedings of the IFIP TC6/WG6.4 Fourth International Conference on High Performance Networking, IFIP Transactions, volume C-14, pages 253–264, Liège, Belgium, December 1992. Elsevier Science B. V.

[singhai1997a] Ashish Singhai, Aamod Sane, and Roy H. Campbell. Reflective ORBs: Supporting robust, time-critical distribution. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Reflective Real-Time Object-Oriented Programming and Systems, volume 1357 of Lecture Notes in Computer Science, pages 55–61, Jyväskylä, Finland, June 1997. Springer-Verlag.

[singhai1998a] Ashish Singhai, Aamod Sane, and Roy H. Campbell. Quarterware for middleware. In 18th IEEE International Conference on Distributed Computing Systems, Amsterdam, The Netherlands, May 1998.

[singhal1989a] Mukesh Singhal. Deadlock detection in distributed systems. Computer, pages 37–48, November 1989.

[smeaton1994a] Alan F. Smeaton and Aidan Gregan. Distributed multimedia QoS parameters from presentation modelling by coloured petri nets. In Peter Brusilovsky, Piet Kommers, and Norbert Streitz, editors, Multimedia, Hypermedia, and Virtual Reality — First International Conference, MHVR’94, volume 1077 of Lecture Notes in Computer Science, pages 47–60, Moscow, Russia, September 1994. Springer-Verlag.

[smith1992a] Gerard J. M. Smith and Paul J. M. Havinga. The architecture of Rattlesnake: A real-time multimedia network. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 15–24, La Jolla, California, USA, November 1992. Springer-Verlag.

[sobel1996a] Jonathan M. Sobel and Daniel P. Friedman. An introduction to reflection-oriented programming. In Gregor Kiczales, editor, Proceedings of the Reflection’96 Conference, April 1996.

[soisalon-soininen1982a] Eljas Soisalon-Soininen. Inessential error entries an their use i LR parser optimization. ACM Transactions on Programming Languages and Systems, 4(2):179–195, April 1982.

[solomon199?a] Richard Jay Solomon. Anything you can do, i can do better, i can do anything better than you. … Technical report, MIT Reaserch Program on Communication Policy, 199?

[sonah1995a] B. Sonah, M. R. Ito, and G. Neufeld. The design and performance of a multimedia server for high-speed networks. In Proceedings of the International Conference on Multimedia Computing and Systems, pages 15–22, Washington, D.C., May 1995. IEEE Computer Society Press.

[spafford1989a] Eugene H. Spafford. The internet worm program: An analysis. Purdue Technical Report CSD-TR-823, Department of Computer Sciences, Purdue University, 1989.

[spencer19??a] Henry Spencer. How to steal code or inventing the wheel only once. Zoology Computer Systems, University of Toronto.

[stabell-kulo1995a] Tage Stabell-Kulø. Some hints on the theory and practice of authentication in distributed systems. Technical report, Department of Computer Science, University of Tromsø, November 1995. Draft.

[stabell-kulo1995b] Tage Stabell-Kulø. Secure communications in FRTP. Computer Science Technical Report 95-01, Department of Computer Science, University of Tromsø, February 1995.

[stabell-kulo1997a] Tage Stabell-Kulø. Security and log structured file systems. Operating Systems Review, 31(2):9–10, April 1997.

[stabell-kulo1997b] Tage Stabell-Kulø, Arne Helme, and Gianluca Dini. Key-dependency vulnerabilities in cryptographic protocols. In NIK’97, 1997.

[stabell-kuloe2004a] Tage Stabell-Kulø, T

le Segtan Skogan, and Per Harald Myrvang. The design and implementation of Obol. In Submitted Usenix 2004, 2004.

[stallings1994a] William Stallings. Confidential speaking. LAN Magazine, 9(8), August 1994.

[steen1997a] Maarten van Steen, Philip Homburg, and Andrew S. Tanenbaum. The architectural design of Globe: A wide-area distributed system. Technical Report IR-422, Faculty of Mathematics and Computer Science, Vrije Universiteit, March 1997.

[steen1998a] Maarten van Steen, Andrew S. Tanenbaum, Ihor Kuz, and Henk J. Sips. A scalable middleware solution for advanced wide-area web services. In Procceding of Middleware’98, The Lake District, UK, September 1998.

[stefani1998a] Jean-Bernard Stefani, Bruno Dumant, Frédéric Dang Tran, and François Horn. The RETINA DPE kernel: A flexible, real-time ORB framework. In Sebastiano Trigila, Al Mullery, Mario Camploargo, Hans Vanderstraeten, and Marcel Mampaey, editors, Intelligence in Services and Networks: Technology for Ubiquitous Telecom Services, volume 1430 of Lecture Notes in Computer Science, pages 287–296, Antwerp, Belgium, May 1998. Springer-Verlag.

[steinbakk1991a] Line Steinbakk, Lars Kristian Vognild, and Svein Arne Solbakk. Networking requirements from distributed multimedia applications. Report R 0526, Forut, IT, June 1991.

[steinbakk1993a] Line Steinbakk. Konferansekontroll i en telemedisinsk multimedia arbeidsstasjon. Notat IT3032/3-93, Norut IT, December 1993.

[steiner1988a] Jennifer G. Steiner, Clifford Neuman, and Jeffrey Schiller. Kerberos: An authentication service for open networks systems. In Proceedings of Usenix Winter Conference 1988, pages 191–202, February 1988.

[steinmetz1995b] Ralf Steinmetz. Analyzing the multimedia operating system. IEEE Multimedia, 2(1):68–84, Spring 1995.

[steinmetz1995c] Ralf Steinmetz and Lars C. Wolf. Evaluation of a CPU scheduling mechanism for synchronized multimedia streams. In Heinz Beilner and Falko Bause, editors, Quantitative Evaluation of Computing and Communication Systems, Performance Tools ’95 and MMB ’95, volume 977 of Lecture Notes in Computer Science, pages 1–10, Heidelberg, Germany, September 1995. Springer-Verlag.

[stephen1992a] Graham A. Stephen. String search. Technical Report TR-92-gas-01, School of Electronic Engineering Science, University College of North Wales, October 1992.

[sterbenz1995a] James P. G. Sterbenz. Protocols for high speed networks: Life after ATM? In Gerald Neufield and Mabo Ito, editors, Protocols for High Speed Networks IV, pages 3–18. IFIP, Chapman & Hall, 1995.

[stiller1995a] Burkhard Stiller. Quality-of-service issues in networking environments. Technical report, University of Cambridge, Computer Laboratory, September 1995.

[stiller1995b] Burkhard Stiller. A framework for QoS updates in a networking environment. Technical Report TR 374, University of Cambridge, Computer Laboratory, July 1995.

[stiller1995c] Burkhard Stiller. CogPiT — configuration of protocols in TIP. Technical Report TR 368, University of Cambridge, Computer Laboratory, June 1995.

[strunk1918a] William Strunk, Jr. Elements of Style. Ithaca, New York, 1918.

[stuttgen1995a] Heinrich J. Stüttgen. Network evolution and multimedia communication. IEEE Multimedia, 2(3):42–59, Fall 1995.

[sun1998a] Sun Microsystems. Java core reflection. Technical report, Sun Microsystems, Inc., 1998.

[sun1999a] Sun Microsystems. Jini technology executive overview. Technical report, Sun Microsystems, Inc., January 1991.

[sun1999b] Sun Microsystems. Jini architecture specification. Technical report, Sun Microsystems, Inc., January 1991.

[sun1999c] Sun Microsystems. Simplified guide to the Java 2 platform, enterprise edition. Technical report, Sun Microsystems, Inc., 1991.

[svobodova1984a] Liba Svobodova. File servers for network-based distributed systems. ACM Computing Surveys, 16(4):353–398, December 1984.

[swick19??a] Ralph R. Swick and Mark S. Ackerman. The X toolkit: More bricks for building user-interfaces -or- widgets for hire. Project Athena.

[sydir1998a] J. Sydir, S. Chatterjee, and B. Sabata. Providing end-to-end QoS assurances in CORBA-based systems. In Proceedings of the First IEEE International Symposium on Object-Oriented Real-Time Distributed Compuring, Kyoto, Japan, April 1998.

[synnes1997a] Kåre Synnes, Serge Lachapelle, Peter Parnes, and Dick Schefström. Distributed education using the mStar environment. Research Report 1997:25, Department of Computer Science/Center for Distance-spanning Technology, Luleå University of Technology, Sweden, December 1997.

[tanenbaum1988a] Andrew Tanenbaum. Brief description of the state-of-the-art in operating systems, 1988.

[tanenbaum1990b] Andrew S. Tanenbaum, Robbert van Renesse, Hans van Staveren, Gregory J. Sharp, Sape J. Mullender, Jack Jansen, and Guido van Rossum. Experiences with the Amoeba distributed operating system. Communications of the ACM, 33(12):46– 63, December 1990.

[tanenbaum1992b] Andrew S. Tanenbaum. Distributed operating systems anno 1992, what have we learned so far? In Distributed Computing — Practice and Experience, Proceedings of the Autumn 1992 OpenForum Techical Conference, pages 1–14, Utrecht, The Netherlands, November 1992. EurOpen and UniForum.

[tanenbaum1993a] Andrew S. Tanenbaum. The amoeba distributed operating system. Technical report, Vrije Universiteit, 1993.

[tarr1999a] P. Tarr, H. Ossher, W. Harrison, and S.M. Sutton, Jr. N degrees of separation: Multi-dimensional separation of concerns. In Proceedings of the International Conference on Software Engineering, May 1999.

[tassel1997a] Jérôme Tassel, Bob Briscoe, and Alan Smith. An end to end price-based QoS control component using Reflective Java. In André Danthine and Christophe Diot, editors, From Multimedia Services to Network Services, Proceedings of the 4th International COST 237 Workshop, volume 1356 of Lecture Notes in Computer Science, pages 18–32, Lisboa, Portugal, December 1997. Springer-Verlag.

[tassel1997b] Jérôme Tassel, Bob Briscoe, and Alan Smith. An adaptive QoS architecture for Internet multimedia application using Reflective Java. Technical report, BT Advanced Research & Technology, UK, 1997.

[tenet1995a] Tenet Group. Recent and current research. Technical report, University of California, Berkeley and International Computer Science Institute, May 1995.

[teraoka1990a] Fumio Teraoka, Yasuhiko Yokote, and Mario Tokoro. Virtual network: Towards location transparent communication i large distributed systems. Technical Report SCSL-TR-90-005, Sony Computer Science Laboratory Inc., June 1990.

[tevanian1989a] Avadis Tevanian, Jr. and Ben Smith. Mach: The model for future Unix. Byte, pages 411–416, November 1989.

[tezuka1996a] Hiroshi Tezuka and Tatsuo Nakajima. Simple continuous media storage server on Real-Time Mach. In Proceedings of the USENIX 1996 annual technical conference, pages 87–98, Berkeley, CA, USA, January 1996. USENIX.

[thai2001b] Thuan Thai. Embracing the web. Technical report, O’Reilly Associates, Inc, July 2001.

[thatte1986a] Satish R. Thatte. Towards a semantic theory for equational programming languages. In Proceedings of the 1986 ACM Conference on Lisp and Functional Programming, pages 332–342, Cambridge, Massachusetts, August 1986. ACM.

[thibault1999a] Scott Thibault, Jérôme Marant, and Gilles Muller. Adapting distributed applications using extensible networks. In Proceedings of the 19th IEEE International Conference on Distributed Computing Systems, Austin, Texas, USA, June 1999.

[thomas1989a] Dave Thomas. What’s in an object? Byte, pages 231–240, March 1989.

[thompson1984a] Ken Thompson. Reflections on trusting trust. Communications of the ACM, 27(8):761–763, August 1984.

[thompson1986a] Simon Thompson. Laws in Miranda. In Proceedings of the 1986 ACM Conference on Lisp and Functional Programming, pages 1–12, Cambridge, Massachusetts, August 1986. ACM.

[thompson1989a] Simon Thompson. Interactive functional programs. Technical report, Computing Laboratory, University of Kent at Canterbury, U.K., September 1989.

[thompson1989b] Tom Thompson. The next step. Byte, pages 265–269, March 1989.

[thompson1998a] Craig Thompson, editor. Workshop Report, Monterey, California, January 1998.

[thompson1998b] Craig Thompson, Ted Linden, and Bob Filman. Thoughts on OMA-NG: The next generation OMG object management architecture (summary). In Workshop on Compositional Software Architectures, Monterey, California, January 1998. (position paper).

[tikka1995a] Sami Tikka, Jaripekka Salminen, and Antti Ylä-Jääski. A real-time vide-on-demand system on World Wide Web and ATM networks. In F. Williams, H. Brä, and J. Nolan, editors, Proceedings of the Fourth International Conference Broadband Island ’95, Global Broadband & Beyond, pages 331–340, Dublin, Ireland, September 1995.

[tokoro1990a] Mario Tokoro. Computational field model: Toward a new computing model/methodology for open distributed environment. Technical Report SCSL-TR-90-006, Sony Computer Science Laboratory Inc., June 1990.

[tokoro1991a] Mario Tokoro. Towards operating systems for the 2000’s. Technical Report SCSL-TR-91-005, Sony Computer Science Laboratory Inc., May 1991.

[tokuda1990a] Hideyuki Tokuda, Tatsuo Nakajima, and Prithvi Rao. Real-Time Mach: Towards a predictable real-time system. In Mach Workshop Conference Proceedings, Carnegie-Mellon University, October 1990. USENIX.

[tokuda1993a] Hide Tokuda and Takuro Kitayama. Dynamic QOS control based on real-time threads. In D. Shepherd, G. Blair, G. Coulson, N. Davies, and F. Garcia, editors, Proceedings of the 4th International Workshop on Network and Operating System Support for Digital Audio and Video, volume 846 of Lecture Notes in Computer Science, pages 114–123, Lancaster, U.K., November 1993. Springer-Verlag.

[tol1992a] Ronald M. Tol. A small real-time kernel proven correct. In Real-Time Systems Symposium, pages 227–230, 1992.

[torbjornsen1988a] Øystein Torbjørnsen. Communication in a failsoft hypercube database machine. In Proceedings of NIK’88, 1988.

[torbjornsen1990a] Øystein Torbjørnsen. Low communication cost parallel join. In Proceedings of NIK’90, 1990.

[toutain1995a] François Toutain and Laurent Toutain. Network support for multimedia communications using distributed media scaling. In D. Hutchison, H. Christiansen, G. Coulson, and A. Danthine, editors, Teleservices and Multimedia Communications — Second International COST 237 Workshop, volume 1052 of Lecture Notes in Computer Science, pages 138–157, Copenhagen, Denmark, November 1995. Springer-Verlag.

[tran1995a] F. Dang Tran and V. Perebaskine. TORBoyau: Architecture and implementation. Note Technique NT/PAA/TSA/TLR/4587, CNET, Centre Paris A, December 1995.

[treese19??a] G. Winfield Treese. Berkley UNIX on 1000 workstations: Athena changes to 4.3BSD. Project Athena.

[trompette1995a] Christophe Trompette and François Merciol. Introducing muldimedia in COOL OODS: An ODP-based approach. In Proceedings of European Research Seminar on Advances in Distributed Systems, ERSADS’95, L’Alpe-d’-Huez, France, April 1995.

[troya1997a] José M. Troya and Antonio Vallecillo. On the addition of properties to components. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Component-Oriented Programming, volume 1357 of Lecture Notes in Computer Science, pages 374–378, Jyväskylä, Finland, June 1997. Springer-Verlag.

[troya1998a] José M. Troya and Antonio Vallecillo. A reflective component model for open systems. In ECOOP’98 Workshop on Reflective Object-Oriented Programming Systems, Brussels, Belgium, July 1998.

[truyen1999a] Eddy Truyen, Bart Vanhaute, Bert Robben, and Wouter Joosen. Integrating flexible middleware solutions with applications through non-functional policies. In Walter Cazzola, Robert J. Stroud, and Francesco Tisato, editors, OOPSLA’99 Workshop on Object Oriented Reflection and Software Engineering, pages 85–93, Denver, Colorado, US, November 1999.

[turek1992a] John Turek and Dennis Shasha. The many faces of consensus in distributed systems. Computer, pages 8–17, June 1992.

[turn1977a] Rein Turn. Impementation of privacy protection requirements. In Bruce Gilchrists, editor, Information Processing 77 — Proceedings of IFIP Congress 77, volume 7 of IFIP Congress Series, Toronto, August 1977. North-Holland Publishing Company.

[turner1985a] D. A. Turner. Miranda: A non-strict functional language with polymorphic types. In Nancy France, editor, Proceedings IFIP International Conference on Functional Programming Languages and Computer Architecture, volume 201 of Lecture Notes in Computer Science. Springer-Verlag, September 1985.

[turner1993a] Kenneth J. Turner. Introduction to LOTOS. In Kenneth J. Turner, editor, Using Formal Description Techniques, Wiley Series in Communication and Distributed Systems, chapter 3, pages 47–83. John Wiley & Sons, 1993.

[ugai1996a] Takanori Ugai and Motoshi Saeki. Reflective extensions to LOTOS and its applications. Research Report ISIS-RR-96-9E, International Institute for Advanced Study of Social Information Science, June 1996.

[valencia199?a] Andrew Valencia. An overview of the VSTa microkernel. Technical report, Valencia Consulting, 199?

[varhol1994a] Peter D. Varhol. Small kernels hit it big. Byte, pages 119–128, January 1994.

[various1988a] from the Net Various. Virus + worm: The internet worm. Collected from different newsgroups/postings on the Net, November 1988.

[various1992a] Various, Department of Computer Science, University College London. Lessons and challenges of distributed computing — culled from the network. Posted on ‘comp.os.research’, 1992.

[vecchi1995a] Mario P. Vecchi. Broadband networks and services: Architecture and control. IEEE Communications Magazine, 33(8):24–32, August 1995.

[venkatasubramanian1998a] Nalini Venkatasubramanian, Gul Agha, and Carolyn Talcott. Composable QoS-based distributed resource management. In Workshop on Compositional Software Architectures, Monterey, California, January 1998. (position paper).

[venkatesh1995a] D. Venkatesh and T. D. C. Little. Dynamic service aggregation for efficient use of resources in interactive video delivery. In 5th International Workshop on Network and Operating System Support for Digital Audio and Video, NOSSDAV’95, pages 119–122, Durham, New Hampshire, April 1995.

[viega1998a] John Viega, Barry Warshaw, and Ken Manheimer. Mailman: The GNU mailing list manager. In Proceedings of the 12th Usenix LISA conference, Bostaon, MA, US, December 1998.

[villazon1999a] Alex Villazón and Jarle Hulaas. Mobility, reflection and meta-objects protocols. In Walter Cazzola, Robert J. Stroud, and Francesco Tisato, editors, OOPSLA’99 Workshop on Object Oriented Reflection and Software Engineering, pages 51–62, Denver, Colorado, US, November 1999.

[vin1994a] Harrick M. Vin. Multimedia system architecture. In Proceedings of the SPIE International Symposium on Photonics for Industrial Applications, Critical Review on Defining Global Information Infrastructure: Infrastructure, Systems, and Services, Boston, US, November 1994.

[vogel1995a] Andreas Vogel, Brigitte Kerhervé, Gregor von Bochman, and Jan Gecsei. Distributed multimedia and QOS: A survey. IEEE Multimedia, 2(2):10–19, Summer 1995.

[vogel1995b] Andreas Vogel. Distributed multimedia research at the DSTC. IEEE Multimedia Newsletter, 3(1), December 1995. http://ada.computer.org/tab/tcmm/.

[vogt1993a] Martin Vogt, Berhard Plattner, Thomas Plagemann, and Thomas Walter. A runtime environment for Da CaPo. In Proceedings of INET’93, International Networking Conference Internet Society, San Francisco, August 1993.

[vogt1995a] Carsten Vogt. Quality-of-service managment for multimedia streams with fixed arrival periods and variable frame sizes. Multimedia Systems, 3(2):66–75, May 1995.

[voydock1983a] Victor L. Voydock and Stephen T. Kent. Security mechanisms in high-level network protocols. ACM Computing Surveys, 15(2):135–171, June 1983.

[wadler1985a] Philip Wadler. How to replace failure by a list of successes — a method for exception handling, backtracking, and pattern matching in lazy functional languages. In Nancy France, editor, Proceedings IFIP International Conference on Functional Programming Languages and Computer Architecture, volume 201 of Lecture Notes in Computer Science. Springer-Verlag, September 1985.

[wadler1990a] Philip Wadler. Linear types can change the world! In Proceedings IFIP International Conference on Functional Programming Languages and Computer Architecture, pages 546–566, 1990.

[walicki1994a] Michal Walicki and Sigurd Meldal. Multialgebras, power algebras and complete calculi of identities and inclusion. In Egidio Astesiano and Gianna Reggio, editors, Recent Trends in Data Type Specification, volume 906 of Lecture Notes in Computer Science, pages 453–468, S. Margherita, Italy, May/June 1994. Springer-Verlag.

[walicki1995a] Michal Walicki and Sigurd Meldal. Generated models and the ω-rule: the nondeterministic case. In Peter D. Mosses, Mogne Nielsen, and Michael I.Schwartzbach, editors, TAPSOFT’95: Theory and Practice of Software Development, volume 915 of Lecture Notes in Computer Science, pages 424–438, Aarhus, Denmark, May 1995. Springer-Verlag.

[walker1983a] Bruce Walker, Gerald Popek, Robert English, Charles Kline, and Greg Thiel. The LOCUS distributed operating system. In Proceedings of the Ninth Symposium on Operating System Principles, pages 49–70, October 1983. (in Operating System Review, vol. 17 no. 5, October 1993).

[walpole19??a] Jonathan Walpole, Jon Inouye, and Ravindranath Konuru. Modularity and interfaces in micro-kernel design and implementation: A case study of Chorus on the HP PA-RISC. Department of Computer Science and Engineering, Oregon Graduate Institute of Science & Technology.

[wang1995a] Farn Wang. Reachability analysis at protocol level through timing coincidence. In Insup Lee and Scott A. Smolka, editors, CONCUR ’95: Concurrency Theory, 6th International Conference, volume 962 of Lecture Notes in Computer Science, pages 284–298, Philadelphia, PA, USA, August 1995. Springer-Verlag.

[wang1998a] Zhenyu Wang. Separate application logic from architectural concerns — beyond object services and frameworks. In Workshop on Compositional Software Architectures, Monterey, California, January 1998. (position paper).

[wang2000a] Nanbor Wang, Kirthika Parameswaran, Michael Kircher, and Douglas C. Schmidt. Applying reflective techniques to optimize a QoS-enabled CORBA component model implementation. In The 24th Annual International Computer Software and Applications Conference, Taipei, Taiwan, 2000.

[wang2000b] Nanbor Wang, Douglas C. Schmidt, and Carlos O’Ryan. Overview of the CORBA component model. In Component-Based Software Engineering: Putting the Pieces Together, chapter 38. Addison-Wesley, 2000.

[wang2001b] Nanbor Wang, Douglas C. Schmidt, Michael Kircher, and Kirthika Parameswaran. Adaptive and reflective middleware for QoS-enabled CCM applications. IEEE Distributed Systems Online, 2(5), 2001.

[watanabe1988a] Takuo Watanabe and Akinori Yonezawa. Reflection in an object-oriented concurrent language. In OOPSLA ’88 Proceeding, volume 28 of Sigplan Notices, pages 306–315. ACM Press, September 1988.

[watanabe1990a] Takuo Watanabe and Akinori Yonezawa. An actor-based metalevel architecture for group-wide reflection. In J. W. de Bakker, W. P. de Roever, and G. Rozenberg, editors, The Proceedings of REX School/Workshop Foundations of Object-Oriented Languages, volume 489 of Lecture Notes in Computer Science, pages 405–425, Noordwijkerhout, The Netherlands, May/June 1990. Springer-Verlag.

[watters1995a] Aaron Watters. The what, why, who, and where of Python. Unix/World Online: Tutorial Article, (5), 1995. (last updated January 25, 1997).

[wegdam2000a] M. Wegdam, D.-J. Plas, A. van Halteren, and B. Nieuwenhuis. Using reflection in a managemenr architecture for CORBA. In Proceedings of the 11th IFIP/IEEE International Workshop on Distributed Systems, Operations & Management (DSOM 2000), Austin, Texas, USA, December 2000.

[wegner1983a] Peter Wegner and Scott A. Smolka. Processes, tasks, and monitors: A comparative study of concurrent programming primitives. IEEE Transactions on Software Engineering, SE-9(4):446–462, July 1983.

[wegner1989a] Peter Wegner. Learning the language. Byte, pages 245–253, March 1989.

[wei1992a] Liming Wei, FongChing Liaw, Deborah Estrin, Allyn Romanow, and Tom Lyon. Analysis of a resequencer model for multicast over ATM networks. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 209–220, La Jolla, California, USA, November 1992. Springer-Verlag.

[welch98a] Ian Welch and Robert Stroud. Adaptation of connectors in software architectures. In ECOOP’98 Workshop on Component-Oriented Programming, July 1998. Position Paper.

[white1997a] Paul P. White and John Crowcroft. The integrated services in the internet: State of the art. Proceedings of the IEEE, 85(12), December 1997.

[widmer1999a] Bernfried Widmer and Wolfgang Lugmayer. A comparison of three corba management tools. In Engineering Distributed Objects Workshop, ICSE’99, Los Angeles, May 1999. Position Paper.

[wileden1998a] Jack Wileden and Alan Kaplan. Middleware as underwear: Toward a more mature approach to compositional software development. In Workshop on Compositional Software Architectures, Monterey, California, January 1998. (position paper).

[wilkes1990a] Maurice V. Wilkes. Computer security in the business world. Communications of the ACM, 33(4):399–401, April 1990.

[williams1991a] Neil Williams, Gordon S. Blair, and Nigel Davis. Distributed multimedia computing: An assessment of the state of the art. Information Services and Use, 11:265–281, 1991.

[williams1994a] Sara Williams and Charlie Kindel. The component object model: A technical overview. Dr. Dobb’s Journal, December 1994.

[wilson19??a] Paul R. Wilson. Operating system support for small objects. Department of Computer Science, University of Texas at Austin.

[wing1997a] Jeanette M. Wing. Subtyping for distributed object stores. In Howard Bowman and John Derrick, editors, IFIP TC6 WG6.1 International Workshop on Formal Methods for Open Object-based Distributed Systems (FMOODS ’97), volume 2, pages 305–318, Canterbury, Kent, UK, July 1997. International Federation for Information Processing (IFIP), Chapman & Hall.

[witten1987a] Ian H. Witten. Computer (in)security: Infiltrating open systems. Abacus, 4(4):7–25, 1987.

[wittenburg1994a] T. M. Wittenburg and T. D. C. Little. An adaptive document management system for shared multimedia data. In Proceedings of the 1st IEEE International Conference on Multimedia Computing and Systems, ICMCS’94, pages 245–254, Boston, May 1994.

[wittig1994a] Hartmut Wittig, Lars C. Wolf, and Carsten Vogt. CPU utilization of multimedia processes: HeiPOET — the Heidelberg predictor of execution times measurement tool. In Ralf Steinmetz, editor, Multimedia: Advanced Teleservices and High-Speed Communication Architectures, Second International Workshop, IWACA ’94, volume 868 of Lecture Notes in Computer Science, pages 92–103, Heidelberg, Germany, September 1994. Springer-Verlag.

[wobber1993a] Edward Wobber, Martín Abadi, Mike Burrows, and Butler Lampson. Authentication in the Taos operating system. SRC Research Reports 117, DEC’s System Research Center, December 1993.

[wohlrab1997a] Lutz Wohlrab. Configuration and adaption management for object-oriented operating systems. In Jan Bosch and Stuart Mitchell, editors, ECOOP’97 Workshops — Object-Orientation and Operating Systems, volume 1357 of Lecture Notes in Computer Science, pages 524–530, Jyväskylä, Finland, June 1997. Springer-Verlag.

[wolf1997a] Lars C. Wolf, Carsten Griwodz, and Ralf Steinmetz. Multimedia communication. Proceedings of the IEEE, 85(12), December 1997.

[wolf199?a] Lars C. Wolf, Ralf Guido Herrtwich, and Luca Delgrossi. Filtering multimedia data in reservation-based internetworks. Technical report, IBM European Networking Center, 199?

[wragg1994a] David Wragg. Design of a more usable (operating) system. Posted on ‘comp.os.research’, May 1994.

[wray1994a] Stuart Wray, Tim Glaubert, and Andy Hopper. Networked multimedia: The Medusa environment. IEEE Multimedia, 1(4):54–63, Winter 1994.

[wu1997a] Zhixue Wu. Design and implementation of a transaction service for Java. Technical Report APM.1923.00.01, ANSA Phase III, APM, January 1997.

[wu1997b] Zhixue Wu and Scarlet Schwiderski. Reflective Java: Making Java even more flexible. External Paper APM.1936.02, ANSA Phase III, APM, February 1997.

[xerox1997a] Xerox Parc Aspect-Oriented Programming Project. Aspect-oriented programming. Technical report, Xerox Parc, 1997.

[xiao1999a] Xipeng Xiao and Lionel M. Ni. Internet QoS: A big picture. IEEE Network, 13(2):8–18, March/April 1999.

[yang1995a] Zhonghua Yang and Keith Duddy. Distributed object computing with CORBA. DSTC Technical Report 23, CRC for Distributed Systems Technology, June 1995. Preliminary Version.

[yeadon1993a] Nicholas J. Yeadon. Supporting quality of service in multimedia communications via the use of filters. Lancaster University Report MPG-94-10, MPG Research Group, Computing Department, Lancaster University, December 1993.

[yeadon1994a] Nicholas Yeadon, Francisco García, Andrew Campbell, and David Hutchison. QoS adaption and flow filtering in ATM networks. In Ralf Steinmetz, editor, Multimedia: Advanced Teleservices and High-Speed Communication Architectures, Second International Workshop, IWACA ’94, volume 868 of Lecture Notes in Computer Science, pages 191–202, Heidelberg, Germany, September 1994. Springer-Verlag.

[yeadon1996a] Nicholas Yeadon, Andreas Mauthe, Francisco García, and David Hutchison. QoS filters: Addressing the hetrogenity gap. In Berthold Butscher, Eckhard Moeller, and Herwart Pusch, editors, Interactive Distributed Multimedia Systems and Services — European Workshop IDMS’96, volume 1045 of Lecture Notes in Computer Science, pages 227–243, Berlin, Germany, March 1996. Springer-Verlag.

[yen1994a] Wei Yen and Ian F. Akyildiz. On the synchronization mechanisms for multimedia integrated services networks. In D. Hutchison, A. Danthine, H. Leopold, and G. Coulson, editors, Multimedia Transport and Teleservice, International COST 237 Workshop, volume 882 of Lecture Notes in Computer Science, pages 168–184, Vienna, Austria, November 1994. Springer-Verlag.

[yeo1995a] Boon-Lock Yeo and Bede Liu. A unified approach to temporal segmentation of motion JPEG and MPEG compressed video. In Proceedings of the International Conference on Multimedia Computing and Systems, pages 81–88, Washington, D.C., May 1995. IEEE Computer Society Press.

[yeo1995b] Boon-Lock Yeo, Minerva M. Yeung, Wayne Wolf, and Bede Liu. Theft-resistant video browsing using filtered versions of compressed sequences. In Proceedings of the International Conference on Multimedia Computing and Systems, pages 50–55, Washington, D.C., May 1995. IEEE Computer Society Press.

[yokote1992a] Y. Yokote. The Apertos reflective operating system: The concept and its implementation. In Proceedingss of OOPSLA’92, volume 28 of Sigplan Notices, pages 414–434. ACM Press, 1992.

[zhang1992a] Hui Zhang and Tom Fisher. Preliminary measurements of the RMTP/RTIP. In P. Venkat Rangan, editor, Proceedings of the Third International Workshop on Network and Operating System Support for Digital Audio and Video, volume 712 of Lecture Notes in Computer Science, pages 185–196, La Jolla, California, USA, November 1992. Springer-Verlag.

[zinky1997a] John A. Zinky, David E. Bakken, and Richard D. Schantz. Architectural support for quality of service for CORBA objects. Theory and Practice of Object Systems, 3(1), 1997.

[znati1995a] Taieb Znati, Robert Simon, and Brian Field. A network-based scheme for synchronization of multimedia streams. In Proceedings of the International Workshop Multimedia Synchronization, Tysons Corner, Virginia, May 1995.